### v1

# Extension principles for the Einstein Yang–Mills system

Abstract.
We establish an extension principle for the spherically symmetric Einstein Yang–Mills system (SSEYM) with $H^{1}$ data. Based on this result, we further prove an extension theorem for developments of weighted $H^{1}$ data. In particular, the weighted $H^{1}$ space allows Hölder continuous data. Therefore, our result is consistence with the conjecture that well-posedness holds for Hölder continuous data (with sufficiently many angular derivatives) in vacuum where the axis is included.
In contrast to a massless scalar field, the purely magnetic Yang–Mills field in spherical symmetry satisfies a wave-type equation with singular potential. The proof of Christodoulou [10], based on an $L^{\infty}-L^{\infty}$ estimate, fails for the Yang–Mills scenario. Instead, we employ the $L^{2}$ method, which works for the (massless or massive) scalar matter field as well.
1. Introduction
In a series of works [9, 10, 11, 13], Christodoulou proved the weak cosmic censorship conjecture [^1] Generic asymptotically flat initial data have a maximal future development possessing a complete future null infinity[12]. for the spherically symmetric Einstein massless scalar field system. It was further demonstrated by Li-Liu [16, 17] and An [1] under gravitational perturbations without symmetries. Recently, An-Tan [2] gave a proof for the system of Einstein charged massless scalar field in spherical symmetry as well.
The instability mechanic of naked singularity in [13] was found out in the topology of bounded variation (BV). From the other perspective, the local existence theory and further the small data global existence in the BV topology were both verified ahead [10]. As shown in [10, 13], the proofs of these (local or global) existence or instability statements were all built on a sharp criterion for $C^{1}$ extension. Along the line of [10], extension criterions were verified as well in the context of Einstein charged massless scalar field and Einstein massless Vlasov system in spherical symmetry [2, 20].
We are now concerned with the Einstein Yang–Mills system, which is significant for various reasons. From a physical perspective, it triggers a wealth of fascinated gravitational collapses, as seen, for instance, [3, 6, 5] for several numerical results. On the other hand, we will show in this paper that the mathematical techniques in the treatment of Einstein Yang–Mills system will be significantly different from those for gravity coupled with massless matter fields [10, 20, 2].
In this paper, we initiate the study of extension principle for the Einstein Yang–Mills system in spherical symmetry. We will first consider the characteristic problem for the spherically symmetric Einstein Yang–Mills (SSEYM) system where initial data are prescribed on a future light cone emanating from vertex at the center of symmetry. In this setting, we demonstrate the local existence theorem for data with a low-regularity norm ($H^{1}$ initial data). We also deduce an extension principle for developments of $H^{1}$ data. Our second result is an extension theorem in a weaker topology. We show that if the interior is vacuum (and hence Minkowski), the spacetime is extendable in the topology of weighted $H^{1}$ space. Following theses extension theorems, the Cauchy stabilities are justified for the Einstein Yang–Mills system under spherically symmetric perturbations, which are initially small with respect to the above norms respectively.
The weighted $H^{1}$ space allows the class of Hölder continuous data, or more precisely, data that are smooth except at one point and Hölder continuous there. It was conjectured by Rodnianski–Shlapentokh-Rothman [21] that well-posedness holds for Hölder continuous data with sufficiently many angular derivatives in vacuum (without symmetries) where the axis is included. Our extension theorem thus supports this conjecture. Previously, in the context of spherically symmetric massless matter field [10, 2], this conjecture is true in BV topology, where the instabilities of naked singularities are stated. However, as will be discussed later, the method used in massless case fails for Einstein–Yang–Mills scenario and hence a well-posed result in BV topology is not expected. The new $L^{2}$-based method we propose in this paper has a closer nature to the vacuum context, and may be relevant for future progress in vacuum.
1.1. Statement of the main results
Let $\Gamma$ be the set of fixed points of the symmetry group $SO(3)$, and $r$ be the area function of the group orbit, a $2$-sphere. Let $m$ be the Hawking mass contained within the sphere. We introduce the double null coordinates $(u,\,\underline{u})$, and denote $C_{u}$ and $\underline{C}_{\underline{u}}$ the incoming and outgoing null cones. For any $u_{0}<\underline{u}_{\ast}\in\mathbb{R}$, we define the domain of the quotient spacetime as
$\mathcal{D}(u_{0};\,\underline{u}_{\ast}):={}$$\{(u,\,\underline{u})\,|\,u_{0}\leq u<\underline{u}_{\ast},\,\,u\leq\underline{u}<\underline{u}_{\ast}\},$ (1.1)
$\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast}):={}$$\{(u,\underline{u})\,|\,u_{0}\leq u\leq u_{1},\,\,u\leq\underline{u}\leq\underline{u}_{\ast}\}.$ (1.2)
As shown in their definitions, $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$ is the domain enclosed by the axis $\Gamma$ and the hypersurfaces $C_{u_{0}}$, $\underline{C}_{\underline{u}_{\ast}}$, while $\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast})$ is the domain bounded by the axis $\Gamma$ and the hypersurfaces $C_{u_{0}}$, $C_{u_{1}}$, $\underline{C}_{\underline{u}_{\ast}}$.
We formulate the SSEYM system with the gauge group $SU(2)$ in Section 2.1, where we assume the spherically symmetric purely magnetic ansatz for the Yang–Mills potential (2.2). The characteristic problem involves prescribing a $C^{0}$ function $\partial_{r}w|_{C_{u_{0}}}$ on an initial outgoing cone $C_{u_{0}}$, where $w$ is related to the Yang–Mills potential (2.2). Throughout this paper, the $C^{0}$ Yang–Mills solution always refers to $\partial_{r}w\in C^{0}$.
Remark 1.1.
Without loss of generality, we assume $w|_{\Gamma}=1$ and set
$\hat{w}:=w-1.$ (1.3)
Then
$\frac{\hat{w}}{r}=\frac{1}{r}\int_{0}^{r}\partial_{r}w\,\mathrm{d}r^{\prime}|_{C_{u}}.$
It suggests choosing the initial datum $w|_{C_{u_{0}}}$ can be replaced by prescribing $\partial_{r}w|_{C_{u_{0}}}$.
Let us begin our statement with the $H^{1}$ case.
1.1.1. Statement for the $H^{1}$ case
Fixing a constant
$0\leq b<\frac{\sqrt{89}-9}{4},$ (1.4)
we define the $H^{1}$ norm for $\partial_{r}w$
$\|\partial_{r}w\|_{H^{1}}:=\int_{C_{u}}\frac{1}{r^{1+b}}\left(\frac{\hat{w}^{2}}{r^{2}}+(\partial_{r}w)^{2}+(r\partial_{r}(\partial_{r}w))^{2}\right)\mathrm{d}r.$ (1.5)
Remark 1.2.
It is remarkable that the norm (1.5) is equivalent to
$\int_{0}^{r}\frac{1}{r^{1+b}}\sum_{0\leq k\leq 2}\left((r\partial_{r})^{k}\frac{\hat{w}}{r}\right)^{2}\mathrm{d}r.$ (1.6)
If we choose a new parameter $s:=\ln r$, (1.6) is alternatively written as
$\int_{0}^{r_{1}}\frac{1}{r^{1+b}}\sum_{0\leq k\leq 2}\left((r\partial_{r})^{k}\frac{\hat{w}}{r}\right)^{2}\mathrm{d}r=\int_{-\infty}^{\ln r_{1}}e^{-bs}\sum_{0\leq k\leq 2}\left(\partial_{s}^{k}\frac{\hat{w}}{r}\right)^{2}\mathrm{d}s.$ (1.7)
When $b=0$, (1.7) is the standard $H^{2}$ norm for $\frac{\hat{w}}{r}$. We abuse the notation a little bit and call (1.7) the $H^{2}$ norm of $\frac{\hat{w}}{r}$, or rather $H^{1}$ norm of $\partial_{r}w$. With respect to this $H^{1}$ space, the density theorem is proved in Appendix C.
Remark 1.3.
As a further remark, the standard energy for the Yang–Mills field does not account for the definition of energy (1.5), even if $b=0$. Actually, the standard first order energy for the Yang–Mills field should be roughly in the form of $\int_{C_{u}}\left(\frac{\hat{w}^{2}}{r^{2}}+(\partial_{r}w)^{2}\right)\mathrm{d}r$, where the weight $\frac{1}{r^{1+b}}$ is absent. The motivation for (1.5) indeed comes from the right hand side of (2.14a), since after integrating (2.14a) in $\underline{u}$, we expect to bound $\frac{m}{r}$ via the $H^{1}$ (in fact $H^{0}$) norm of $\partial_{r}w$.
We present the $H^{1}$-extension theorem. One may refer to Theorem 3.1 for a comprehensive statement.
Theorem 1.4.
Suppose we have the initial data $\partial_{r}w|_{C_{u_{0}}}$ with bounded $H^{1}$ norm, and an $H^{1}$ solution (that is, $\partial_{r}w\in H^{1}$) to the SSEYM system on $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$. There is a constant $\varepsilon>0$ such that if
$\lim_{u\rightarrow\underline{u}_{\ast}}\sup_{\mathcal{D}(u;\,\underline{u}_{\ast})}\frac{m}{r}<\varepsilon^{2},$
the solution extends as an $H^{1}$ solution on $\mathcal{D}(u_{0};\,\underline{u}_{1})$ for some $\underline{u}_{1}>\underline{u}_{\ast}$.
Following the extension theorem, there is the Cauchy stability statement, for which a more precise version is given in Theorem 3.9.
Theorem 1.5.
Suppose the data $\partial_{r}w|_{C_{u_{0}}}$ admit bounded $H^{1}$ norm that is small enough. Then there is a global and unique solution on $\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{0})$ with small $H^{1}$ norm, for any $\underline{u}_{0}\in(u_{0},+\infty)$ and $u_{1}\in(u_{0},\,\underline{u}_{0}]$.
We next proceed to the weighted $H^{1}$ case.
1.1.2. Statement for the weighted $H^{1}$ case
We will consider a class of weighted energies in the domain $\mathcal{D}(u_{0},u;\underline{u}_{\ast}+\delta)$ with $u\in[u_{0},\,\underline{u}_{\ast})$ and $\delta>0$. The weighted energies are defined with the assistance of a non-negative function
$\xi:=\frac{r-r_{\ast}}{r}\quad\text{with}\quad r_{\ast}(u):=r(u,\,\underline{u}_{\ast}),$ (1.8)
for $\underline{u}\geq\underline{u}_{\ast}$. We fix a constant $\gamma\in(b,2]$ with $b$ given in (1.4) and define the following weighted $H^{1}$ norm for $\partial_{r}w$
$\int_{C_{u}\cap\{\underline{u}\geq\underline{u}_{\ast}\}}\frac{1}{r^{1+b}}\left(\xi^{-3-\gamma}\frac{\hat{w}^{2}}{r^{2}}+\xi^{-1-\gamma}(\partial_{r}w)^{2}+\xi^{1-\gamma}\left(r\partial_{r}(\partial_{r}w)\right)^{2}\right)\mathrm{d}r.$ (1.9)
Remark 1.6.
The weighted $H^{1}$ norm (1.9) is equivalent to
$\int_{\{r_{\ast}\leq r\}}\xi^{-3-\gamma}r^{-1-b}\sum_{0\leq k\leq 2}\left((\xi r\partial_{r})^{k}\frac{\hat{w}}{r}\right)^{2}\mathrm{d}r|_{C_{u}}.$ (1.10)
For the density theorem in the weighted $H^{1}$ space, please refer to Appendix C for a proof.
Theorem 1.7.
Given datum $\partial_{r}w|_{C_{u_{0}}}$ that is supported in $(\underline{u}_{\ast},\,+\infty)$ (namely, $\partial_{r}w|_{C_{u_{0}}\cap\{\underline{u}\leq\underline{u}_{\ast}\}}\equiv 0$), we assume $|r-r_{\ast}|^{-\frac{\gamma}{2}}\partial_{r}w|_{C_{u_{0}}}$ is continuous at $\underline{u}=\underline{u}_{\ast}$ and the weighted $H^{1}$ norm of $\partial_{r}w|_{C_{u_{0}}}$ on $C_{u_{0}}\cap\{\underline{u}_{\ast}\leq\underline{u}\leq\underline{u}_{\ast}+\delta\}$ for some $\delta>0$ is bounded. If $\delta$ is small enough, the solution extends as a solution with bounded, weighted $H^{1}$ norm in the domain of $\mathcal{D}(u_{0},\,u;\,\underline{u}_{\ast}+\delta)$, where $u\in[u_{0},\,\underline{u}_{\ast})$.
Remark 1.8.
We remark that our data requirement allows the initially second derivative $\partial^{2}_{r}w|_{C_{u_{0}}}$ blows up [^2] The $C^{1}$ case ($\gamma=2$) is also included. like
$(r-r_{\ast})^{-1+\frac{\gamma}{2}},\quad 0\leq b<\gamma\leq 2.$
Thus, the data in Theorem 1.7 are, in fact, Hölder continuous. Roughly speaking, Theorem 1.7 states that if the interior $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$ is Minkowski, and the data are Hölder continuous, then the solution can be extended beyond the last incoming cone $\underline{C}_{\underline{u}_{\ast}}$. Furthermore, we will show that the extended solution is still Hölder continuous when it goes across $\underline{C}_{\underline{u}_{\ast}}$. Please refer to Theorem 4.1 for a detailed description.
The next one is a Cauchy stability statement in the weighted norm space. One may refer to Theorem 4.10 for more details.
Theorem 1.9.
Given datum $\partial_{r}w|_{C_{u_{0}}}$ that is supported in $(\underline{u}_{\ast},\,+\infty)$, suppose $|r-r_{\ast}|^{-\frac{\gamma}{2}}\partial_{r}w|_{C_{u_{0}}}$ is continuous at $\underline{u}=\underline{u}_{\ast}$, and the weighted $H^{1}$ norm on $C_{u_{0}}\cap\{\underline{u}_{\ast}\leq\underline{u}\leq\underline{u}_{1}\}$, $\underline{u}_{1}<+\infty$, is small enough. Then there is a unique solution with bounded, weighted $H^{1}$ norm on $\mathcal{D}(u_{0},\,u;\,\underline{u}_{1})$, where $u\in[u_{0},\,\underline{u}_{\ast})$.
Remark 1.10.
In the subsection followed, we will sketch the main idea for the proofs. We propose the $L^{2}$ method, which is tailed for the Yang–Mills field, or rather, the wave type equation with a positive, singular potential. Indeed we can also adapt the $L^{2}$ scheme to a massless or massive scalar field so that all the preceding statements (Theorem 1.41.9) hold for the Einstein (massive or massless) scalar field system in spherical symmetry. Since our proof is independ of $C^{1}$ norm, but only depends on $H^{1}$ norm, we have in fact achieved an extension principle in a weaker topology ($H^{1}$ topology) than that in Christodoulou [10] ($C^{1}$ topology), and thus less regularity on the axis is obtained as well.
To facilitate the extension theorems, we verify the associated local existence theory for the characteristic problem of SSEYM system in Appendix A. We also state an extension principle away from the axis in Appendix B for completeness. Thus, it confirms that the “first singularity” for the SSEYM system must emanate from the axis.
1.2. Comments on the proof
The purely magnetic Yang–Mills field obeys a wave-type equation with a singular potential $\frac{1}{r}$, which makes itself significantly different from a standard massless scalar field. To highlight the difference, it is useful to compare their formulations in Minkowski space. Notice that, the massless sclar field equation in Minkowski takes the form of
$\partial_{u}\partial_{\underline{u}}(r\phi)=0,$
while the Yang–Mills equation reads
$\partial_{u}\partial_{\underline{u}}w+\frac{w(w^{2}-1)}{r^{2}}=0.$
In terms of $\hat{w}$ (1.3), the Yang–Mills equation can be reformulated as follows
$\partial_{u}\partial_{\underline{u}}\hat{w}+\frac{2}{r}\frac{\hat{w}}{r}=\frac{\hat{w}^{2}}{r^{2}}\left(\hat{w}+3\right).$
It manifests itself as a nonlinear wave equation with an additional linear, singular term
$\frac{2}{r}\frac{\hat{w}}{r},$
compared to the scalar wave equation. As a consequence, the techniques for proving the extension principles are necessarily different.
1.2.1. The $H^{1}$ case
To begin with, let us illustrate the proof for the massless scalar field [10]. Let $\alpha_{s}=\partial_{r}(r\phi)$, where $\partial_{r}$ is the null vector parallel to $\partial_{\underline{u}}$ and parametrized by $r$. The wave equation for $\phi$ leads to the following equation for $\partial_{r}\alpha_{s}$
$\partial_{u}(\partial_{r}\alpha_{s})={}$ $-2\frac{\partial_{u}r}{r}\frac{\frac{2m}{r}}{1-\frac{2m}{r}}\partial_{r}\alpha_{s}+3\frac{\partial_{u}r}{r^{2}}\frac{\frac{2m}{r}}{1-\frac{2m}{r}}r\partial_{r}\phi-\frac{\partial_{u}r}{r^{2}}\frac{1}{1-\frac{2m}{r}}(r\partial_{r}\phi)^{3}.$
Remarkably, this equation contains no linear terms on the right hand side. Namely, terms that are linear on $\partial_{r}\alpha_{s}$, $\alpha_{s}$, $r\partial_{r}\phi$ and $\frac{2m}{r}$ are absent. It is exactly this feature that allows an usage of $L^{\infty}-L^{\infty}$ estimate in proving the extension theorem [10, Theorem 5.1]. Without loss of generality, we take the first term $-2\frac{\partial_{u}r}{r}\frac{\frac{2m}{r}}{1-\frac{2m}{r}}\partial_{r}\alpha_{s}$ for instance to explain the main ingredient of this approach. In this quadratic term, the $\frac{2m}{r}$ offers smallness $\varepsilon$ so that the whole term is negligible and a type of inequality
$\sup|\partial_{r}\alpha_{s}|\leq\text{data}+\varepsilon\sup|\partial_{r}\alpha_{s}|$
yields the bound for $\partial_{r}\alpha_{s}$. In principle, this kind of $L^{\infty}-L^{\infty}$ estimate works well for ”generic” massless matter fields that share the common feature of vanishing linear terms. Along with the massless scalar field, as far as we know, the $L^{\infty}-L^{\infty}$ strategy truns out to be valid as well for the charged massless scalar field [2] and the massless Vlasov system [20].
In contrast, for the Yang–Mills scenario, due to the additional potential, the main equation for $\partial_{r}\alpha$ (2.3), $\alpha:=\partial_{r}w$, involves linear terms that are linear on $\frac{\hat{w}}{r}$ and $\alpha$ respectively on the right hand side. Therefore, the strategy of $L^{\infty}-L^{\infty}$ estimate fails in the Yang–Mills case, since we are unable to gain extra smallness from these linear terms. Nevertheless, noticing the positive sign of the potential in the Yang–Mills equation, we propose the $L^{2}$ method, which allows us to treat the quadratic terms originated from the linear terms through integration by parts. After that, terms that have no additional smallness give rise to quadratic terms with definite signs. Since our energy is not the standard energy, it is in fact $r^{-1-b}$ weighted, these quadratic terms do not necessarily admit favorable signs. Specifically, we are unable to derive any energy bound using the first-order energy identities; see (3.1.4) and (3.1.4). To proceed, the idea is, after performing integration by parts on these linear-related terms, the second order energy identity generates positive or negative quadratic terms that are potential to compensate the corresponding terms with opposite signs in the first order energy identities, see (3.1.4). Then we play with these two energy identities by assigning different weights on them to render all the quadratic spacetime integrals positive. Eventually, we obtain the $H^{1}$ energy bounds, see Section 3.1.4, and the pointwise estimates follow from a Sobolev type inequality in the $H^{1}$ space.
The scheme employed for the Yang–Mills case applies to the (massless or massive) scalar field so that an extension principle in a low regularity ($H^{1}$) can be established. Moreover, using our method, we establish the extension principle with less regularity on the axis: $r\partial_{r}\phi\sim r^{b}$ with $0\leq b<\frac{\sqrt{89}-9}{4}$, while in Christodoulou [10], $r\partial_{r}\phi$ is roughly $r$ near the axis. Besides, the $H^{1}$ energy estimate is crucial to promote a local existence theorem, Theorem A.1.
1.2.2. The weighted $H^{1}$ case
The proof in the weighted $H^{1}$ case heavily depends on the energy identities derived in the previous $H^{1}$ case and its modification through introducing a weight function $\xi$ (1.8). Remarkably, the weight function $\xi$ is designed in such a way that in the domain concerned,
$0\leq\xi<1,\quad\partial_{u}\xi\geq 0,\quad\partial_{\underline{u}}\xi\geq 0$
More specifically, noting the negative power on $\xi$ in the first order weighted energy, the sign of $\partial_{u}\xi$ and $\partial_{\underline{u}}\xi$ helps to produce positive spacetime integrals which will cancel the negative terms in the originally unweighted energy identities. Once the first order estimate is completed, the second order one follows.
1.3. Outline of this paper
This paper is organized as follows. In section 2, we formulate the SSEYM system. Section 3 is devoted to establishing the extension theorem and Cauchy stability in $H^{1}$ space. The counterparts in the weighted $H^{1}$ space are proved in Section 4. We also demonstrate the local existence theorem in Appendix A, the extension principle away from the axis in Appendix B, and provide a density theorem in Appendix C.
1.4. Acknowledgement
J.L. is supported by National Key R&D Program of China (No. 2022YFA1005400) and NSFC (12326602, 12141106). J.W. is supported by NSFC (Grant No. 12271450).
2. Preliminaries
2.1. Spherically symmetric Einstein Yang–Mills equations
In this section, we shall formulate the SSEYM system with certain gauge assumptions.
On a spherical symmetric spacetime $(\mathcal{M},g)$, there exists an isometric action of the rotational group $SO(3)$, where the orbits are $2$-dimensional spheres. Define the axis $\Gamma$ of $(\mathcal{M},g)$ to be the set of fixed points of the $SO(3)$ action. Let us introduce the double null coordinates $(u,\underline{u})$, and denote the level sets by $C_{u}$ and $\underline{C}_{\underline{u}}$, which are the incoming and outgoing null cones. The intersection of $C_{u}$ and $\underline{C}_{\underline{u}}$ is a $2$-sphere denoted by $S_{\underline{u},u}$. We define the area function $r$ by
$\text{Area}(S_{\underline{u},u})=4\pi r^{2}.$
Then the axis is characterized by
$r=0\quad\text{on}\,\,\Gamma.$
The metric on $\mathcal{M}$ takes the form of
$g=-4\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}+r^{2}\mathrm{d}\sigma_{S^{2}},$
where $\mathrm{d}\sigma_{S^{2}}$ is the standard metric on $S^{2}$ and $\Omega$ is the lapse function such that $-2\Omega^{2}=g(\partial_{u},\partial_{\underline{u}})$.
We define
$h=\Omega^{-2}\partial_{\underline{u}}r,\quad\underline{h}=\partial_{u}r.$
The Hawking mass contained within the $2$-sphere $S_{\underline{u},u}$ is defined via
$m=\frac{r}{2}\left(1+h\underline{h}\right).$
In the analysis approaching the axis, it is useful to define the mass ratio
$\mu:=\frac{2m}{r}.$ (2.1)
In what follows, we turn to the Yang–Mills field with gauge group $SU(2)$. The group $SU(2)$ is the real Lie group of unitary matrices of determinant one, and its associated Lie algebra is $su(2)$, the anti-hermitian traceless $2\times 2$ matrices. Let $\tau_{i}$, $i\in\{1,2,3\}$, be the following real basis of $su(2)$,
$\tau_{1}=\frac{i}{2}\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\,\,\tau_{2}=\frac{1}{2}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix},\,\,\tau_{3}=\frac{i}{2}\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},$
with $[\tau_{i},\,\tau_{j}]=\epsilon_{ijk}\tau_{k}$, $\epsilon_{ijk}$ being the Levi-Civita symbol, $\{i,j,k\}\subset\{1,2,3\}$. We assume the spherically symmetric purely magnetic ansatz [4, 6] for the Yang–Mills potential,
$A=w\tau_{1}\mathrm{d}\theta+\left(\cot\theta\tau_{3}+w\tau_{2}\right)\sin\theta\mathrm{d}\phi,$ (2.2)
where $(\theta,\,\phi)$ are coordinates on $S^{2}$. Then the Yang–Mills field takes
$F={}$ $\mathrm{d}A+A\wedge A$
[…]
As followed from the ansatz (2.2), $F_{u\underline{u}}=0$, namely the electric charge vanishes; $F_{AB}=\frac{w^{2}-1}{r^{2}}\tau_{3}\varepsilon_{AB}$ [^3] The capital Latin letters $\{A,B\}$ refer to indices on $S^{2}$. and we define the magnetic charge within the sphere $S_{\underline{u},u}$ to be
$Q:=w^{2}-1.$ (2.3)
The vacuum of the Yang–Mills field are given by $w=\pm 1$.
The Einstein Yang–Mills system takes the form of
$R_{\alpha\beta}-\frac{1}{2}Rg_{\alpha\beta}={}$ $2T_{\alpha\beta},$
$\hat{D}^{\mu}F_{\mu\nu}={}$ $0,$
[…]
$T_{\alpha\beta}=\left\langle F_{\alpha}{}^{\nu},\,F_{\beta\nu}\right\rangle-\frac{1}{4}g_{\alpha\beta}\left\langle F_{\sigma\rho},\,F^{\sigma\rho}\right\rangle.$
For more detailed descriptions for the Einstein Yang–Mills system, please refer to [7] and [18, 19].
We are now ready to write down the spherical symmetric Einstein Yang–Mills system (SSEYM). With the ansatz (2.2) and the double gauge, the SSEYM system consistes of the following equations for $(r,\Omega,w)$,
$\partial_{\underline{u}}h={}$ $-2\Omega^{-2}\frac{(\partial_{\underline{u}}w)^{2}}{r},$ (2.4a)
$\partial_{u}(\Omega^{-2}\underline{h})={}$ $-2\Omega^{-2}\frac{(\partial_{u}w)^{2}}{r},$ (2.4b)
$\partial_{\underline{u}}\underline{h}={}$ $-\Omega^{2}\frac{\mu}{r}+\Omega^{2}\frac{Q^{2}}{r^{3}},$ (2.4c)
$\partial_{u}(\Omega^{2}h)={}$ $-\Omega^{2}\frac{\mu}{r}+\Omega^{2}\frac{Q^{2}}{r^{3}},$ (2.4d)
$\partial_{\underline{u}}\partial_{u}\ln\Omega={}$ $\partial_{u}\partial_{\underline{u}}\ln\Omega=\Omega^{2}\frac{\mu}{r^{2}}-2\Omega^{2}\frac{Q^{2}}{r^{4}},$ (2.4e)
and the Yang–Mills equation
$\partial_{u}\partial_{\underline{u}}w+\Omega^{2}\frac{Q}{r^{2}}w=0.$ (2.5)
Moreover, we can derive from the above system the subsequent equations for the mass function
$\partial_{\underline{u}}m={}$ $\Omega^{2}h\frac{Q^{2}}{2r^{2}}-\Omega^{-2}\underline{h}(\partial_{\underline{u}}w)^{2},$ (2.6a)
$\partial_{u}m={}$ $\underline{h}\frac{Q^{2}}{2r^{2}}-h(\partial_{u}w)^{2}.$ (2.6b)
A characteristic initial data set for the SSEYM system composes of a triplet $(r,\Omega,\partial_{r}w)|_{C_{u_{0}}}$ defined on an outgoing cone $C_{u_{0}}$, see Remark 1.1. In addition to the initial condition, the SSEYM system is incorporated with the following boundary condition
$r|_{\Gamma}=0,\quad m|_{\Gamma}=0,\quad Q|_{\Gamma}=0.$ (2.7)
Note that, $Q|_{\Gamma}=0$ ensures the regularity of (2.5) on the axis.
We can reformulate the SSEYM system by eliminating $\Omega$ and replacing it by $m$, so that it is equivalent to the system of $(r,m,w)$ which is composed of
[…]
with the boundary condition (2.7).
2.2. Gauge choice
We fix $\underline{u}$ on the initial outgoing cone $C_{u_{0}}$ by requiring that $\underline{u}=2r+u_{0}$ on $C_{u_{0}}$, and it suggests
$\Omega^{2}h|_{C_{u_{0}}}=\frac{1}{2}.$ (2.8)
Next, we choose the double null coordinates $(u,\,\underline{u})$ satisfying the additional gauge condition
$u=\underline{u}\quad\text{on}\,\,\Gamma.$
As a consequence, in view of $r=0$ on $\Gamma$, it follows that
$\Omega^{2}h=-\underline{h}\quad\text{on}\,\,\Gamma.$ (2.9)
2.3. More equations
For notational convenience, we define
$\alpha:={}$ $\Omega^{-2}h^{-1}\partial_{\underline{u}}w,$ (2.10)
$\alpha^{\prime}:={}$ $r\Omega^{-2}h^{-1}\partial_{\underline{u}}\alpha,$ (2.11)
$\Phi:={}$ $\frac{Q}{r},$ (2.12)
$\Phi^{\prime}:={}$ $r\Omega^{-2}h^{-1}\partial_{\underline{u}}\Phi.$ (2.13)
The transport equations for $m$ (2.6a)–(2.6b) help to deduce the subsequent equations for $\mu$,
$\partial_{\underline{u}}\mu={}$ $-\Omega^{2}h\frac{\mu}{r}+\Omega^{2}h\frac{\Phi^{2}}{2r}+\left(1-\mu\right)\Omega^{2}h\frac{\alpha^{2}}{r},$ (2.14a)
$\partial_{u}\mu={}$ $-\underline{h}\frac{\mu}{r}+\underline{h}\frac{\Phi^{2}}{2r}-h\frac{(\partial_{u}w)^{2}}{r}.$ (2.14b)
Due to the Yang–Mills equation (2.5), we have
$\partial_{u}\alpha=-h^{-1}w\frac{Q}{r^{2}}+h^{-1}\left(\frac{\mu}{r}-\frac{Q^{2}}{r^{3}}\right)\alpha.$ (2.15)
Along with (2.15), we need a higher order equation
$\partial_{u}(\partial_{r}\alpha)={}$ $\underbrace{-2w^{2}\frac{h^{-1}}{r^{2}}\alpha+2w\frac{h^{-1}}{r^{2}}\frac{Q}{r}}_{\text{Linear terms}}$
$+2\frac{h^{-1}}{r}\left(\mu-\Phi^{2}\right)\partial_{r}\alpha-\frac{h^{-1}}{r}\Phi\alpha-6w\frac{h^{-1}}{r^{2}}\Phi\alpha^{2}$
$+2\frac{h^{-1}}{r^{2}}\left(2\Phi^{2}-\mu\right)\alpha+2\frac{h^{-1}}{r^{2}}\left(1-\Phi^{2}\right)\alpha^{3}.$ (2.16)
We always denote
$\partial_{r}:=\Omega^{-2}h^{-1}\partial_{\underline{u}}$
the outgoing null vector parametrized by $r$.
3. The $H^{1}$ case
We aim to establish an extension theorem for the solution to the SSEYM system in the class of $\left(r,\,m,\,\alpha\right)\in C^{1}\times C^{1}\times H^{1}$. Recall (1.5) for the definition of the $H^{1}$ norm, where the fixed constant $b$ is given in (1.4).
3.1. Extension theorem in $H^{1}$ space
For any $u_{0}<\underline{u}_{\ast}\in\mathbb{R}$, the domain $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$ bounded by the axis $\Gamma$, $C_{u_{0}}$ and $\underline{C}_{\underline{u}_{\ast}}$, is defined in (1.1).
Theorem 3.1.
Given initial data $\alpha|_{C_{u_{0}}}$, assume the initial energy
$\int_{C_{u_{0}}\cap\{u_{0}\leq\underline{u}\leq\underline{u}_{0}\}}\frac{1}{r^{1+b}}\left(\alpha^{2}+(\alpha^{\prime})^{2}+r^{-2}\hat{w}^{2}\right)\mathrm{d}r,\quad\text{for any}\,\,\,\underline{u}_{0}>u_{0}$
is finite. Suppose we have in the domain $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$ a solution to the SSEYM system $(r,\,m,\,\alpha)\in C^{1}\times C^{1}\times H^{1}$ (if $b>0$) or $(r,\,m,\,\alpha)\in C^{1}\times C^{0}\times H^{1}$ (if $b=0$). There is a constant $\varepsilon>0$ such that if
$\lim_{u\rightarrow\underline{u}_{\ast}}\sup_{\mathcal{D}(u;\,\underline{u}_{\ast})}\mu<\varepsilon^{2},$
then the following conclusions hold.
Case I: $0<b<\frac{\sqrt{89}-9}{4}$. The solution $(r,\,m,\,\alpha)$ extends as a $C^{1}\times C^{1}\times H^{1}$ solution on $\mathcal{D}(u_{0};\,\underline{u}_{1})$, for some $\underline{u}_{1}>\underline{u}_{\ast}$. Moreover, the solution obeys the estimates $\mathcal{D}(u_{0};\,\underline{u}_{1})$
$|r^{-1}\hat{w}|+|\alpha|+|\partial_{u}w|\lesssim{}$ $r^{b},$ (3.1)
$|r^{-1}\hat{w}|+|\alpha|\lesssim{}$ $\varepsilon^{2},$ (3.2)
and the energy bound holds
$\int_{C_{u}\cap\mathcal{D}(u_{0};\,\underline{u}_{1})}\frac{1}{r^{1+b}}\left(\Phi^{2}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}r$
$+\int_{\underline{C}_{\underline{u}}\cap\mathcal{D}(u_{0};\,\underline{u}_{1})}\frac{h^{-1}}{r^{1+b}}\left(\Phi^{2}+\alpha^{2}+h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u<C.$ (3.3)
Case II: $b=0$. There exists $\underline{u}_{1}>\underline{u}_{\ast}$ such that the solution extends to $\mathcal{D}(u_{0};\,\underline{u}_{1})$ in the class of $(r,m,\alpha)\in C^{1}\times C^{0}\times H^{1}$, and the estimates (3.2)–(3.1) hold with $b=0$.
In the rest of this section, we will prove Theorem 3.1.
Let us introduce some conventions for this section. For any $u_{0}<u_{1}\leq\underline{u}_{\ast}$, we recall the definition of the domain $\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast})$ (1.2) that is bounded by $\Gamma$ and the hypersurfaces $C_{u_{0}}$, $C_{u_{1}}$, $\underline{C}_{\underline{u}_{\ast}}$. We make the following simplification
$\mathcal{D}:=\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast}+\zeta),\quad\text{with some constant}\,\,\zeta>0,$ (3.4)
and denote the truncated null cones by
$C_{u}^{[\underline{u}_{0},\,\underline{u}_{1}]}:=C_{u}\cap\{\underline{u}_{0}\leq\underline{u}\leq\underline{u}_{1}\},\quad\underline{C}_{\underline{u}}^{[u_{0},\,u_{1}]}:=\underline{C}_{\underline{u}}\cap\{u_{0}\leq u\leq u_{1}\}.$
Besides, we abbreviate the outgoing cone emanating from the vertex $C_{u}^{[u,\,\underline{u}]}$ by $C_{u}$, and $\underline{C}_{\underline{u}}^{[u_{0},\,u]}$ by $\underline{C}_{\underline{u}}$, whenever the interval of integration is clear in the context.
For the fixed constant $b$ (1.4), we define the energy norms,
$E_{0}(u,\,\underline{u}):={}$ $\int_{C_{u}}\frac{\Omega^{2}h}{r^{1+b}}\left(r^{-2}\hat{w}^{2}+\alpha^{2}\right)\mathrm{d}\underline{u}^{\prime},$
$E_{1}(u,\,\underline{u}):={}$ $E_{0}(u,\,\underline{u})+\int_{C_{u}}\frac{\Omega^{2}h}{r^{1+b}}(\alpha^{\prime})^{2}\mathrm{d}\underline{u}^{\prime},$
$\underline{E}_{0}(u,\,\underline{u}):={}$ $\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(r^{-2}\hat{w}^{2}+h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u^{\prime},$
$\underline{E}_{0}^{\prime}(u,\,\underline{u}):={}$ $\underline{E}_{0}(u,\,\underline{u})+\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\alpha^{2}\mathrm{d}u^{\prime}.$ (3.5)
If we use $r$ as the parameter on $C_{u}$, $\underline{C}_{\underline{u}}$ respectively, then the energies in (3.1) are rewritten as
$E_{0}={}$ $\int_{0}^{r}\frac{r^{-2}\hat{w}^{2}+\alpha^{2}}{r^{1+b}}\mathrm{d}r|_{C_{u}},$
$E_{1}={}$ $\int_{0}^{r}\frac{(\alpha^{\prime})^{2}+\alpha^{2}+r^{-2}\hat{w}^{2}}{r^{1+b}}\mathrm{d}r|_{C_{u}},$
$\underline{E}_{0}={}$ $\int_{r}^{r_{0}}\frac{1}{1-\mu}\frac{r^{-2}\hat{w}^{2}+h^{2}(\partial_{u}w)^{2}}{r^{1+b}}\mathrm{d}r|_{\underline{C}_{\underline{u}}},$
$\underline{E}_{0}^{\prime}={}$ $\underline{E}_{0}+\int_{r}^{r_{0}}\frac{1}{1-\mu}\frac{\alpha^{2}}{r^{1+b}}\mathrm{d}r|_{\underline{C}_{\underline{u}}},$
where $r_{0}:=r(u_{0},\underline{u})$.
3.1.1. Energy estimates theorem
The proof of Theorem 3.1 follows from the energy estimates which we state as below.
Lemma 3.2.
Suppose $(r,\,m,\,\alpha)$ is any solution to the SSEYM system in $\mathcal{D}$ for any $u_{0}<u_{1}\leq\underline{u}_{\ast}$ and some constant $0<\zeta<+\infty$. Given initial data with the bounded energy
$E_{1}(u_{0},\,\underline{u}_{\ast})\leq I_{1},\quad\text{for some}\,I_{1}\in\mathbb{R}_{>0}.$
The subscript $1$ in $I_{1}$ denotes the number of derivatives used in the energy norms. There is an $\varepsilon>0$ such that if
$\mu\leq\varepsilon^{2},\quad\text{in}\,\,\mathcal{D},$
the following statements are concluded.
Case I: $0<b<\frac{\sqrt{89}-9}{4}$. The solution $(r,m,\alpha)\in C^{1}\times C^{1}\times H^{1}$ in $\mathcal{D}$ and it holds that
$|r^{-1}\hat{w}|^{2}\leq C(I_{1})\varepsilon^{2},\quad|\alpha|^{2}\leq C(I_{1})\varepsilon,$ (3.6)
and
$|r^{-1}\hat{w}|^{2}+|\alpha|^{2}+|\partial_{u}w|^{2}\leq C(I_{1})r^{b},$ (3.7)
and the energies are uniformly bounded
$E_{1}(u,\,\underline{u})+\underline{E}_{0}^{\prime}(u,\,\underline{u})\leq C(I_{1}),\quad\text{where}\,\,\,u\in[u_{0},\,u_{1}],\,\,\underline{u}\in[u,\,\underline{u}_{\ast}+\zeta]$ (3.8)
for some constant $C(I_{1})$ depending only on $I_{1}$.
Case II: $b=0$. The solution $(r,m,\alpha)\in C^{1}\times C^{0}\times H^{1}$ in $\mathcal{D}$ and both of (3.6) and (3.8) hold with $b=0$.
Remark 3.3.
The energy arguments leading to Lemma 3.2 are closed in $H^{1}$ space. This fact enables us to prove a local well-posedness theorem with $H^{1}$ data, see Appendix A.
3.1.2. Preliminary estimates
As a preliminary step, we have
$\underline{h}<0,\quad h>0\quad\text{in}\,\,\mathcal{D}.$ (3.9)
The fact $\underline{h}<0$ follows from the monotonicity of $\Omega^{-2}\underline{h}$ in $u$ (see (2.4b)) and the initial requirement $\Omega^{-2}\underline{h}|_{u_{0}}<0$, and moreover, $h>0$ is due to the smallness of $\mu$ which indicates $01-\mu=-\underline{h}h>$ 0.
3.1.3. Continuity argument
We make the following bootstrap assumption. Let $\Lambda$ be a large constant that will be determined (depending on the initial energy). We assume
$|r^{-1}\hat{w}|+|\alpha|\leq\varepsilon\Lambda,\quad\text{in}\,\,\mathcal{D}.$ (3.10)
We will show that the same inequality with $\Lambda$ being replaced by $\frac{\Lambda}{2}$ holds.
Since $r$ is finite in $\mathcal{D}$, it is straightforward to know from (3.10) that
$|\Phi|\lesssim\varepsilon\Lambda,\,\,|w^{2}-1|\lesssim\varepsilon\Lambda,\quad\text{in}\,\,\mathcal{D}.$ (3.11)
3.1.4. Main energy estimates
We will use the bootstrap assumption (3.10) to carry out the energy estimates in details.
Let us start with the first order energy identities. We multiply $2\partial_{\underline{u}}w$ on the Yang–Mills equation (2.5) to obtain the identity
$\partial_{u}(\partial_{\underline{u}}w)^{2}+\frac{\Omega^{2}}{2r^{2}}\partial_{\underline{u}}Q^{2}=0.$ (3.12)
We further multiply $\Omega^{-2}h^{-1}r^{-1-b}$ on (3.12), then
$\partial_{u}\left(r^{-1-b}\Omega^{-2}h^{-1}(\partial_{\underline{u}}w)^{2}\right)-\partial_{u}r^{-1-b}\Omega^{-2}h^{-1}(\partial_{\underline{u}}w)^{2}$
$\qquad-\partial_{u}(\Omega^{-2}h^{-1})\frac{(\partial_{\underline{u}}w)^{2}}{r^{1+b}}+\partial_{\underline{u}}\left(\frac{h^{-1}}{2r^{3+b}}Q^{2}\right)-\partial_{\underline{u}}\left(\frac{h^{-1}}{2r^{3+b}}\right)Q^{2}=0.$
Using the equations (2.4a) and (2.4d), and integrating over the domain $\mathcal{D}$, we arrive at the following energy identity
$\int_{C_{u}}\Omega^{2}h\frac{\alpha^{2}}{r^{1+b}}\mathrm{d}\underline{u}+\int_{\underline{C}_{\underline{u}}}h^{-1}\frac{\Phi^{2}}{2r^{1+b}}\mathrm{d}u$
$+\iint_{\mathcal{D}}\left(\frac{3+b}{2}\frac{\Phi^{2}}{r^{2+b}}-(1+b)\frac{\alpha^{2}}{r^{2+b}}+b\mu\frac{\alpha^{2}}{r^{2+b}}\right)\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
[…]
Note that, among the spacetime integrals on the left hand side of (3.1.4), there is a negative term, which prevents us from deducing any energy bound at this stage.
On the other hand, multiplying $2\partial_{u}w$ on the Yang–Mills equation (2.5) yields
$\partial_{\underline{u}}(\partial_{u}w)^{2}+\frac{\Omega^{2}}{2r^{2}}\partial_{u}Q^{2}=0.$ (3.14)
Multiplied with $hr^{-1-b}$, (3.14) is turned into the following identity with the assistance of (2.4a) and (2.4d),
$\partial_{\underline{u}}\left(h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\right)+(1+b)\Omega^{2}h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}+2\Omega^{-2}(\partial_{\underline{u}}w)^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}$
$\quad+\partial_{u}\left(\frac{\Omega^{2}h}{2}\frac{Q^{2}}{r^{3+b}}\right)+\frac{3+b}{2}\Omega^{2}\underline{h}h\frac{Q^{2}}{r^{4+b}}+\frac{\Omega^{2}}{2}\left(\frac{\mu}{r}-\frac{Q^{2}}{r^{3}}\right)\frac{Q^{2}}{r^{3+b}}=0.$
We then obtain an energy identity, after integrating over the domain $\mathcal{D}$,
$\int_{\underline{C}_{\underline{u}}}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\mathrm{d}u+\int_{C_{u}}\frac{\Omega^{2}h}{2}\frac{\Phi^{2}}{r^{1+b}}\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}\left((1+b)+2\alpha^{2}\right)\Omega^{2}h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}\left(-\frac{3+b}{2}+\left(2+\frac{b}{2}\right)\mu\right)\Omega^{2}\frac{\Phi^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$
[…]
Analogous to (3.1.4), the above identity has a nonpositive spacetime integral on the left hand side and therefore fails to conclude any energy bound. What is more, cooperating the energy identity (3.1.4) with (3.1.4) does not give any bound for the first order energy neither.
In what follows, we proceed to investigate the second order energy identities, which we expect to cancel those terms with unfavourable signs in (3.1.4) and (3.1.4).
By virtue of (2.3), there is
$\partial_{u}\alpha^{\prime}={}$ $\frac{\underline{h}}{r}\alpha^{\prime}-2w^{2}\frac{h^{-1}}{r}\alpha+2w\frac{h^{-1}}{r}\Phi$
$+2\frac{h^{-1}}{r}\left(\mu-\Phi^{2}\right)\alpha^{\prime}-h^{-1}\Phi\alpha-6w\frac{h^{-1}}{r}\Phi\alpha^{2}$
$+2\frac{h^{-1}}{r}\left(2\Phi^{2}-\mu\right)\alpha+2\frac{h^{-1}}{r}\left(1-\Phi^{2}\right)\alpha^{3},$ (3.16)
where we denote the linear terms on the first line as follows
$\ell_{1}=\frac{\underline{h}}{r}\alpha^{\prime},\quad\ell_{2}=-2w^{2}\frac{h^{-1}}{r}\alpha,\quad\ell_{3}=2w\frac{h^{-1}}{r}\frac{Q}{r}.$
Before going into the lengthy calculations for the second order energy estimates, here we sketch the main idea of proof. To obtain an energy identity from (3.1.4), we will multiply on this equation
$r^{-1-b}\cdot\Omega^{2}h\alpha^{\prime}=r^{-b}\partial_{\underline{u}}\alpha$
and integrate over $\iint_{\mathcal{D}}\mathrm{d}u\mathrm{d}\underline{u}$. Regarding to the first terms on the right hand side of the first line of (3.1.4),
$\iint_{\mathcal{D}}\Omega^{2}h\frac{\alpha^{\prime}}{r^{1+b}}\cdot\ell_{1}\mathrm{d}u\mathrm{d}\underline{u}=-\iint_{\mathcal{D}}\left(1-\mu\right)\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$ (3.17)
[…]
$\iint_{\mathcal{D}}\Omega^{2}h\frac{\alpha^{\prime}}{r^{1+b}}\cdot\ell_{i}\,\mathrm{d}u\mathrm{d}\underline{u}$
$\lesssim$ $\left(\iint_{\mathcal{D}}\Omega^{2}\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}\right)^{\frac{1}{2}}\cdot\left(\iint_{\mathcal{D}}\Omega^{2}\left(\frac{\alpha^{2}}{r^{2+b}}+\frac{\Phi^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}\right)^{\frac{1}{2}}$
is not allowed, since we have no yet any improved bounds for the energies on the right hand side of the inequality above. The difficulty coming from these quadratic terms is resolved via applying integration by parts. In this way, we expect to eventually gain spacetime integrals, such as $\iint_{\mathcal{D}}\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$, $\iint_{\mathcal{D}}\Omega^{2}\frac{\Phi^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$, with favourable signs.
Among all the terms on the right side of (3.1.4), the nonlinear terms in the second and third lines can be in principle estimated by means of (after multiplying $r^{-1-b}\cdot\Omega^{2}h\alpha^{\prime}$)
$\lesssim\varepsilon\Lambda\iint_{\mathcal{D}}\left(\frac{\Phi^{2}}{r^{2+b}}+\frac{\alpha^{2}}{r^{2+b}}+\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\right)\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}.$ (3.18)
The extra smallness of (3.18) suggests they could be absorbed, since in the preceding steps of treating the linear terms $\ell_{i}$, we have managed to extract a positive bulk terms as $\iint_{\mathcal{D}}\left(\frac{\Phi^{2}}{r^{2+b}}+\frac{\alpha^{2}}{r^{2+b}}+\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\right)\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$ on the left hand side of the energy inequality through integration by parts.
We now go on with the detailed proof. First of all, let us turn to the left hand side of (3.1.4). After multiplied with $\Omega^{2}h\frac{\alpha^{\prime}}{r^{1+b}}$, it becomes,
$\partial_{u}\alpha^{\prime}\cdot\Omega^{2}h\frac{\alpha^{\prime}}{r^{1+b}}$
[…]
We notice that the last term
$\frac{1+b}{2}\underline{h}h\Omega^{2}\frac{(\alpha^{\prime})^{2}}{r^{2+b}}=-\frac{1+b}{2}(1-\mu)\Omega^{2}\frac{(\alpha^{\prime})^{2}}{r^{2+b}}$
which indeed lies on the left hand side the energy identity has a bad sign. Nevertheless, this term is compensated by (3.17) and their summation still contributes a positive term
$\iint_{\mathcal{D}}\frac{1-b}{2}\left(1-\mu\right)\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u},\quad 0\leq b<\frac{\sqrt{89}-9}{4},$
on the left side of the energy identity.
[…]
$r^{-1-b}\Omega^{2}h\alpha^{\prime}\cdot\ell_{2}=-w^{2}h^{-1}r^{-1-b}\partial_{\underline{u}}\alpha^{2}$
[…]
$+w^{2}r^{-1-b}\alpha^{2}\partial_{\underline{u}}h^{-1}+w^{2}\partial_{\underline{u}}r^{-1-b}\alpha^{2}h^{-1}$
[…]
$+2w^{2}\Omega^{2}r^{-2-b}\alpha^{4}-(1+b)w^{2}\Omega^{2}r^{-2-b}\alpha^{2},$
where we have made use of the equation (2.4a) in the last identity. In particular, we notice that the last term $-(1+b)w^{2}\Omega^{2}r^{-2-b}\alpha^{2}$ which lies on the right hand side the energy identity admits a favourable sign. When $i=3$, we have
$r^{-1-b}\Omega^{2}h\alpha^{\prime}\cdot\ell_{3}=2h^{-1}r^{-2-b}(w^{2}-1)w\cdot\partial_{\underline{u}}\alpha$
[…]
$-\partial_{\underline{u}}w\cdot 2r^{-2-b}h^{-1}(w^{2}-1)\alpha$
[…]
$-4h^{-2}\Omega^{-2}\frac{(\partial_{\underline{u}}w)^{2}}{r^{3+b}}(w^{2}-1)w\alpha-4\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}-6(w^{2}-1)\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}.$
Here the quadratic term $-4\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}$ admits a favourable sign, while the sign of $(4+2b)\Omega^{2}\frac{w^{2}-1}{r^{3+b}}w\alpha$ is not definite. Nevertheless, we can perform the same trick on this quadratic term as follows
$(4+2b)\Omega^{2}\frac{w^{2}-1}{r^{3+b}}w\alpha=(2+b)h^{-1}\frac{w^{2}-1}{r^{3+b}}\partial_{\underline{u}}(w^{2}-1)$
[…]
$+(3+b)\left(1+\frac{b}{2}\right)\Omega^{2}\frac{(w^{2}-1)^{2}}{r^{4+b}}.$
We then combine these calculations to obtain
$r^{-1-b}\Omega^{2}h\alpha^{\prime}\cdot\ell_{3}={}$ $\partial_{\underline{u}}\left(2wh^{-1}\frac{\Phi\alpha}{r^{1+b}}\right)+\left(1+\frac{b}{2}\right)\partial_{\underline{u}}\left(h^{-1}\frac{\Phi^{2}}{r^{1+b}}\right)$
$-4\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}+(3+b)\left(1+\frac{b}{2}\right)\Omega^{2}\frac{\Phi^{2}}{r^{2+b}}$
$-2\left(1+\frac{b}{2}\right)\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}\Phi^{2}-4\Omega^{2}\frac{\Phi}{r^{2+b}}w\alpha^{3}-6\Omega^{2}\Phi\frac{\alpha^{2}}{r^{1+b}},$
where in the second line, the quadratic term $-4\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}$ contributes a positive term in the energy identity, while $(3+b)\left(1+\frac{b}{2}\right)\Omega^{2}\frac{\Phi^{2}}{r^{2+b}}$ offers a bad sign. Nevertheless, noting the positive integral $\iint_{\mathcal{D}}\frac{\Phi^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$ on the left hand side of the first order energy identity (3.1.4), the bad-sign term $(3+b)\left(1+\frac{b}{2}\right)\Omega^{2}\frac{\Phi^{2}}{r^{2+b}}$ will be absorbed when (3.1.4) is added into the second order energy identity.
Taking all the calculations into accounts, we integrate these identities over $\iint_{\mathcal{D}}\mathrm{d}u\mathrm{d}\underline{u}$ to derive the energy identity
$\int_{C_{u}}\frac{(\alpha^{\prime})^{2}}{2r^{1+b}}\Omega^{2}h\mathrm{d}\underline{u}+\int_{\underline{C}_{\underline{u}}}\frac{1}{r^{1+b}}\left(\left(w\alpha-\Phi\right)^{2}-\left(2+\frac{b}{2}\right)\Phi^{2}\right)h^{-1}\mathrm{d}u$
$+\iint_{\mathcal{D}}\left(\frac{1-b}{2}+\frac{3\Phi^{2}}{2}\right)\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}\left((5+b)+(2+b)\Phi^{2}\right)\frac{\alpha^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
$-\iint_{\mathcal{D}}(3+b)\left(1+\frac{b}{2}\right)\frac{\Phi^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
[…]
$+\iint_{\mathcal{D}}\left(2(1-\Phi^{2})\alpha^{2}+4\Phi^{2}-2\mu-6w\Phi\alpha-r\Phi\right)\frac{\alpha\alpha^{\prime}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}\left(2\left(wr\alpha+w^{2}\alpha^{2}\right)-4w\Phi\alpha-(7+b)r\Phi\right)\frac{\alpha^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}.$ (3.19)
Now in view of the identities (3.1.4), (3.1.4) and (3.1.4), to eliminate the negative term $-\iint_{\mathcal{D}}(3+b)\left(1+\frac{b}{2}\right)\Omega^{2}\frac{\Phi^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$ on the left hand side of (3.1.4), and obtain coercive energy on $\underline{C}_{\underline{u}}$, we consider the linear combination
[…]
which leads to
$\int_{C_{u}}\frac{\Omega^{2}h}{r^{1+b}}\left((9+2b)\alpha^{2}+(\alpha^{\prime})^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}$
$+\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(2\left(w\alpha-\Phi\right)^{2}+\frac{\Phi^{2}}{2}+2h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u$
$+\iint_{\mathcal{D}}\left((1-b)+3\Phi^{2}\right)\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}\left((1-9b-2b^{2})+2(2+b)\Phi^{2}+(9b+2b^{2})\mu\right)\frac{\alpha^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}\left(\frac{3(3+b)}{2}+(4+b)\mu\right)\frac{\Phi^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}\left(2(1+b)+4\alpha^{2}\right)\frac{h^{2}(\partial_{u}w)^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
[…]
$+\iint_{\mathcal{D}}\left(\mu+r|\Phi|+\Phi^{2}+\alpha^{2}\right)\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}\left(\mu+r|\Phi|+\Phi^{2}+\alpha^{2}+r|\alpha|\right)\frac{\alpha^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}.$ (3.20)
We note that, since $0\leq b<\frac{\sqrt{89}-9}{4}$ and hence $01-9b-2b^{2}>$ 0, terms on the left hand side of (3.1.4) are all positive. In particular, the energy on $\underline{C}_{\underline{u}}$ is coercive (the energy on $C_{u}$ is obviously coercive), and there is
$\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(\Phi^{2}+\alpha^{2}+h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u$
$\lesssim$ $\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(2\left(w\alpha-\Phi\right)^{2}+\frac{\Phi^{2}}{2}+2h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u.$
In view of the bootstrap assumption (3.10) and the fact that $\mu$ is small, and $r$ is finite in $\mathcal{D}$, the spacetime integrals on the right hand side of (3.1.4) can be absorbed by the bulk integrals on the left hand side. We thus achieve the energy bound
$\int_{C_{u}}\frac{\Omega^{2}h}{r^{1+b}}\left(\Phi^{2}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}\underline{u}$
$+\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(\Phi^{2}+\alpha^{2}+h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u\lesssim I_{1},$ (3.21)
and then (3.8) follows.
3.1.5. Pointwise estimates
With the help of the energy estimates obtained in the previous step, we will deduce various $L^{\infty}$ estimates to improve the bootstrap assumption.
Lemma 3.4.
We have the enhanced estimates for $\mu$
$\mu\lesssim r^{b}\quad\text{and}\quad\lim_{u\rightarrow\underline{u}_{\ast}}\sup_{\mathcal{D}(u,\,\underline{u}_{\ast})}\frac{\mu}{r^{b}}=0.$
In particular, since $b\geq 0$, there is
$\mu|_{\Gamma}=0.$
Proof.
Integrating (2.6a) in $\underline{u}$, we have the mass formula,
$m=\int_{0}^{r}\left(\left(1-\mu\right)\alpha^{2}+\Phi^{2}\right)\mathrm{d}r^{\prime}|_{C_{u}}.$
Due to the smallness of $\mu$, there is the equivalence
$m\sim\int_{0}^{r}\left(\alpha^{2}+\Phi^{2}\right)\mathrm{d}r^{\prime}|_{C_{u}}.$ (3.22)
Since $0\leq r^{\prime}\leq r$, and hence $\frac{1}{r}\leq\frac{1}{r^{\prime}}$, the energy bound (3.1.4) leads to
$\frac{m}{r^{1+b}}\leq\int_{0}^{r}\frac{\left(\left(1-\mu\right)\alpha^{2}+\Phi^{2}\right)}{r^{\prime 1+b}}\mathrm{d}r^{\prime}|_{C_{u}}\leq E_{0}\lesssim I_{0}.$
This lemma is concluded.
As a sequel, we will improve the pointwise estimates for $\alpha$ and $r^{-1}\hat{w}$.
Lemma 3.5.
The estimates for $\alpha$ and $r^{-1}\hat{w}$ are enhanced as follows
$\alpha^{2}\lesssim{}$ $r^{b}\quad\text{and}\quad\alpha^{2}\lesssim\varepsilon,$
$r^{-2}\hat{w}^{2}\lesssim{}$ $r^{b}\quad\text{and}\quad r^{-2}\hat{w}^{2}\lesssim\varepsilon^{2}.$
Remark 3.6.
If $b>0$, Lemma 3.5 suggests
$\alpha|_{\Gamma}=0\quad\text{and}\quad r^{-1}\hat{w}|_{\Gamma}=0.$
Proof of Lemma 3.5.
Multiplying with $2\alpha$ on (2.15) yields
$\partial_{u}\alpha^{2}=-2\frac{h^{-1}}{r}w\Phi\alpha+2\frac{h^{-1}}{r}\left(\mu-\Phi^{2}\right)\alpha^{2}.$
Then integrate in $u$ on each $\underline{C}_{\underline{u}}$,
$\big{|}\alpha^{2}-\alpha^{2}|_{C_{u_{0}}}\big{|}\lesssim{}$ $\int_{\underline{C}_{\underline{u}}}h^{-1}\frac{\Phi^{2}}{r}\mathrm{d}u+\int_{\underline{C}_{\underline{u}}}h^{-1}w^{2}\frac{\alpha^{2}}{r}\mathrm{d}u$
$+\int_{\underline{C}_{\underline{u}}}2h^{-1}\left(\mu+\Phi^{2}\right)\frac{\alpha^{2}}{r}\mathrm{d}u$
[…]
where the bootstrap assumption (3.10) and the energy bound (3.1.4) are used. That is, we have derived a preliminary estimate for $\alpha$,
$|\alpha|^{2}\lesssim\alpha^{2}|_{C_{u_{0}}}+I_{1}.$
Based on the boundedness of $\alpha$, we can further refine its estimate. Noting that the boundedness of $\alpha$ tells $r\alpha^{2}|_{\Gamma}=0$, we integrate $\partial_{r}(r\alpha^{2})$ in $r$ on each $C_{u}$ to yield
$r\alpha^{2}={}$ $\int_{0}^{r}\left(\alpha^{2}+2\alpha\alpha^{\prime}\right)\mathrm{d}r^{\prime}|_{C_{u}}$
[…]
Here in the last inequality, we have access to the equivalence (3.22) and the energy bound (3.1.4). As a result,
$\alpha^{2}\lesssim\mu+\mu^{\frac{1}{2}}r^{\frac{b}{2}}.$
Combined with the smallness of $\mu$ and Lemma 3.4, the first part of Lemma 3.5 is justified.
The estimate for $r^{-1}\hat{w}$ ( or $\Phi$) follows in an analogous fashion. We will first show that $r^{-1}\hat{w}$ is bounded by means of integrating $\partial_{u}\left(r^{-2}\hat{w}^{2}\right)$ in $u$ on each $\underline{C}_{\underline{u}}$,
$\frac{\hat{w}^{2}}{r^{2}}\Big{|}_{u_{0}}^{u}={}$ $\int_{\underline{C}_{\underline{u}}}\left(\partial_{u}w\frac{2\hat{w}}{r^{2}}-2\underline{h}\frac{\hat{w}^{2}}{r^{3}}\right)\mathrm{d}u$
[…]
Therefore, $r^{-2}\hat{w}^{2}$ is bounded as
$r^{-2}\hat{w}^{2}\lesssim r^{-2}\hat{w}^{2}\Big{|}_{C_{u_{0}}}+I_{0},$
which entails $r\cdot\left(r^{-2}\hat{w}^{2}\right)|_{\Gamma}=0$. It enables us to integrate $\partial_{r}\left(r\cdot\left(r^{-2}\hat{w}^{2}\right)\right)$ in $r$ on each $C_{u}$ and then an application of the Cauchy-Schwarz inequality leads to
$r\frac{\hat{w}^{2}}{r^{2}}={}$ $\int_{0}^{r}\left(\frac{2\alpha\hat{w}}{r}-\frac{\hat{w}^{2}}{r^{2}}\right)\mathrm{d}r^{\prime}|_{C_{u}}$
[…]
We arrive at
$r^{-2}\hat{w}^{2}\lesssim\mu,$
and thus obtain the concluding estimate for $r^{-1}\hat{w}$ with the help of Lemma 3.4. ∎
With Lemma 3.5 concluded, we have improved the pointwise estimates for $r^{-1}\hat{w}$ and $\alpha$ and hence close the bootstrap argument. In the subsequent analysis, we will derive more precise estimates to provide a detailed description for the solution.
3.1.6. Estimates for the geometry
Lemma 3.7.
Given the assumptions in Lemma 3.2, the following estimates hold in $\mathcal{D}$,
$|\ln(\Omega^{2}h)|+|\ln(-\underline{h})|+|\ln h|\lesssim I_{0}.$
Proof.
Before going into the pointwise estimates, we will deduce some integral bounds for $\mu$.
Denote $r_{0}=r(u_{0},\,\underline{u})$. Based on (2.14b), an integration in $u$ along each $\underline{C}_{\underline{u}}$ yields
$\mu|_{r}^{r_{0}}+\int_{r}^{r_{0}}\frac{\mu}{r^{\prime}}\mathrm{d}r^{\prime}|_{\underline{C}_{\underline{u}}}$
[…]
where the energy bound (3.8) is employed in the last inequality. Since $\mu\leq\varepsilon^{2}$, we thus obtain
$\int_{\underline{C}_{\underline{u}}}(1-\mu)\frac{\mu}{r}h^{-1}\mathrm{d}u=\int_{r}^{r_{0}}\frac{\mu}{r^{\prime}}\mathrm{d}r^{\prime}\big{|}_{\underline{C}_{\underline{u}}}\lesssim I_{0}+\varepsilon^{2}$ (3.24)
along each incoming cone $\underline{C}_{\underline{u}}\cap\mathcal{D}$, $\underline{u}\in(u_{0},\,\underline{u}_{\ast}+\zeta]$.
In the same manner, due to (2.14a), there is, along each $C_{u}\cap\mathcal{D}$, $u\in[u_{0},\,u_{1}]$,
$\mu|_{0}^{r}+\int_{0}^{r}\frac{\mu}{r^{\prime}}\mathrm{d}r^{\prime}|_{C_{u}}$
[…]
which leads to
$\int_{C_{u}}\frac{\mu}{r}\Omega^{2}h\,\mathrm{d}\underline{u}=\int_{0}^{r}\frac{\mu}{r^{\prime}}\mathrm{d}r^{\prime}\big{|}_{C_{u}}\lesssim I_{0}+\varepsilon^{2}$ (3.26)
along each outgoing cone $C_{u}\cap\mathcal{D}$, $u\in[u_{0},\,u_{1}]$.
Now to prove this lemma, let us start with $\ln(\Omega^{2}h)$. We rewrite the equation (2.4d) alternatively as
$\partial_{u}\ln(\Omega^{2}h)=-h^{-1}\left(\frac{\mu}{r}-\frac{\Phi^{2}}{r}\right).$
Using the gauge choice (2.8), the integral bound (3.24) and the energy bound (3.8), we derive via an integration in $u$ that,
$\Big{|}\ln\left(2\Omega^{2}h\right)\Big{|}\lesssim\int^{u}_{u_{0}}\left(\frac{\mu}{r}+\frac{\Phi^{2}}{r}\right)h^{-1}\mathrm{d}u\lesssim I_{0}.$
After the estimate for $\ln(\Omega^{2}h)$, we turn to $\ln(-\underline{h})$. The equation (2.4c) indicates
$\partial_{\underline{u}}\left(\ln(-\underline{h})\right)=\frac{\Omega^{2}h}{1-\mu}\left(\frac{\mu}{r}-\frac{\Phi^{2}}{r}\right)$
which together with the energy bound (3.8) and (3.26) gives rise to
$|\ln(-\underline{h})|_{\Gamma}-\ln(-\underline{h})|\lesssim\int_{0}^{r}\left(\frac{\mu}{r^{\prime}}+\frac{\Phi^{2}}{r^{\prime}}\right)\mathrm{d}r^{\prime}|_{C_{u}}\lesssim I_{0}+\varepsilon^{2}.$
Noting the value of $\ln(-\underline{h})|_{\Gamma}$ (2.9), we then obtain the bound for $|\ln(-\underline{h})|$.
In the end, for $\ln h$, the equation (2.4a) suggests
$\partial_{\underline{u}}\ln h=-2\Omega^{2}h\frac{\alpha^{2}}{r},$
and as a consequence of the energy bound (3.8),
$|\ln h|_{\Gamma}-\ln h|\lesssim\int_{0}^{r}\frac{\alpha^{2}}{r^{\prime}}\mathrm{d}r^{\prime}|_{C_{u}}\lesssim I_{0}^{2}.$
In view of Lemma 3.4, $01+h\underline{h}|_{\Gamma}=\mu|_{\Gamma}=$ 0, and hence $h|_{\Gamma}=-\underline{h}^{-1}|_{\Gamma}=(\Omega^{2}h)^{-1}|_{\Gamma}$, the estimate for $\ln h$ is proved.
In particular, if $b>0$, we can further show the solution $w\in C^{1}$.
Lemma 3.8.
In the case of $0<b<\frac{\sqrt{89}-9}{4}$, we have the following estimate for $\partial_{u}w$,
$|\partial_{u}w|^{2}\lesssim r^{b}.$
Proof.
The fact $u=\underline{u}$ on $\Gamma$ and the boundary condition $w|_{\Gamma}=\pm 1$, plus Remark 3.6 entail
$\partial_{u}w|_{\Gamma}=-\partial_{\underline{u}}w|_{\Gamma}=0.$
Now we can integrate the YM equation (2.5) from the axis and use the $L^{\infty}$ estimate for $\Phi$ (Lemma 3.5) to conclude
$|\partial_{u}w|\lesssim{}$ $|\partial_{u}w|_{\Gamma}+\int_{0}^{r}h^{-1}|w|\frac{|\Phi|}{r^{\prime}}\mathrm{d}r^{\prime}|_{C_{u}}$
[…]
3.1.7. Proof of Theorem 3.1
For the SSEYM system, the first singularity must emanate from the axis. One may refer to Theorem B.2 for more details. Therefore, it suffices to focus on a small neighbourhood of the tip $u=\underline{u}=\underline{u}_{\ast}$, and we may assume $u_{0}$ is close to $\underline{u}_{\ast}$ and
$\mu\leq\varepsilon,\quad\text{on}\,\,\mathcal{D}(u_{0};\,\underline{u}_{\ast})$
As suggested by Lemma 3.2, the solution ($\alpha$) extends as an $H^{1}$ solution on the closure of the domain $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$. To complete the proof Theorem 3.1, we have to show that the solution can be extended beyond $\underline{C}_{\underline{u}_{\ast}}$. More explicitly, we shall verify that the solution can be extended to $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\,\underline{u}_{\ast}+\zeta)$ if $\zeta>0$ is small enough.
We assume that
$\mu\leq\frac{\varepsilon_{1}+\varepsilon}{2}\,\,\text{on}\,\,\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast}+\zeta),\quad\text{for any}\,\,u_{1}\in(u_{0},\,\underline{u}_{\ast}],$ (3.27)
where $\varepsilon_{1}>\varepsilon$ is a small constant.
Define
$u_{2}=\sup\,\{u_{1}\in(u_{0},\,\underline{u}_{\ast}]:\,\,\sup_{\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast}+\zeta)}\mu\leq\frac{\varepsilon_{1}+\varepsilon}{2}\}.$
Then either $u_{2}=\underline{u}_{\ast}$ or $u_{2}<\underline{u}_{\ast}$ and $\sup_{\mathcal{D}(u_{0},\,u_{2};\,\underline{u}_{\ast}+\zeta)}\mu=\frac{\varepsilon_{1}+\varepsilon}{2}$. Now we shall exclude the second possibility. This will be proved by showing that the bootstrap assumption (3.27) can be improved as
$\mu<\frac{\varepsilon_{1}+\varepsilon}{2},\quad\text{on}\,\,\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast}+\zeta),\quad\text{for any}\,\,u_{1}\in(u_{0},\,\underline{u}_{\ast}]$
if $\zeta$ is small enough.
Preliminarily, by virtue of the local existence theorem (Appendix A) and Lemma 3.2, the bootstrap assumption (3.27) leads to an $H^{1}$ solution on $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\,\underline{u}_{\ast}+\zeta)$. In particular, the $H^{1}$ norm
$\int_{C_{u}^{[\underline{u}_{\ast},\,\underline{u}_{\ast}+\zeta]}}\frac{\Omega^{2}h}{r}\left(\Phi^{2}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}\underline{u},\quad u\in[u_{0},\,u_{1}],$ (3.28)
is uniformly bounded.
On any $C_{u}$, $u\in[u_{0},\,\underline{u}_{\ast}]$, an integration in $\underline{u}^{\prime}\in[\underline{u}_{\ast},\,\underline{u}]$ as (3.1.6) yields the identity
$\mu\Big{|}_{\underline{u}_{\ast}}^{\underline{u}}+\int_{r(u,\,\underline{u}_{\ast})}^{r(u,\,\underline{u})}\frac{\mu}{r}\mathrm{d}r|_{C_{u}}=\int_{\underline{u}_{\ast}}^{\underline{u}}\frac{\Omega^{2}h}{r}\left(\left(1-\mu\right)\alpha^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}^{\prime},$
where $\underline{u}\in[\underline{u}_{\ast},\,\underline{u}_{\ast}+\zeta]$. Then $\mu$ is bounded as
$\mu(u,\,\underline{u})<\mu(u,\,\underline{u}_{\ast})+\int_{\underline{u}_{\ast}}^{\underline{u}}\frac{\Omega^{2}h}{r}\left(\alpha^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}^{\prime},\quad\underline{u}\in[\underline{u}_{\ast},\,\underline{u}_{\ast}+\zeta].$
Due to the uniform boundedness of the $H^{1}$ norm (3.28), we can take $\zeta$ small enough, such that
$\int_{\underline{u}_{\ast}}^{\underline{u}}\frac{\Omega^{2}h}{r}\left(\alpha^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}^{\prime}<\frac{\varepsilon_{1}-\varepsilon}{2},\quad\forall\,\underline{u}\in[\underline{u}_{\ast},\,\underline{u}_{\ast}+\zeta].$
Thus, making use of the bootstrap assumption (3.27) improves the upper bound for $\mu$
$\mu(u,\,\underline{u})<\mu(u,\,\underline{u}_{\ast})+\frac{\varepsilon_{1}-\varepsilon}{2}<\frac{\varepsilon_{1}+\varepsilon}{2}.$
In summary, we have proved (3.27) holds with $u_{1}=\underline{u}_{\ast}$. As a result, by the proof leading to Lemma 3.2, we know that the solution extends as an $H^{1}$ solution on $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\,\underline{u}_{\ast}+\zeta)$. Combined with the local existence theorem (Appendix A) for $H^{1}$ data, it further allows us to extend the solution to $\mathcal{D}(\underline{u}_{\ast};\,\underline{u}_{\ast}+\zeta^{\prime})$ for some $\zeta^{\prime}\in(0,\,\zeta]$. We complete the proof of Theorem 3.1.
3.2. Cauchy stability in $H^{1}$ space
Having established the extension theorem 3.1, we will justify the Cauchy stability within the same topology. This amounts to proving a small data global existence in the finite domain $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\underline{u}_{\ast})$, where $\underline{u}_{\ast}\in(u_{0},\,+\infty)$. Recalling that the energies $E_{i}$, $i=0,1$, and $\underline{E}_{0}$, $\underline{E}_{0}^{\prime}$, are defined as in (3.1), we now state the Cauchy stability theorem.
Theorem 3.9.
Suppose on the initial hypersurface $C_{u_{0}}$ the initial energy is small
$E_{1}(u_{0},\,\underline{u}_{\ast})<\varepsilon^{2},$
with $\varepsilon$ being small enough.
Case I: $0<b<\frac{\sqrt{89}-9}{4}$. There is a unique solution $\left(r,m,\alpha\right)\in C^{1}\times C^{1}\times H^{1}$ on $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\underline{u}_{\ast})$ such that
$r^{-2}|\hat{w}|^{2}+|\alpha|^{2}+|\partial_{u}w|^{2}+\mu\lesssim\varepsilon^{2}r^{b},$
and for $u_{0}<u\leq\underline{u}_{\ast}$, $u_{0}<\underline{u}\leq\underline{u}_{\ast}$,
$E_{1}(u,\,\underline{u})+\underline{E}^{\prime}_{0}(u,\,\underline{u})\lesssim\varepsilon^{2}.$ (3.29)
Case II: $b=0$. There is a unique solution $\left(r,m,\alpha\right)\in C^{1}\times C^{0}\times H^{1}$ on $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\underline{u}_{\ast})$ such that
$r^{-2}|\hat{w}|^{2}+|\alpha|^{2}+\mu\lesssim\varepsilon^{2},$
and (3.29) holds as well.
Proof.
The proof goes through in line with the one leading to Theorem 3.1 and thus we omit the details.
4. The weighted $H^{1}$ case
Given two constants $u_{0}$ and $\underline{u}_{\ast}$ with $u_{0}<\underline{u}_{\ast}$, for any $\delta>0$, we will consider in the domain $\mathcal{D}(u_{0},\,u;\underline{u}_{\ast}+\delta)$ with $u\in[u_{0},\,\underline{u}_{\ast})$, a class of weighted energies. The weighted norm is defined through a scale invariant function $\xi$ (1.8). Indeed, we will prove later in Proposition 4.4 that in the domain concerned,
$10\leq\xi<$ 1 .
Before the statement of the main theorem, let us introduce some additional notations for this section. We denote the strip
$\mathcal{S}(u,\,\underline{u}):=[u_{0},\,u]\times[\underline{u}_{\ast},\,\underline{u}],\quad u\in(u_{0},\,\underline{u}_{\ast}),\,\,\underline{u}\in(\underline{u}_{\ast},\,\underline{u}_{\ast}+\delta],$ (4.1)
and $\mathcal{S}(u,\,\underline{u})$ will be simplified as $\mathcal{S}$. Besides, we abbreviate $C_{u}^{[\underline{u}_{\ast},\,\underline{u}]}$ by $C_{u}$, and $\underline{C}_{\underline{u}}^{[u_{0},\,u]}$ by $\underline{C}_{\underline{u}}$. We also denote the value of $m$ at $\underline{u}=\underline{u}_{\ast}$ by $m_{\ast}$. Similar conventions apply to $\underline{h}_{\ast},\,\mu_{\ast},\cdots$.
Fix $b$ as in (1.4), and a constant $\gamma\in(b,2]$. We define the $\xi$ weighted energy norms
$\|\alpha\|^{2}_{H^{0}_{w}(C_{u})}:={}$ $\int_{C_{u}}\frac{\Omega^{2}h}{r^{1+b}}\left(\frac{r^{-2}\hat{w}^{2}}{\xi^{3+\gamma}}+\frac{\alpha^{2}}{\xi^{1+\gamma}}\right)\mathrm{d}\underline{u},$ (4.2a)
$\|\hat{w}\|^{2}_{H^{1}_{w}(\underline{C}_{\underline{u}})}:={}$ $\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(\frac{(\partial_{u}w)^{2}}{\xi^{3+\gamma}}+\frac{r^{-2}\hat{w}^{2}}{\xi^{1+\gamma}}\right)\mathrm{d}u,$ (4.2b)
$\|\alpha\|^{2}_{H^{1}_{w}(C_{u})}:={}$ $\|\alpha\|^{2}_{H^{0}_{w}(C_{u})}+\int_{C_{u}}\xi^{1-\gamma}\frac{(\alpha^{\prime})^{2}}{r^{1+b}}\Omega^{2}h\mathrm{d}\underline{u},$ (4.2c)
$\|\hat{w}\|^{2}_{H^{1^{\prime}}_{w}(\underline{C}_{\underline{u}})}:={}$ $\|\hat{w}\|^{2}_{H^{1}_{w}(\underline{C}_{\underline{u}})}+\int_{\underline{C}_{\underline{u}}}\xi^{1-\gamma}\frac{\alpha^{2}}{r^{1+b}}h^{-1}\mathrm{d}u.$ (4.2d)
4.1. Extension theorem in the weighted $H^{1}$ space
Theorem 4.1.
Fix a number $b$ as in (1.4), and a constant $\gamma\in(b,2]$. Given datum $\alpha|_{C_{u_{0}}}$ that is supported in $(\underline{u}_{\ast},\,+\infty)$ (namely, $\alpha|_{C_{u_{0}}\cap\{\underline{u}\leq\underline{u}_{\ast}\}}=0$), we assume the initial energy
$\|\alpha\|^{2}_{H^{1}_{w}(C^{[\underline{u}_{\ast},\,\underline{u}_{\ast}+\delta]}_{u_{0}})}$
is bounded for some $\delta>0$, and $|r-r_{\ast}|^{-\frac{\gamma}{2}}\alpha|_{C_{u_{0}}}$ is continuous at $\underline{u}=\underline{u}_{\ast}$. If $\delta$ is small enough, we might suppose
$\|\alpha\|^{2}_{H^{1}_{w}(C^{[\underline{u}_{\ast},\,\underline{u}_{\ast}+\delta]}_{u_{0}})}\leq\varepsilon^{2},$
where $\varepsilon$ is small enough (depending on $\delta$).
Case I: $b>0$. If $\delta$ is small enough, the solution $(r,m,\alpha)$ extends as a $C^{1}\times C^{1}\times H^{1}_{w}$ solution in the domain of $\mathcal{D}(u_{0},\,u;\,\underline{u}_{\ast}+\delta)$ with $u\in[u_{0},\,\underline{u}_{\ast})$. Moreover, for each $u\in[u_{0},\,\underline{u}_{\ast})$, $\underline{u}\in[\underline{u}_{\ast},\,\underline{u}_{\ast}+\delta]$, the solution admits the energy bound
$\|\hat{w}\|^{2}_{H^{1\prime}_{w}(\underline{C}_{\underline{u}})}+\|\alpha\|^{2}_{H^{1}_{w}(C_{u})}\lesssim\varepsilon^{2}.$
In the strip $\mathcal{S}(u,\,\underline{u}_{\ast}+\delta)$, $u\in[u_{0},\,\underline{u}_{\ast})$, there is the pointwise estimate
$\xi^{-2-\gamma}r^{-2}\hat{w}^{2}+\xi^{-\gamma}\alpha^{2}+\xi^{-3-\gamma}|\partial_{u}w|^{2}+\xi^{-1-\gamma}\mu\lesssim\varepsilon^{2}r^{b}.$
Moreover, if $|(r-r_{\ast})^{1-\frac{\gamma}{2}}\alpha^{\prime}|_{L^{\infty}(C^{[\underline{u}_{\ast},\,\underline{u}_{\ast}+\delta]}_{u_{0}})}$ is bounded, then
$\xi^{2-\gamma}(\alpha^{\prime})^{2}\lesssim r^{b},\quad\text{in}\,\,\mathcal{S}(u,\,\underline{u}_{\ast}+\delta),\,\,u\in[u_{0},\,\underline{u}_{\ast}).$ (4.3)
Case II: $b=0$. If $\delta$ is small enough, the solution $(r,m,\alpha)$ extends as a $C^{1}\times C^{0}\times H^{1}_{w}$ solution in the domain of $\mathcal{D}(u_{0},\,u;\,\underline{u}_{\ast}+\delta)$ with $u\in[u_{0},\,\underline{u}_{\ast})$. The estimates as in Case I, except the $L^{\infty}$ estimate for $\partial_{u}w$, hold as well with $b=0$.
Remark 4.2.
Since the data $\alpha|_{C_{u_{0}}}\equiv 0$ in $\underline{u}\leq\underline{u}_{\ast}$, the domain $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$ is flat and hence
$m=0,\quad\text{on}\,\,\mathcal{D}(u_{0};\,\underline{u}_{\ast}).$
Therefore, it suffices to establish the energy estimates in the strip $\mathcal{S}(u,\,\underline{u}_{\ast}+\delta)$ with $u\in[u_{0},\,\underline{u}_{\ast})$.
In the proof, we have made the following continuity assumptions on the flat boundary cone $\underline{C}_{\underline{u}_{\ast}}$. Since the domain $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$ is flat and hence
$\lim_{\underline{u}\rightarrow\underline{u}_{\ast}^{-}}\int_{\underline{C}_{\underline{u}}}\left(|\xi|^{-1-\gamma}h^{-1}\frac{\Phi^{2}}{r^{1+b}}+|\xi|^{-3-\gamma}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\right)\mathrm{d}u=0,$
$\lim_{\underline{u}\rightarrow\underline{u}_{\ast}^{-}}|\xi|^{-1-\gamma}\frac{\mu}{r^{b}}+\xi^{-2-\gamma}\frac{\Phi^{2}}{r^{b}}+\xi^{-3-\gamma}\frac{(\partial_{u}w)^{2}}{r^{b}}=0.$
We assume that these quantities are continuous at $\underline{u}=\underline{u}_{\ast}$. That is, we assume
$\lim_{\underline{u}\rightarrow\underline{u}_{\ast}^{+}}\int_{\underline{C}_{\underline{u}}}\left(|\xi|^{-1-\gamma}h^{-1}\frac{\Phi^{2}}{r^{1+b}}+|\xi|^{-3-\gamma}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\right)\mathrm{d}u=0,$
$\lim_{\underline{u}\rightarrow\underline{u}_{\ast}^{+}}|\xi|^{-1-\gamma}\frac{\mu}{r^{b}}+\xi^{-2-\gamma}\frac{\Phi^{2}}{r^{b}}+\xi^{-3-\gamma}\frac{(\partial_{u}w)^{2}}{r^{b}}=0.$
Remark 4.3.
The estimate (4.3) indicates the singularity emanating from the initial sphere $S_{\underline{u}_{\ast},u_{0}}$ (with singular $\alpha^{\prime}$) is propagated along the incoming cone $\underline{C}_{\underline{u}_{\ast}}$.
We intend to carry out the proof of Theorem 4.1 in the remainder of this section.
4.1.1. Bootstrap assumptions
Let $\Lambda$ be a large constant that will be determined (depending on the initial energy). We suppose in $\mathcal{S}$ there are
$|r^{-1}\hat{w}|^{2}+|\alpha|^{2}+\mu\leq{}$ $\varepsilon^{2}\Lambda^{2},$ (4.4a)
$|rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}|\leq{}$ $\varepsilon^{2}\Lambda^{2}.$ (4.4b)
It follows from the bootstrap assumption for $\hat{w}$ (4.4a) that
$w^{2}\lesssim 1+\varepsilon\Lambda,\quad\text{in}\,\,\mathcal{S}.$ (4.5)
Using (2.4b) and the bootstrap assumption for $\mu$, we have, as stated in (3.9), that
$\underline{h}<0,\quad h>0,\quad\text{in}\,\,\mathcal{S}.$
In regard to the weight $\xi$, we list some basic facts followed from the bootstrap assumptions.
Proposition 4.4.
We have the identities concerning the derivatives of $\xi$
$r\partial_{\underline{u}}\xi=(1-\xi)\Omega^{2}h,$
$rh\frac{\partial_{u}\xi}{\xi}+\underline{h}h=rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}.$
As a result, under the bootstrap assumption (4.4a), there is
$0\leq\xi<1,\quad\partial_{u}\xi\geq 0,\quad\partial_{\underline{u}}\xi>0,\quad\text{in}\,\,\mathcal{S}.$
Proof.
For any $u\in[u_{0},\,\underline{u}_{\ast})$, $\underline{u}\in[\underline{u}_{\ast},\,\underline{u}_{\ast}+\delta]$, there is always, in the strip $\mathcal{S}$, $r\geq r_{\ast}>0$ since $h>0$, and thus
$10\leq\xi<$ 1 .
Moreover, it follows that
$\partial_{\underline{u}}\xi=\frac{\Omega^{2}hr_{\ast}}{r^{2}}=(1-\xi)\frac{\Omega^{2}h}{r}>0.$
A computation yields $\partial_{u}\xi=\frac{\underline{h}-\underline{h}_{\ast}}{r}-\frac{\underline{h}\xi}{r}$, and hence
$\frac{\partial_{u}\xi}{\xi}=\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}-\frac{\underline{h}}{r}.$
Alternatively, we have the identity
$\partial_{u}\xi=\frac{-\underline{h}_{\ast}}{r}+\frac{\underline{h}r_{\ast}}{r^{2}}=-\frac{r_{\ast}}{r}\left(\frac{\underline{h}_{\ast}}{r_{\ast}}-\frac{\underline{h}}{r}\right).$
It is suggested from this formulation that $\partial_{u}\xi$ will be nonnegative if we can show that $\frac{\underline{h}}{r}$ is increasing in $\underline{u}$. In fact, using the equation (2.4c) and the bootstrap assumption (4.4a),
$\partial_{\underline{u}}\frac{\underline{h}}{r}=\frac{\partial_{\underline{u}}\underline{h}}{r}-\frac{\Omega^{2}h\underline{h}}{r^{2}}=-\left(\mu-\Phi^{2}\right)\frac{\Omega^{2}}{r^{2}}+(1-\mu)\frac{\Omega^{2}}{r^{2}}>0,$
and we finish the proof. ∎
4.1.2. Weighted energy estimates
To begin with, we multiply $\Omega^{-2}h^{-1}r^{-1-b}\xi^{-1-\gamma}$ on (3.12) and follow the procedures leading to (3.1.4). Consequently, we obtain an identity in the strip $\mathcal{S}$ for any $u\in[u_{0},\,\underline{u}_{\ast})$, $\underline{u}\in[\underline{u}_{\ast},\,\underline{u}_{\ast}+\delta]$,
$\int_{C_{u}}\xi^{-1-\gamma}\Omega^{2}h\frac{\alpha^{2}}{r^{1+b}}\mathrm{d}\underline{u}+\int_{\underline{C}_{\underline{u}}}\xi^{-1-\gamma}h^{-1}\frac{\Phi^{2}}{2r^{1+b}}\mathrm{d}u$
$+\iint_{\mathcal{S}}\xi^{-1-\gamma}\Omega^{2}\left(\frac{3+b}{2}\frac{\Phi^{2}}{r^{2+b}}-(1+b)\frac{\alpha^{2}}{r^{2+b}}+b\mu\frac{\alpha^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}$
[…]
$+\iint_{\mathcal{S}}\left(\partial_{u}\xi^{-1-\gamma}\Omega^{2}h\frac{\alpha^{2}}{r^{1+b}}+\partial_{\underline{u}}\xi^{-1-\gamma}h^{-1}\frac{\Phi^{2}}{2r^{1+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}.$
Note that, the energy on the initial cone $\underline{C}_{\underline{u}_{\ast}}$, $\int_{\underline{C}_{\underline{u}_{\ast}}}\xi^{-1-\gamma}h^{-1}\frac{\Phi^{2}}{2r^{1+b}}\mathrm{d}u$, vanishes due to the continuity assumption on $\underline{C}_{\underline{u}_{\ast}}$, as shown in Remark 4.2.
With the help of Proposition 4.4, the above energy identity is further calculated as
$\int_{C_{u}}\xi^{-1-\gamma}\Omega^{2}h\frac{\alpha^{2}}{r^{1+b}}\mathrm{d}\underline{u}+\int_{\underline{C}_{\underline{u}}}\xi^{-1-\gamma}h^{-1}\frac{\Phi^{2}}{2r^{1+b}}\mathrm{d}u$
$+\iint_{\mathcal{S}}\xi^{-1-\gamma}\Omega^{2}\left(\frac{2+b-\gamma}{2}\frac{\Phi^{2}}{r^{2+b}}+(\gamma-b)\frac{\alpha^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{S}}(1+\gamma)\Omega^{2}\xi^{-2-\gamma}\frac{\Phi^{2}}{2r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$
[…]
$-\iint_{\mathcal{S}}(1+\gamma)\xi^{-1-\gamma}\Omega^{2}\left(rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\right)\frac{\alpha^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}.$
Since $\gamma\in(b,2]$, terms on the left hand side are all positive. At the same time, due to the bootstrap assumptions (4.4a)–(4.4b), the spacetime integrals on the right hand side have extra smallness and hence can be absorbed by the positive term $\iint_{\mathcal{S}}(\gamma-b)\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$ on the left side. Therefore, we conclude
$\int_{C_{u}}\xi^{-1-\gamma}\Omega^{2}h\frac{\alpha^{2}}{r^{1+b}}\mathrm{d}\underline{u}+\int_{\underline{C}_{\underline{u}}}\xi^{-1-\gamma}h^{-1}\frac{\Phi^{2}}{2r^{1+b}}\mathrm{d}u$
$+\iint_{\mathcal{S}}\xi^{-1-\gamma}\Omega^{2}\left(\frac{2+b-\gamma}{2}\frac{\Phi^{2}}{r^{2+b}}+\frac{\gamma-b}{2}\frac{\alpha^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{S}}(1+\gamma)\xi^{-2-\gamma}\Omega^{2}\frac{\Phi^{2}}{2r^{2+b}}\lesssim\int_{C_{u_{0}}}\xi^{-1-\gamma}\Omega^{2}h\frac{\alpha^{2}}{r^{1+b}}\mathrm{d}\underline{u}.$ (4.6)
Meanwhile, multiplying $\xi^{-3-\gamma}hr^{-1-b}$ on (3.14) and following the proof leading to (3.1.4), we derive another energy identity
$\int_{\underline{C}_{\underline{u}}}\xi^{-3-\gamma}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\mathrm{d}u+\int_{C_{u}}\xi^{-3-\gamma}\Omega^{2}h\frac{\Phi^{2}}{2r^{1+b}}\mathrm{d}\underline{u}$
$+\iint_{\mathcal{S}}\xi^{-3-\gamma}\Omega^{2}\left((1+b)h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}-\frac{3+b}{2}\frac{\Phi^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{S}}\xi^{-3-\gamma}\Omega^{2}\left(2\alpha^{2}h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}+\left(2+\frac{b}{2}\right)\mu\frac{\Phi^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}$
[…]
$+\iint_{\mathcal{S}}\left(\partial_{\underline{u}}\xi^{-3-\gamma}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}+\partial_{u}\xi^{-3-\gamma}\Omega^{2}h\frac{\Phi^{2}}{2r^{1+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}.$
Note that, we have used the fact that the energy $\int_{\underline{C}_{\underline{u}_{\ast}}}\xi^{-3-\gamma}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\mathrm{d}u$ vanishes, in view of Remark 4.2. Then substituting the calculations in Proposition 4.4 into the last line of the above identity, we have
$\int_{\underline{C}_{\underline{u}}}\xi^{-3-\gamma}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\mathrm{d}u+\int_{C_{u}}\xi^{-3-\gamma}\Omega^{2}h\frac{\Phi^{2}}{2r^{1+b}}\mathrm{d}\underline{u}$
$+\iint_{\mathcal{S}}\xi^{-3-\gamma}\Omega^{2}\left((1+b)h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}+\frac{\gamma-b}{2}\frac{\Phi^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{S}}(3+\gamma)\xi^{-4-\gamma}\left(1-\xi\right)\Omega^{2}h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{S}}2\xi^{-3-\gamma}\Omega^{2}\alpha^{2}h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$
[…]
$+\iint_{\mathcal{S}}(3+\gamma)\xi^{-3-\gamma}\Omega^{2}\left(\mu+\Phi^{2}+\big{|}rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\big{|}\right)\frac{\Phi^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}.$
By the bootstrap assumptions (4.4a)–(4.4b), the spacetime integrals on the right hand side are in fact small error terms and can be absorbed by $\iint_{\mathcal{D}}\frac{\gamma-b}{2}\Omega^{2}\frac{\Phi^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$ on the left side. Consequently, we achieve the following energy bound
$\int_{\underline{C}_{\underline{u}}}\xi^{-3-\gamma}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\mathrm{d}u+\int_{C_{u}}\xi^{-3-\gamma}\Omega^{2}h\frac{\Phi^{2}}{r^{1+b}}\mathrm{d}\underline{u}$
$+\iint_{\mathcal{S}}\xi^{-3-\gamma}\Omega^{2}\left((1+b)h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}+\frac{\gamma-b}{4}\frac{\Phi^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{S}}(3+\gamma)\xi^{-4-\gamma}\left(1-\xi\right)\Omega^{2}h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$
[…]
In particular, the energy estimates (4.1.2) and (4.1.2) indicates that
$\|\alpha\|^{2}_{H^{0}_{w}(C_{u})}+\|\hat{w}\|^{2}_{H^{1}_{w}(\underline{C}_{\underline{u}})}\lesssim\|\alpha\|^{2}_{H^{0}_{w}(C_{u_{0}})}\leq\varepsilon^{2}.$
This completes the first order weighted energy estimate.
The second order weighted energy estimate follows from multiplying $2\xi^{1-\gamma}r^{-1-b}\cdot\Omega^{2}h\alpha^{\prime}$ on (3.1.4), and modifying the proof leading to (3.1.4). Moreover, using Proposition 4.4 and the bootstrap assumptions (4.4a)–(4.4b), we obtain
$\int_{C_{u}}\xi^{1-\gamma}\frac{\Omega^{2}h}{r^{1+b}}\left(\Phi^{2}+(9+2b)\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}\underline{u}$
$+\int_{\underline{C}_{\underline{u}}}\xi^{1-\gamma}\frac{h^{-1}}{r^{1+b}}\left(2\left(w\alpha-\Phi\right)^{2}+\frac{\Phi^{2}}{2}+2h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u$
$+\iint_{\mathcal{S}}(\gamma-b)\xi^{1-\gamma}\frac{\Omega^{2}}{r^{2+b}}(\alpha^{\prime})^{2}\mathrm{d}u\mathrm{d}\underline{u}$
[…]
$+\iint_{\mathcal{S}}\varepsilon\Lambda\xi^{1-\gamma}\frac{\Omega^{2}}{r^{2+b}}(\alpha^{\prime})^{2}\mathrm{d}u\mathrm{d}\underline{u}+\iint_{\mathcal{S}}\varepsilon\Lambda\xi^{1-\gamma}\frac{\Omega^{2}}{r^{2+b}}\left(\alpha^{2}+\Phi^{2}\right)\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{S}}\xi^{-\gamma}\frac{\Omega^{2}}{r^{2+b}}\left(\xi\alpha^{2}+\left(w\alpha-\Phi\right)^{2}+\Phi^{2}+2h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u\mathrm{d}\underline{u}.$
Due to the choice $\gamma>b$, the second term on the right hand side having the small constant $\varepsilon\Lambda$ can be absorbed by the spacetime integral on the left hand side. To bound the other spacetime integrals on the right hand side, we appeal to the first order estimates (4.1.2) and (4.1.2) so that we are able to deduce the energy bound
$\int_{C_{u}}\xi^{1-\gamma}\frac{\Omega^{2}h}{r^{1+b}}\left(\Phi^{2}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}\underline{u}$
$+\int_{\underline{C}_{\underline{u}}}\xi^{1-\gamma}\frac{h^{-1}}{r^{1+b}}\left(\left(w\alpha-\Phi\right)^{2}+\Phi^{2}+h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u$
$+\iint_{\mathcal{S}}\frac{(\gamma-b)}{2}\xi^{1-\gamma}\frac{\Omega^{2}}{r^{2+b}}(\alpha^{\prime})^{2}\mathrm{d}u\mathrm{d}\underline{u}$
[…]
Once the energy estimates are completed, we are able to improve the pointwise estimates.
4.1.3. Pointwise estimates
In the following lemmas 4.54.7, we will improve the $L^{\infty}$ estimates for $r^{-1}\hat{w},\,\alpha$, $\mu$ and $\big{|}rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\big{|}$ to close the bootstrap argument.
Lemma 4.5.
We have the refine estimate
$\xi^{-1-\gamma}\mu\lesssim\varepsilon^{2}r^{b},\quad\text{in}\,\,\mathcal{S}.$
Proof.
The proof is analogous to that of Lemma 3.4. By the mass formula,
$\frac{m}{r^{1+b}\xi^{1+\gamma}}={}$ $\frac{1}{r^{1+b}\xi^{1+\gamma}}\int_{r_{\ast}}^{r}\left((1-\mu)\alpha^{2}+\Phi^{2}\right)\mathrm{d}r^{\prime}|_{C_{u}},$
[…]
where in the second line we have used the fact $\partial_{r}\xi\geq 0$ along $C_{u}\cap\mathcal{S}$. ∎
Lemma 4.6.
We have in $\mathcal{S}$
$\xi^{-\gamma}\alpha^{2}\lesssim\varepsilon^{2}r^{b},\quad\xi^{-2-\gamma}r^{-2}\hat{w}^{2}\lesssim\varepsilon^{2}r^{b}.$
Proof.
We will first show that $\|\alpha\|^{2}_{H^{1}_{w}(C_{u_{0}})}\lesssim\varepsilon^{2}$ implies
$\|\xi^{-\gamma}r^{-b}\alpha^{2}\|_{L^{\infty}(C_{u_{0}}\cap\mathcal{S})}\lesssim\varepsilon^{2}.$ (4.9)
Integrating along $\partial_{r}$ on $C_{u_{0}}$, we obtain the following identity
$\xi^{-\gamma}r^{-b}\alpha^{2}\big{|}_{C_{u_{0}}}(\underline{u})-\xi^{-\gamma}r^{-b}\alpha^{2}\big{|}_{C_{u_{0}}}(\underline{u}_{\ast})$
[…]
Noting that $\xi^{-\gamma}r^{-b}\alpha^{2}\big{|}_{C_{u_{0}}}(\underline{u}_{\ast})=0$, we have
$\xi^{-\gamma}r^{-b}\alpha^{2}\big{|}_{C_{u_{0}}}(\underline{u})+\int_{C_{u_{0}}}\gamma\frac{\alpha^{2}}{\xi^{1+\gamma}r^{1+b}}\mathrm{d}r$
[…]
which justifies (4.9).
Next, we prove pointwise estimate for $\alpha$, using (4.9).
Multiplying $2\xi^{-\gamma}r^{-b}\alpha$ on (2.15) and integrating along $\partial_{u}$ on each $\underline{C}_{\underline{u}}$, noting that $\gamma>b$, we have
$\big{|}\xi^{-\gamma}r^{-b}\alpha^{2}-\xi^{-\gamma}r^{-b}\alpha^{2}|_{u_{0}}\big{|}+\int_{\underline{C}_{\underline{u}}}\left(\gamma-b\right)(1-\mu)\xi^{-\gamma}\frac{\alpha^{2}}{r^{1+b}}h^{-1}\mathrm{d}u$
[…]
$+\int_{\underline{C}_{\underline{u}}}\gamma\Big{|}r\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\Big{|}\xi^{-\gamma}\frac{\alpha^{2}}{r^{1+b}}\mathrm{d}u$
[…]
$+\int_{\underline{C}_{\underline{u}}}\left(2\mu+\gamma\big{|}rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\big{|}\right)\xi^{-\gamma}\frac{\alpha^{2}}{r^{1+b}}h^{-1}\mathrm{d}u.$
By the bootstrap assumptions, $\big{|}rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\big{|}$ and $\mu$ are both small enough, then
$(\gamma-b)(1-\mu)>\mu+\gamma\big{|}rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\big{|}.$
After absorbing the error terms on the right hand side, we conclude
$\xi^{-\gamma}r^{-b}\alpha^{2}+\int_{\underline{C}_{\underline{u}}}\frac{\gamma-b}{2}\xi^{-\gamma}\frac{\alpha^{2}}{r^{1+b}}h^{-1}\mathrm{d}u$
[…]
where in the last inequality we have appealed to (4.9) and the energy estimates (4.1.2) and (4.1.2).
Meanwhile, integrating $\xi^{-2-\gamma}r^{-b-2}\hat{w}^{2}$ along $\partial_{\underline{u}}$ on each $C_{u}$ gives rise to
$\xi^{-2-\gamma}r^{-b-2}\hat{w}^{2}{}$ $+(2+\gamma)\int_{C_{u}}\xi^{-3-\gamma}(1-\xi)\frac{r^{-2}\hat{w}^{2}}{r^{1+b}}\Omega^{2}h\mathrm{d}\underline{u}$
$+(2+b)\int_{C_{u}}\xi^{-2-\gamma}\frac{r^{-2}\hat{w}^{2}}{r^{1+b}}\Omega^{2}h\mathrm{d}\underline{u}$
$\lesssim$ $\int_{C_{u}}\left(\xi^{-1-\gamma}\frac{\alpha^{2}}{r^{1+b}}+\xi^{-3-\gamma}\frac{r^{-2}\hat{w}^{2}}{r^{1+b}}\right)\Omega^{2}h\mathrm{d}\underline{u}\lesssim\varepsilon^{2}.$
We complete the proof of this lemma. ∎
Lemma 4.7.
We have the improvement,
$\Big{|}rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\Big{|}\lesssim\varepsilon^{2},\quad\text{in}\,\,\mathcal{S}.$
Proof.
By means of a straightforward calculation, using the mean value theorem, we find that
$rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}={}$ $\frac{rh\underline{h}-r_{\ast}h_{\ast}\underline{h}_{\ast}+r_{\ast}h_{\ast}\underline{h}_{\ast}-rh\underline{h}_{\ast}}{r-r_{\ast}}$
[…]
which is small in view of Lemma 4.6 and the fact
$\big{|}\ln\frac{h|_{r_{2}}}{h_{\ast}}\big{|}\lesssim\int^{r_{2}}_{r_{\ast}}\frac{\alpha^{2}}{r}\mathrm{d}r\lesssim\varepsilon^{2}.$
What is more, along the lines of the subsection 3.1.6, we are able to prove the boundedness of the following geometric quantities,
$|\ln(\Omega^{2}h)|+|\ln(-\underline{h})|+|\ln h|.$
In the end, we present the pointwise estimates for $\partial_{u}w$ and $\alpha^{\prime}$.
Lemma 4.8.
In the case of $0<b<\frac{\sqrt{89}-9}{4}$, there is
$\xi^{-3-\gamma}|\partial_{u}w|^{2}\lesssim\varepsilon^{2}r^{b}.$
The proof resembles the one in the second part of Lemma 4.6.
Lemma 4.9.
Suppose $\xi^{2-\gamma}r^{-b}(\alpha^{\prime})^{2}\big{|}_{C_{u_{0}}}$ is bounded, then we have the following $L^{\infty}$ estimate for $\alpha^{\prime}$,
$\xi^{2-\gamma}(\alpha^{\prime})^{2}\lesssim r^{b}.$
Proof.
Multiply $2\xi^{2-\gamma}r^{-b}\alpha^{\prime}$ on the second order equation (3.1.4) and integrate along $\partial_{u}$. Then
$\xi^{2-\gamma}r^{-b}(\alpha^{\prime})^{2}+\int_{\underline{C}_{\underline{u}}}(\gamma-b)\xi^{2-\gamma}\frac{(\alpha^{\prime})^{2}}{r^{1+b}}h^{-1}\mathrm{d}u$
[…]
$+\int_{\underline{C}_{\underline{u}}}\left(c+\varepsilon\right)\xi^{2-\gamma}\frac{(\alpha^{\prime})^{2}}{r^{1+b}}h^{-1}\mathrm{d}u,$
where we take the constant $c>0$ to be small, so that $c+\varepsilon$ can be absorbed by the positive integral on the left hand side, noting $02\geq\gamma>b>$ 0. Thus this lemma is concluded.
4.2. Cauchy stability in weighted $H^{1}$ space
The proof of Theorem 4.1 itself leads to the Cauchy stability theorem in $H^{1}_{w}$ space.
Theorem 4.10.
[…]
$\int_{C^{[\underline{u}_{\ast},\underline{u}_{1}]}_{u_{0}}}\frac{\Omega^{2}h}{r^{1+b}}\left(\frac{r^{-2}\hat{w}^{2}}{\xi^{3+\gamma}}+\frac{\alpha^{2}}{\xi^{1+\gamma}}+\xi^{1-\gamma}\frac{(\alpha^{\prime})^{2}}{2}\right)\mathrm{d}\underline{u}<\varepsilon^{2},$
and assume $|r-r_{\ast}|^{-\frac{\gamma}{2}}\alpha|_{C_{u_{0}}}$ is continuous at $\underline{u}=\underline{u}_{\ast}$. There exists an $\varepsilon_{0}>0$ such that if $\varepsilon^{2}\leq\varepsilon^{2}_{0}$, the following conclusions hold.
Case I: $b>0$. There is a unique solution
$(r,\,m,\,\alpha)\in C^{1}\times C^{1}\times H^{1}_{w}\quad\text{in}\,\,\mathcal{D}(u_{0},\,u;\,\underline{u}_{1}),\,\,u\in[u_{0},\,\underline{u}_{\ast}).$
Moreover, for any $(u,\,\underline{u})\in[u_{0},\,\underline{u}_{\ast})\times[\underline{u}_{\ast},\,\underline{u}_{1}]$, there is
$\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(\frac{(\partial_{u}w)^{2}}{\xi^{3+\gamma}}+\frac{r^{-2}\hat{w}^{2}}{\xi^{1+\gamma}}+\xi^{1-\gamma}\alpha^{2}\right)\mathrm{d}u$
$+\int_{C_{u}}\frac{\Omega^{2}h}{r^{1+b}}\left(\frac{r^{-2}\hat{w}^{2}}{\xi^{3+\gamma}}+\frac{\alpha^{2}}{\xi^{1+\gamma}}+\xi^{1-\gamma}(\alpha^{\prime})^{2}\right)\mathrm{d}\underline{u}\lesssim\varepsilon^{2},$
and we have the following pointwise estimates
$\xi^{-2-\gamma}r^{-2}\hat{w}^{2}+\xi^{-\gamma}\alpha^{2}+\xi^{-3-\gamma}|\partial_{u}w|^{2}+\xi^{-1-\gamma}\mu\lesssim\varepsilon^{2}r^{b}.$
Case II: $b=0$. There is a unique solution
$(r,\,m,\,\alpha)\in C^{1}\times C^{0}\times H^{1}_{w},\quad\text{in}\,\,\mathcal{D}(u_{0},\,u;\,\underline{u}_{1}),\,\,u\in[u_{0},\,\underline{u}_{\ast}).$
The estimates in Case I hold (with $b=0$) except the $L^{\infty}$ estimate for $\partial_{u}w$.
Appendix A Local existence theorem
For the characteristic problem where initial data are prescribed on a future light cone with vertex at the center of symmetry, the local existence theorem for the classical solution to the spherically symmetric Einstein scalar field system was proved in [8]. Regarding to the proof, Christodoulou reduced the whole system in Bondi coordinates to a scalar equation and utilized the contracting mapping principle to obtain the existence theorem. However, in this section, we mean to provide an alternative proof based on the energy method. To be more precisely, we work within the double null gauge and demonstrate a local existence theory for the SSEYM system in $H^{1}$ (and $H^{1}_{w}$) space, using an iteration method.
Theorem A.1.
Let $\underline{u}_{1}\in(u_{0},\,+\infty)$ be a finite constant. Suppose on the initial slice $C_{u_{0}}$, the following energy is bounded
$\int_{C_{u_{0}}}r^{-1-b}\left(\frac{\hat{w}^{2}}{r^{2}}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}r<I_{1}.$
Case I: $0<b<\frac{\sqrt{89}-9}{4}$. i) There exists $u_{\ast}>u_{0}$ such that the SSEYM system has a unique solution $\left(r,\,m,\,\alpha\right)\in C^{1}\times C^{1}\times H^{1}$ on the domain $\mathcal{D}(u_{0},u_{\ast};\underline{u}_{1})$, which enjoys the subsequent bound,
$\int_{C_{u}}r^{-1-b}\left(\frac{\hat{w}^{2}}{r^{2}}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}r\lesssim I_{1},$
$\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(\frac{\hat{w}^{2}}{r^{2}}+\alpha^{2}+h^{2}(\partial_{u}\hat{w})^{2}\right)\mathrm{d}u\lesssim I_{1},$
$|r^{-1}\hat{w}|^{2}+|\partial_{u}\hat{w}|+|\alpha|^{2}+\mu\lesssim I_{1}r^{b},$
$|\ln\left(\Omega^{2}h\right)|,\,|\ln\left(-\underline{h}\right)|,\,|\ln h|\lesssim I_{1}.$
ii) The solution depends continuously on data in the following sense. Suppose $\left(r_{n},\,m_{n},\,\hat{w}_{n}\right)$ and $\left(r_{m},\,m_{m},\,\hat{w}_{m}\right)$ are solutions with data $\hat{w}_{0n}$, $\hat{w}_{0m}$ respectively, then
$\int_{C_{u}}\left(\frac{|\hat{w}_{n}-\hat{w}_{m}|^{2}}{r^{3+b}_{n}}+\frac{|\alpha_{n}-\alpha_{m}|^{2}}{r^{1+b}_{n}}+\frac{|\alpha^{\prime}_{n}-\alpha^{\prime}_{m}|^{2}}{r^{1+b}_{n}}\right)\mathrm{d}r_{n}$
$+\int_{\underline{C}_{\underline{u}}}\frac{h_{n}^{-1}}{r^{1+b}_{n}}\left(\frac{\left(\hat{w}_{n}-\hat{w}_{m}\right)^{2}}{r_{n}^{2}}+\left(\alpha_{n}-\alpha_{m}\right)^{2}+h_{n}^{2}(\partial_{u}\hat{w}_{n}-\partial_{u}\hat{w}_{m})^{2}\right)\mathrm{d}u$
[…]
Case II: $b=0$. The statement in Case I holds with $b=0$, except that the pointwise estimate for $|\partial_{u}w|$ is absent and hence in this case the solution $\left(r,\,m,\,\alpha\right)\in C^{1}\times C^{0}\times H^{1}$.
Remark A.2.
We can prove the local existence theorem in $H^{1}_{w}$ space (or rather the space of data in Theorem 4.1) with $0\leq b<\frac{\sqrt{89}-9}{4}$ and $b<\gamma\leq 2$ straightforwardly in the same manner as Theorem A.1. Thus, we omit the details.
Alternatively, since $H^{1}$ is dense in $H^{1}_{w}$ (actually, smooth functions with compact support are dense in $H^{1}_{w}$, see Appendix C), we can make use of the property ii) that the solution continuously depends on data, to verify the local existence in $H^{1}_{w}$ space.
Proof.
For the sake of simplicity, we only carry out the proof for the case of $b=0$ in details. The proof for the other case $0<b<\frac{\sqrt{89}-9}{4}$ goes through in an analogous fashion.
It suffices to prove the local existence theorem near the axis, which essentially turns the local existence problem into the one with small data. Away from the axis, the local existence follows in a standard way.
Choose $\underline{u}_{1}>u_{0}$ and $\underline{u}_{1}-u_{0}$ sufficiently small so that the initial energy is sufficiently small. We denote the size of data by $\varepsilon^{2}I_{1}$, where $\varepsilon^{2}$ is small enough, namely,
$\int_{C_{u_{0}}^{[u_{0},\,\underline{u}_{1}]}}r^{-1-b}\left(\frac{\hat{w}^{2}}{r^{2}}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}r\leq\varepsilon^{2}I_{1}.$
We shall find a local solution for the SSEYM system using the method of Picard iteration. Notice that, the Yang–Mills equation can be alternatively written in terms of $\hat{w}$ (1.3)
$\partial_{u}\partial_{\underline{u}}\hat{w}+2\frac{\Omega^{2}}{r^{2}}\hat{w}=\Omega^{2}\frac{\hat{w}^{2}}{r^{2}}\left(\hat{w}+3\right).$ (A.1)
[…]
$\partial_{u}\partial_{\underline{u}}\hat{w}_{k+1}+{}$ $2\Omega_{k}^{2}\frac{\hat{w}_{k+1}}{r_{k}^{2}}=\Omega_{k}^{2}\frac{\hat{w}_{k}^{2}}{r_{k}^{2}}\left(\hat{w}_{k}+3\right),$ (A.2)
$\partial_{\underline{u}}m_{k+1}={}$ $\Omega^{2}_{k}h_{k}\left(\frac{\Phi_{k+1}^{2}}{2}-h_{k}\underline{h}_{k}\alpha_{k+1}^{2}\right),$ (A.3)
$\partial_{u}m_{k+1}={}$ $\underline{h}_{k}\left(\frac{\Phi_{k+1}^{2}}{2}-\left(h_{k}\underline{h}_{k}\right)^{-1}h^{2}_{k}(\partial_{u}\hat{w}_{k+1})^{2}\right),$ (A.4)
$\partial_{u}\left(\Omega^{2}_{k+1}h_{k+1}\right)={}$ $\Omega^{2}_{k+1}\underline{h}^{-1}_{k+1}\underline{h}_{k}\left(-\frac{2m_{k+1}}{r^{2}_{k}}+\frac{\Phi_{k+1}^{2}}{r_{k}}\right),$ (A.5)
$\partial_{\underline{u}}\left(\underline{h}_{k+1}\right)={}$ $\Omega^{2}_{k}h_{k}h_{k+1}^{-1}\left(-\frac{2m_{k+1}}{r^{2}_{k}}+\frac{\Phi_{k+1}^{2}}{r_{k}}\right),$ (A.6)
with initial data prescribed in Theorem A.1 and the boundary conditions
$m_{k+1}|_{\Gamma}=0,\quad r_{k+1}|_{\Gamma}=0,\quad\hat{w}_{k+1}|_{\Gamma}=0.$
Here we set
$\Omega_{k+1}^{2}h_{k+1}:={}$ $\partial_{\underline{u}}r_{k+1},$ $\underline{h}_{k+1}:=\partial_{u}r_{k+1},$
$\Phi_{k+1}:={}$ $\frac{(w_{k+1}^{2}-1)}{r_{k}},$ $\alpha_{k+1}:=\Omega^{-2}_{k}h_{k}^{-1}\partial_{\underline{u}}\hat{w}_{k+1},$
and $h_{k+1}$ is related to $m_{k+1}$ via
$2m_{k+1}:=r_{k}\left(1+h_{k+1}\underline{h}_{k+1}\right).$
We further define
$\mu_{k+1}:=\frac{2m_{k+1}}{r_{k}}=\left(1+h_{k+1}\underline{h}_{k+1}\right),\quad\alpha^{\prime}_{k+1}:=r_{k}\Omega^{-2}_{k}h_{k}^{-1}\partial_{\underline{u}}\alpha_{k+1}.$
We insist the choice of gauge in Section 2.2 so that
$\Omega_{k+1}^{2}h_{k+1}|_{C_{u_{0}}}=\frac{1}{2}.$ (A.7)
Besides, our gauge choice
$u=\underline{u}\quad\text{on}\,\,\Gamma$
together with the boundary conditions $r_{k+1}|_{\Gamma}=0$, $\hat{w}_{k+1}|_{\Gamma}=0$ entails
$\left(\underline{h}_{k+1}+\Omega^{2}_{k+1}h_{k+1}\right)|_{\Gamma}=0,$
$\left(\partial_{u}\hat{w}_{k+1}+\partial_{\underline{u}}\hat{w}_{k+1}\right)|_{\Gamma}=0.$
We remark that, for the first generation, we have $\underline{h}_{0}=-\frac{1}{2}$, and $h_{0}=2$, $\Omega^{2}_{0}=\frac{1}{4}$. In other words, the first generation is the vacuum spacetime.
In the linearized system (A.2)-(A.6), we first use (A.2) to solve $\hat{w}_{k+1}$. Next, substituting the $\hat{w}_{k+1}$ into (A.3) and (A.4), we can further solve $m_{k+1}$. After that, substituting both of $\hat{w}_{k+1}$ and $m_{k+1}$ into (A.5) and (A.6), we are able to solve $\Omega^{2}_{k+1}h_{k+1}$ and $\underline{h}_{k+1}$, and hence obtain $r_{k+1}$.
Auxiliary equations. Along with the main equations, we can derive the following two equations from (A.3) and (A.6), (A.4) and (A.5).
$\partial_{\underline{u}}h_{k+1}={}$ $2\Omega^{2}_{k}h_{k}(1-\mu_{k})\underline{h}_{k+1}^{-1}\frac{(\alpha_{k+1})^{2}}{r_{k}},$ (A.8)
$\partial_{u}(\Omega^{-2}_{k+1}\underline{h}_{k+1})={}$ $-2\Omega^{-2}_{k+1}h^{-1}_{k+1}h_{k}\frac{(\partial_{u}\hat{w}_{k+1})^{2}}{r_{k}}.$ (A.9)
With the above setup, we now continue the proof with $H^{1}$ estimates.
Uniform bound. We claim that if $\varepsilon$ (or rather $\underline{u}_{1}-u_{0}$) is small enough, there is a constant $M<+\infty$ such that we have the uniform bounds, for all $i\geq 0$,
$A_{i+1}(u,\underline{u}):=$ $\int_{C_{u}}\Omega^{2}_{i}h_{i}\left(\frac{(\alpha^{\prime}_{i+1})^{2}}{r_{i}}+\frac{\alpha_{i+1}^{2}}{r_{i}}+\frac{\hat{w}_{i+1}^{2}}{r^{3}_{i}}\right)\mathrm{d}u$
$+\int_{\underline{C}_{\underline{u}}}h_{i}^{-1}\left(\frac{\alpha_{i+1}^{2}}{r_{i}}+h_{i}^{2}\frac{(\partial_{u}\hat{w}_{i+1})^{2}}{r_{i}}+\frac{\hat{w}_{i+1}^{2}}{r^{3}_{i}}\right)\mathrm{d}u$
[…]
When $i=0$, an energy argument (in line with the proof leading to (3.1.4) in the subsection 3.1.4) for the linearized equation
$\partial_{u}\partial_{\underline{u}}\hat{w}_{1}+2\Omega^{2}_{0}\frac{\hat{w}_{1}}{r_{0}^{2}}=0,$
yields
$\int_{C_{u}}\frac{\Omega^{2}_{0}h_{0}}{r_{0}}\left((\alpha^{\prime}_{1})^{2}+9\alpha_{1}^{2}+4r_{0}^{-2}\hat{w}_{1}^{2}\right)\mathrm{d}\underline{u}$
$+\int_{\underline{C}_{\underline{u}}}\frac{h_{0}^{-1}}{r_{0}}\left(2\alpha_{1}^{2}+2h_{0}^{2}(\partial_{u}\hat{w}_{1})^{2}+8r_{0}^{-2}\hat{w}_{1}^{2}\right)\mathrm{d}u$
$+\iint_{\mathcal{D}}\frac{\Omega^{2}_{0}}{r^{2}_{0}}\left(18r^{-2}_{0}\hat{w}_{1}^{2}+2h_{0}^{2}(\partial_{u}\hat{w}_{1})^{2}+\alpha^{2}_{1}+(\alpha^{\prime}_{1})^{2}\right)\mathrm{d}u\mathrm{d}\underline{u}$
[…]
Thus, (A) holds for $i=0$, which further implies
$\mu_{1}+\alpha^{2}_{1}+\frac{\hat{w}^{2}_{1}}{r_{0}^{2}}\lesssim\varepsilon^{2}I_{1},$
$|\ln(\Omega_{1}h_{1})|,\,|\ln(-\underline{h}_{1})|,\,|\ln h_{1}|\lesssim I_{1}.$
To proceed, let us suppose the estimate (A) holds for $i\leq k-1$ with the constant $M$ to be determined later. Then we can derive the pointwise estimates, as in the subsections 3.1.5 and 3.1.6
$\mu_{i+1}+\alpha^{2}_{i+1}+\frac{\hat{w}^{2}_{i+1}}{r_{i}^{2}}\lesssim\varepsilon^{2}M,\quad i\leq k-1,$
$|\ln(\Omega_{i+1}h_{i+1})|,\,|\ln(-\underline{h}_{i+1})|,\,|\ln h_{i+1}|\lesssim I_{1},\quad i\leq k-1.$
Moreover, using the equation of (A.5) with $k$ replaced by $k-1$ or $k-2$, and noting the gauge choice (A.7), we have
$\Big{|}\ln\frac{\Omega_{k-1}^{2}h_{k-1}}{\Omega_{k}^{2}h_{k}}\Big{|}\lesssim\varepsilon^{2}M,\quad\Big{|}\ln\frac{\Omega_{k-2}^{2}h_{k-2}}{\Omega_{k-1}^{2}h_{k-1}}\Big{|}\lesssim\varepsilon^{2}M.$
Here we set $r_{i}=r_{0}$, $\Omega^{2}_{i}h_{i}=\Omega^{2}_{0}h_{0}$ if $i<0$. These estimates plus the identity
$\Omega_{k-1}^{-2}h^{-1}_{k-1}\partial_{\underline{u}}(r_{k}-r_{k-1})=\left(-1+\frac{\Omega_{k}^{2}h_{k}}{\Omega_{k-1}^{2}h_{k-1}}\right),$
suggest the equivalence of $r_{k}$, $r_{k-1}$, $r_{k-2}$, if $\varepsilon$ is small enough,
$\frac{1}{2}<\big{|}\frac{r_{k}}{r_{k-1}}\big{|},\,\,\big{|}\frac{r_{k-1}}{r_{k-2}}\big{|}<\frac{3}{2}.$
In addition, taking the equation of (A.6) (with $k$ replaced by $k-1$) and the gauge condition of $\underline{h}_{k}|_{\Gamma}=-\Omega^{2}_{k}h_{k}|_{\Gamma}$ into account, we find that
$\Big{|}\ln\frac{\underline{h}_{k-1}}{\underline{h}_{k}}\Big{|}\lesssim\varepsilon^{2}M,$
and hence
$\Big{|}\ln\frac{h_{k-1}}{h_{k}}\Big{|}\lesssim\varepsilon^{2}M$
follows, since $\frac{h_{k-1}}{h_{k}}=\frac{h_{k-1}\underline{h}_{k-1}}{h_{k}\underline{h}_{k}}\cdot\frac{\underline{h}_{k}}{\underline{h}_{k-1}}$.
We shall then prove (A) holds for $i=k$. Thanks to the preceding estimates and an energy argument in line with the subsection 3.1.4, we are able to derive
$\int_{C_{u}}\Omega^{2}_{k}h_{k}\left(\frac{(\alpha^{\prime}_{k+1})^{2}}{r_{k}}+9\frac{\alpha_{k+1}^{2}}{r_{k}}+4\frac{\hat{w}_{k+1}^{2}}{r_{k}}\right)\mathrm{d}\underline{u}$
$+\int_{\underline{C}_{\underline{u}}}\frac{h_{k}^{-1}}{r_{k}}\left(2\left(\alpha_{k+1}-2r_{k}^{-1}\hat{w}_{k+1}\right)^{2}+2h_{k}^{2}(\partial_{u}\hat{w}_{k+1})^{2}+2r_{k}^{-2}\hat{w}_{k+1}^{2}\right)\mathrm{d}u$
$+\iint_{\mathcal{D}}\frac{1}{2}\Omega^{2}_{k}r^{-2}_{k}\left(18r^{-2}_{k}\hat{w}_{k+1}^{2}+2h_{k}^{2}(\partial_{u}\hat{w}_{k+1})^{2}+\alpha^{2}_{k+1}+(\alpha^{\prime}_{k+1})^{2}\right)\mathrm{d}u\mathrm{d}\underline{u}$
[…]
where the universal constant $C$ is independent of $k$. If we choose $M>CI_{1}$, then (A) with $i=k$ is verified.
Convergence. In the compact domain concerned, since $\{r_{k}\}$ is equi-continuous, there is a subsequence, which we still denote by $\{r_{k}\}$, such that $r_{k}\rightarrow r$ on the compact neighbourhood. In what follows, we always use this converging subsequence to replace the original one.
Based on (A.2), we derive the following linear equation,
$\partial_{u}\partial_{\underline{u}}(\hat{w}_{k+1}-\hat{w}_{k})+2\frac{\Omega_{k-1}^{2}}{r_{k-1}^{2}}(\hat{w}_{k+1}-\hat{w}_{k})$
[…]
where $U_{k}-U_{k-1}$ involves
$\frac{r_{k}-r_{k-1}}{r_{k-1}},\,\,h^{-1}_{k}-h^{-1}_{k-1},\,\,\Omega^{2}_{k}h_{k}-\Omega^{2}_{k-1}h_{k-1},\,\,\frac{\hat{w}_{k}}{r_{k-1}}-\frac{\hat{w}_{k-1}}{r_{k-2}},$
and
$\Big{|}f\left(\frac{\hat{w}_{k+1}}{r_{k}},\,\frac{\hat{w}_{k}}{r_{k}},\,\frac{\hat{w}_{k-1}}{r_{k-1}}\right)\Big{|}\lesssim\big{|}\frac{\hat{w}_{k+1}}{r_{k}}\big{|}+\big{|}\frac{\hat{w}_{k}}{r_{k}}\big{|}+\big{|}\frac{\hat{w}_{k-1}}{r_{k-1}}\big{|}.$
In general, we expect $|U_{k}-U_{k-1}|$ can be bounded by $\|\hat{w}_{k}-\hat{w}_{k-1}\|_{H^{1}}$.
We will prove the convergence by induction. Suppose
$\|U_{k}-U_{k-1}\|_{L^{\infty}}\leq M2^{-k+1},\quad\|\mu_{k}-\mu_{k-1}\|_{L^{\infty}}\leq M2^{-k+1},$ (A.13)
$\Big{|}\int_{C_{u}}\frac{m_{k}}{r_{k-1}^{2}}\mathrm{d}r_{k-1}|_{C_{u}}-\int_{C_{u}}\frac{m_{k-1}}{r_{k-2}^{2}}\mathrm{d}r_{k-2}|_{C_{u}}\Big{|}\leq M2^{-k+1},$ (A.14)
$\Big{|}\int_{\underline{C}_{\underline{u}}}\frac{m_{k}}{r_{k-1}^{2}}\mathrm{d}r_{k-1}|_{\underline{C}_{\underline{u}}}-\int_{\underline{C}_{\underline{u}}}\frac{m_{k-1}}{r_{k-2}^{2}}\mathrm{d}r_{k-2}|_{\underline{C}_{\underline{u}}}\Big{|}\leq M2^{-k+1},$ (A.15)
for a large constant $M$ to be determined. When $k=1$, (A.13)-(A.15) hold by the uniform boundedness in the previous step. We will prove the same inequalities hold with $k$ replaced by $k+1$.
As is obvious from the expression, the linear equation (A) shares the same structure as (A.2). Thus, we can establish an analogous energy identity, noting that data vanishes now
$\int_{C_{u}}h_{k-1}\Omega^{2}_{k-1}\frac{|X_{k+1}-X_{k}|^{2}}{r_{k-1}}\mathrm{d}\underline{u}+\int_{\underline{C}_{\underline{u}}}h^{-1}_{k-1}\frac{|Y_{k+1}-Y_{k}|^{2}}{r_{k-1}}\mathrm{d}u$
$+\iint_{\mathcal{D}}\Omega^{2}_{k-1}r^{-2}_{k-1}|V_{k+1}-V_{k}|^{2}\mathrm{d}u\mathrm{d}\underline{u}$
[…]
where, for simplicity, we denote
$X_{k+1}-X_{k}={}$ $\left(r_{k-1}\partial_{r_{k-1}}^{2}(\hat{w}_{k+1}-\hat{w}_{k}),\,\partial_{r_{k-1}}(\hat{w}_{k+1}-\hat{w}_{k}),\,r_{k-1}^{-1}(\hat{w}_{k+1}-\hat{w}_{k})\right),$
$Y_{k+1}-Y_{k}={}$ $\left(\partial_{r_{k-1}}(\hat{w}_{k+1}-\hat{w}_{k}),\,\partial_{u}(\hat{w}_{k+1}-\hat{w}_{k}),\,r_{k-1}^{-1}(\hat{w}_{k+1}-\hat{w}_{k})\right),$
$V_{k+1}-V_{k}={}$ $\left(X_{k+1}-X_{k},\,\partial_{u}(\hat{w}_{k+1}-\hat{w}_{k})\right).$
Then we estimate the right hand side of the inequality as follows,
[…]
$\qquad\cdot\left(\iint_{\mathcal{D}}\Omega_{k-1}^{2}\frac{|V_{k+1}-V_{k}|^{2}}{r^{2}_{k-1}}\mathrm{d}u\mathrm{d}\underline{u}\right)^{\frac{1}{2}}$
[…]
As a result, an application of the Cauchy-Schwarz leads to
$\int_{C_{u}}h_{k-1}\Omega^{2}_{k-1}\frac{|X_{k+1}-X_{k}|^{2}}{r_{k-1}}\mathrm{d}\underline{u}+\int_{\underline{C}_{\underline{u}}}h^{-1}_{k-1}\frac{|Y_{k+1}-Y_{k}|^{2}}{r_{k-1}}\mathrm{d}u$
$+\iint_{\mathcal{D}}\frac{1}{2}\Omega^{2}_{k-1}r^{-2}_{k-1}|V_{k+1}-V_{k}|^{2}\mathrm{d}u\mathrm{d}\underline{u}$
[…]
For notational convenience, in what follows, we will denote
$\|\hat{w}_{k+1}-\hat{w}_{k}\|^{2}_{H^{1}(C_{u})}$
[…]
$\|\hat{w}_{k+1}-\hat{w}_{k}\|^{2}_{H^{2}(C_{u})}$
[…]
Convergence of $\{\hat{w}_{k}\}$ uniformly in the energy norm. By (A.13) and the fact $\varepsilon$ is small, the estimate (A) tells that
$\|\hat{w}_{k+1}-\hat{w}_{k}\|^{2}_{H^{2}}(C_{u})+\int_{\underline{C}_{\underline{u}}}h^{-1}_{k-1}\frac{|Y_{k+1}-Y_{k}|^{2}}{r_{k-1}}\mathrm{d}u$
$\qquad+\iint_{\mathcal{D}}\Omega^{2}_{k-1}r^{-2}_{k-1}|V_{k+1}-V_{k}|^{2}\mathrm{d}u\mathrm{d}\underline{u}$
[…]
Moreover, by virtue of the fact $\frac{|\hat{w}_{k+1}-\hat{w}_{k}|^{2}}{r_{k-1}^{2}}\lesssim\|\hat{w}_{k+1}-\hat{w}_{k}\|^{2}_{H^{1}(C_{u})}$, the previous $L^{\infty}$ estimates and (A.13), it follows that
$\Big{|}\frac{\hat{w}_{k+1}}{r_{k}}-\frac{\hat{w}_{k}}{r_{k-1}}\Big{|}\lesssim{}$ $\Big{|}\frac{\hat{w}_{k+1}}{r_{k}}\Big{|}\Big{|}1-\frac{r_{k}}{r_{k-1}}\Big{|}+\|\hat{w}_{k+1}-\hat{w}_{k}\|_{H^{1}}$
[…]
if we choose $\varepsilon$ small enough such that $2\sqrt{\varepsilon}C<\frac{1}{2}$ and $M>1$.
Convergence of $\mu_{k}$ (uniformly in $C^{0}$ norm). Noting that
$\frac{m_{k+1}}{r_{k}}-\frac{m_{k}}{r_{k-1}}={}$ $\frac{m_{k+1}-m_{k}}{r_{k}}+\frac{m_{k}}{r_{k}}\left(1-\frac{r_{k}}{r_{k-1}}\right),$ (A.19)
and using (A.3), we obtain
$\frac{m_{k+1}-m_{k}}{r_{k}}\lesssim{}$ $\|\hat{w}_{k+1}\|^{2}_{H^{1}}\left(|U_{k}-U_{k-1}|+|\mu_{k}-\mu_{k-1}|\right)$
$+\left(\|\hat{w}_{k+1}\|_{H^{1}}+\|\hat{w}_{k}\|_{H^{1}}\right)\|\hat{w}_{k+1}-\hat{w}_{k}\|_{H^{1}}.$
Then it follows from the inductive assumptions (A.13) and (A) that
$|\mu_{k+1}-\mu_{k}|\leq\varepsilon CM2^{-k+1}.$ (A.20)
We conclude through choosing $\varepsilon$ and $M$ properly that
$|\mu_{k+1}-\mu_{k}|\leq M2^{-k}.$
Furthermore, in view of the formulae (3.1.6) and (3.1.6), it follows from (A) and (A.20) that (A.14)–(A.15) hold with $k$ replaced by $k+1$.
Convergence of $\Omega^{2}_{k}h_{k}$, $\underline{h}_{k}$, $h_{k}$, $m_{k}$ (uniformly in $C^{0}$ norm). Similarly, using the equations (A.5) and (A.6), we prove that
$\Big{|}\ln\frac{\Omega^{2}_{k+1}h_{k+1}}{\Omega^{2}_{k}h_{k}}\Big{|}\leq\varepsilon CM2^{-k+1}\leq M2^{-k},$
$\Big{|}\ln\frac{\underline{h}_{k+1}}{\underline{h}_{k}}\Big{|}\leq\varepsilon CM2^{-k+1}\leq M2^{-k},$
which together with (A.20) implies $|h_{k+1}^{-1}-h^{-1}_{k}|\leq\varepsilon M2^{-k}$. The inductive proof is thus complete.
Finally, substituting all the previous results into the equation of mass (A.3), we obtain
$m_{k}-m_{k-1}=O(2^{-k}).$
Integrating $\Omega_{k}^{-2}h^{-1}_{k}\partial_{\underline{u}}(r_{k}-r_{k+1})=\left(1-\frac{\Omega_{k+1}^{2}h_{k+1}}{\Omega_{k}^{2}h_{k}}\right)$, we have
$|r_{k+1}-r_{k}|\lesssim r_{k}\big{|}1-\frac{\Omega_{k+1}^{2}h_{k+1}}{\Omega_{k}^{2}h_{k}}\big{|}_{C^{0}}.$
Then
$\big{|}\frac{r_{k+1}}{r_{k}}-1\big{|}=O(2^{-k}).$
Due to the uniform convergences, there exist $\mu$, $m$ and $\Omega^{2}h$, $\underline{h}$ such that
$\mu_{k}\rightarrow\mu\quad\text{uniformly in}\,C^{0}\,\text{norm},$
$m_{k}\rightarrow m\quad\text{uniformly in}\,C^{0}\,\text{norm},$
$\Omega_{k}^{2}h_{k}\rightarrow\Omega^{2}h\quad\text{uniformly in}\,C^{0}\,\text{norm},$
$\underline{h}_{k}\rightarrow\underline{h}\quad\text{uniformly in}\,C^{0}\,\text{norm}.$
Meanwhile, we have already known that $r_{k}\rightarrow r$ on this compact domain, hence
$\partial_{u}r=\underline{h},\quad\partial_{\underline{u}}r=\Omega^{2}h.$
The fact that $\mu_{k}$ is bounded entails that $m_{k}=O(r_{k-1})$ and $m=O(r)$. We then have for the sake of uniqueness,
$\mu=\frac{m}{r}.$
Along with the formula of $\mu$, $m$ and $r$, the relation $h_{k}=\frac{\mu_{k}-1}{\underline{h}_{k}}$ gives rise to an identity
$h=\frac{\mu-1}{\underline{h}}.$
As a consequence, it follows that
$\mu\in C^{0},\,\,\ln\Omega^{2}h\in C^{0},\,\,\ln(-\underline{h})\in C^{0},\,\,\ln h\in C^{0},\,\,\text{and}\,\,r\in C^{1}.$
In addition, (A) suggests
$\int_{C_{u}}\Omega^{2}hr^{-1}\left(\frac{\left(\hat{w}_{k+1}-\hat{w}_{k}\right)^{2}}{r^{2}}+\left(\partial_{r}\left(\hat{w}_{k+1}-\hat{w}_{k}\right)\right)^{2}+\left(r\partial^{2}_{r}\left(\hat{w}_{k+1}-\hat{w}_{k}\right)\right)^{2}\right)\mathrm{d}\underline{u}$
$+\int_{\underline{C}_{\underline{u}}}h^{-1}r^{-1}\left(\frac{\left(\hat{w}_{k+1}-\hat{w}_{k}\right)^{2}}{r^{2}}+\left(\partial_{r}\left(\hat{w}_{k+1}-\hat{w}_{k}\right)\right)^{2}+h^{2}\left(\partial_{u}\left(\hat{w}_{k+1}-\hat{w}_{k}\right)\right)^{2}\right)\mathrm{d}u$
[…]
Hence, there exists $\hat{w}\in C^{0}$ such that
$\hat{w}_{k}\rightarrow\hat{w}\quad\text{uniformly in}\,H^{2}\,\text{norm},$
and
$\frac{\hat{w}}{r},\,\partial_{r}\hat{w}\in C^{0}.$
Setting
$w:=\hat{w}+1,$
we conclude $(r,\,m,\,w)$ satisfies the SSEYM equations. However, since $\partial_{u}\partial_{r}w$ is not necessarily bounded on the axis, the Yang–Mills equation (and also the whole SSEYM system) holds in the sense of distribution.
Continuous dependence on data. Suppose $\hat{w}_{n}$ and $\hat{w}_{m}$ are solutions with data $\hat{w}_{0n}$, $\hat{w}_{0m}$ respectively. The following equation holds
$\partial_{u}\partial_{\underline{u}}(\hat{w}_{n}-\hat{w}_{m})+2\frac{\Omega_{n}^{2}}{r_{n}^{2}}(\hat{w}_{n}-\hat{w}_{m})=f\left(\frac{\hat{w}_{n}}{r_{n}},\,\frac{\hat{w}_{m}}{r_{m}}\right)\cdot\frac{(U_{n}-U_{m})}{r_{n}},$
where $U_{n}-U_{m}$ involves $\frac{r_{n}-r_{m}}{r_{n}}$, $h^{-1}_{n}-h^{-1}_{m}$, $\Omega^{2}_{n}h_{n}-\Omega^{2}_{m}h_{m}$ and $r^{-1}_{n}\hat{w}_{n}-r^{-1}_{m}\hat{w}_{m}$, and
$\big{|}f\left(\frac{\hat{w}_{n}}{r_{n}},\,\frac{\hat{w}_{m}}{r_{m}}\right)\big{|}\lesssim\big{|}\frac{\hat{w}_{n}}{r_{n}}\big{|}+\big{|}\frac{\hat{w}_{m}}{r_{m}}\big{|}.$
Following the procedures in proving uniform bounds, we can show that property ii) holds.
In the end, the uniqueness follows in a similar way.
Appendix B An extension principle away from the axis
The extension principle away from the axis had been confirmed for systems, such as Einstein-Vlasov and Einstein charged scalar field in spherical symmetry [14, 15]. A prior to an extension theorem, a local existence theorem away from the axis is needed and it was addressed in [14] as well. The proof in [14] was based on the fixed point method and the double null gauge. This scheme indeed works well when the Einstein equations are coupled with other matter fields. In this section, we follow the approaches in [14, 15] to develop the local theory associated to the SSEYM system.
We rewrite the SSEYM system (2.4a)–(2.4e) and (2.5) in terms of $(r,\Omega,w)$ as follows,
$r\partial_{u}\partial_{\underline{u}}r={}$ $-\Omega^{2}-\partial_{u}r\partial_{\underline{u}}r+\Omega^{2}r^{-2}Q^{2},$ (B.1a)
$r^{2}\partial_{u}\partial_{\underline{u}}\ln\Omega={}$ $\Omega^{2}+\partial_{u}r\partial_{\underline{u}}r-2\Omega^{2}r^{-2}Q^{2},$ (B.1b)
$\partial_{u}\left(\Omega^{-2}\partial_{u}r\right)={}$ $-2\Omega^{-2}r^{-2}\left(\partial_{u}w\right)^{2},$ (B.1c)
$\partial_{\underline{u}}\left(\Omega^{-2}\partial_{\underline{u}}r\right)={}$ $-2\Omega^{-2}r^{-2}\left(\partial_{\underline{u}}w\right)^{2},$ (B.1d)
$\partial_{u}\partial_{\underline{u}}w+{}$ $\Omega^{2}r^{-2}wQ=0,$ (B.1e)
where $Q=w^{2}-1$. In view of the structures of the system (B.1a)-(B.1e), a local existence theorem away from the axis follows in a standard way [14].
Theorem B.1.
Let $k\geq 0$ and $\Sigma:=[0,d]\times\{0\}\cup\{0\}\times[0,d]$. On $\Sigma$, let $r,\Omega$ be positive functions such that $r\in C^{k+2}(\Sigma)$ and $\Omega\in C^{k+1}(\Sigma)$, and let $w\in C^{k+1}(\Sigma)$. Suppose (B.1c)–(B.1d) hold initially on $[0,d]\times\{0\}$ and $\{0\}\times[0,d]$, respectively. Define
$N_{u}:={}$ $\sup_{[0,d]\times\{0\}}\{|\Omega|,\,|\Omega^{-1}|,\,|\partial_{u}\Omega|,\,|r|,\,|r^{-1}|,\,|\partial_{u}r|,\,|\partial_{u}^{2}r|,\,|w|,\,|\partial_{u}w|\},$
$N_{\underline{u}}:={}$ $\sup_{\{0\}\times[0,d]}\{|\Omega|,\,|\Omega^{-1}|,\,|\partial_{\underline{u}}\Omega|,\,|r|,\,|r^{-1}|,\,|\partial_{\underline{u}}r|,\,|\partial_{\underline{u}}^{2}r|,\,|w|,\,|\partial_{\underline{u}}w|\},$
$N:={}$ $\sup\{N_{u},\,N_{\underline{u}}\}.$
Then there exists a $\delta>0$ depending only on $N$ such that there are functions $(r,\Omega,w)\in C^{k+2}\times C^{k+1}\times C^{k+1}$ satisfying equations (B.1a)–(B.1e) in $[0,\delta^{\ast}]\times[0,\delta^{\ast}]$, where $\delta^{\ast}=\min\{d,\,\delta\}$, with the above prescribed data on $\Sigma$.
With respect to the above local theorem, we can prove a generalized extension principle away from the axis (i.e. “strongly tame” in the language of [15]) following [15].
Theorem B.2.
Let $\left(\mathcal{Q}^{+},g_{\mu\nu},w\right)$ be the maximal future development of smooth spherical symmetric initial data for the SSEYM system. For $p\in\overline{\mathcal{Q}^{+}}$ and $q\in I^{-}(p)\cap\mathcal{Q}^{+}/\{p\}$ such that if $\mathcal{D}=J^{+}(q)\cap J^{-}(p)/\{p\}\subset\mathcal{Q}^{+}$ has finite spacetime volume and there are constant $r_{0}$ and $R$ such that
$0<r_{0}\leq r(p^{\prime})\leq R<\infty,\quad p^{\prime}\in\mathcal{D},$
then $p\in\mathcal{Q}^{+}$.
The proof relies on the dominant energy condition and the special structure of Einstein Yang–Mills system. This structure manifests itself in that to bound all the variables, it suffices to bound
$\iint_{\mathcal{D}}r^{2}T_{u\underline{u}}\mathrm{d}u\mathrm{d}\underline{u}\quad\text{and}\quad\iint_{\mathcal{D}}\Omega^{2}T_{AB}g^{AB}\mathrm{d}u\mathrm{d}\underline{u},$
which in the SSEYM case takes the form of
$\iint_{\mathcal{D}}\Omega^{2}\frac{Q^{2}}{r^{2}}\mathrm{d}u\mathrm{d}\underline{u}.$
Nevertheless, this quantity can be bounded straightforwardly through integrating (B.1a), please refer to [15] for more details.
Theorem B.2 shows that the “first singularity” in the maximal future development of spherically symmetric asymptotically flat data for the SSEYM system must emanate from the axis.
Appendix C Density theorem
As explained in Remark 1.6 (or Remark 1.2), regarding to the energy spaces in the extension theorems, we consider their equivalence, the weighted Sobolev space $H^{\prime s}_{w}$ with finite $H^{\prime s}_{w}$ norm [^4] If we take $r_{\ast}=0$ and $s=2$, the $H^{\prime s}_{w}$ space reduces to the $H^{1}$ space (1.5). , where
$\|\phi\|_{H^{\prime s}_{w}}:=\sum_{k\leq s}\int_{\{r_{\ast}<r\}}\xi^{-3-\gamma}r^{-1-b}\left((\xi r\partial_{r})^{k}\phi\right)^{2}\mathrm{d}r,\quad s\in\mathbb{Z}_{\geq 0}.$
We denote
$\rho:=\xi^{-1}r^{-1},\quad\text{that is}\,\,\rho=(r-r_{\ast})^{-1}.$
If $r_{\ast}=0$, $\xi=1$, and
$\xi^{-3-\gamma}r^{-1-b}=\rho^{1+b};$
If $r_{\ast}>0$, there is
$\xi^{-3-\gamma}r^{-1-b}\sim\rho^{3+\gamma}+\rho^{1+b}.$
Thus the $H^{\prime s}_{w}$ space is equivalent to the Sobolev space $H^{s}_{w}$ which has finite $H^{s}_{w}$ norm with
$\|\phi\|^{2}_{H^{s}_{w}}:=\sum_{k\leq s}\int_{\{r_{\ast}<r\}}f(\rho)\left((\rho^{-1}\partial_{r})^{k}\phi\right)^{2}\mathrm{d}r.$
Here $f(\rho)=(\rho^{3+\gamma}+\rho^{1+b})$ if $r_{\ast}>0$, and $f(\rho)=\rho^{1+b}$ if $r_{\ast}=0$. One can refer to [22, Section 3.2.4.] for more details of the $H^{s}_{w}$ space. In the remainder of this section, we aim to establish the density theorem in the weighted Sobolev space $H^{s}_{w}$.
Theorem C.1.
Smooth functions with compact support in $(r_{\ast},+\infty)$ is dense in $H^{s}_{w}\left((r_{\ast},+\infty)\right)$, $s\in\mathbb{Z}_{\geq 0}$.
Remark C.2.
If we take $r_{\ast}=0$, Theorem C.1 amounts to stating the density theorem in the $H^{1}$ space (1.5).
Proof.
The domain $(r_{\ast},+\infty)$ has an open covering
$(r_{\ast},+\infty)=\cup_{i\geq 0}\mathcal{D}_{1i},$
where we define
$\mathcal{D}_{1i}:=\{x\in(r_{\ast},+\infty)|\quad 2^{i}<\rho(x)<2^{i+3}\},\quad i\geq 1,$
and
$\mathcal{D}_{10}:=\{x\in(r_{\ast},+\infty)|\quad\rho(x)<2^{3}\}.$
On the other hand, letting
$\mathcal{D}_{2i}:=\{x\in(r_{\ast},+\infty)|\quad 2^{-i+1}<\rho(x)<2^{-i+1+3}\},\quad i\geq 1,$
we have
$\mathcal{D}_{10}\subset\cup_{i=1}^{+\infty}\mathcal{D}_{2i}.$
Therefore, $(r_{\ast},+\infty)$ is also covered by
$(r_{\ast},+\infty)=\cup_{i\geq 1}\left(\mathcal{D}_{1i}\cup\mathcal{D}_{2i}\right).$
Suppose the open covering $\cup_{i\geq 1}\left(\mathcal{D}_{1i}\cup\mathcal{D}_{2i}\right)$ which is locally finite, is associated to the partition of unity $\{\psi_{1i},\,\psi_{2i}\}_{i\geq 1}$. Namely,
$0\leq\psi_{ki}\leq 1,\,\psi_{ki}\in C_{0}^{\infty}(\mathcal{D}_{ki}),\,\,k=1,2,$
$\sum_{i\geq 1}\psi_{1i}+\psi_{2i}=1,$
and for all $m\in\mathbb{Z}_{>0}$,
[…]
The existence of such partition of unity is proved as follows. Note that, $\mathcal{D}_{1i}$ and $\mathcal{D}_{2i}$ can be alternatively written as
$\mathcal{D}_{1i}:=\{r\in(r_{\ast},+\infty)|\quad 2^{-i-3}<r-r_{\ast}<2^{-i}\},$
and
$\mathcal{D}_{2i}:=\{r\in(r_{\ast},+\infty)|\quad 2^{i-1-3}<r-r_{\ast}<2^{i-1}\},$
for $i\geq 1$. Take a smooth function $\phi$ such that
$\text{supp}\,\phi\subset\left(2^{-3},\,1\right).$
Then we define
$\phi_{1i}:=\phi\left(2^{i}(r-r_{\ast})\right),\quad\phi_{2i}:=\phi\left(2^{-i+1}(r-r_{\ast})\right)$
and let
$\psi_{1i}=\frac{\phi_{1i}}{\sum_{i\geq 1}\phi_{1i}+\phi_{2i}},\quad\psi_{2i}=\frac{\phi_{2i}}{\sum_{i\geq 1}\phi_{1i}+\phi_{2i}}.$
Now $\{\psi_{1i},\,\psi_{2i}\}_{i\geq 1}$ is the required partition of unity.
We can approximate $f\in H^{s}_{w}$ by the sequence $\sum_{j=1}^{K}\left(\psi_{1j}+\psi_{2j}\right)f$, $K\geq 1$, so that
$\|f-\sum_{j=1}^{K}\left(\psi_{1j}+\psi_{2j}\right)f\|_{H^{s}_{w}}\rightarrow 0,\quad\text{if}\,\,K\rightarrow+\infty.$
Furthermore, each of $\psi_{1j}f,\,\psi_{2j}f$, $1\leq i\leq K$, which belongs to the standard Sobolev space $H^{s}(\mathcal{D}_{1j})$, $H^{s}(\mathcal{D}_{2j})$ respectively can be approximated by $C_{0}^{\infty}$ functions. We thus conclude $C_{0}^{\infty}\left((r_{\ast},\,+\infty)\right)$ is dense in $H^{s}_{w}\left((r_{\ast},\,+\infty)\right)$. ∎

### v2

@@
 Abstract.
-We establish an extension principle for the spherically symmetric Einstein Yang–Mills system (SSEYM) with $H^{1}$ data.
-Based on this result, we further prove an extension theorem for developments of weighted $H^{1}$ data.
-In particular, the weighted $H^{1}$ space allows Hölder continuous data.
-Therefore, our result is consistence with the conjecture that well-posedness holds for Hölder continuous data (with sufficiently many angular derivatives) in vacuum where the axis is included.
-In contrast to a massless scalar field, the purely magnetic Yang–Mills field in spherical symmetry satisfies a wave-type equation with singular potential.
-The proof of Christodoulou [10], based on an $L^{\infty}-L^{\infty}$ estimate, fails for the Yang–Mills scenario.
-Instead, we employ the $L^{2}$ method, which works for the (massless or massive) scalar matter field as well.
+We prove the local existence theorem and establish an extension principle for the spherically symmetric Einstein Yang–Mills system (SSEYM) with $H^{1}$ data.
+This in addition implies Cauchy stability for the system.
+In contrast to a massless scalar field, the purely magnetic Yang–Mills field in spherical symmetry satisfies a wave-type equation with a singular potential.
+As a consequence, the proof of Christodoulou [10], based on an $L^{\infty}-L^{\infty}$ estimate, fails in the Yang–Mills context.
+Instead, we employ an $L^{2}$-based method, which is valid for both massless and massive scalar matter fields as well.
 1. Introduction
-In a series of works [9, 10, 11, 13], Christodoulou proved the weak cosmic censorship conjecture [^1] Generic asymptotically flat initial data have a maximal future development possessing a complete future null infinity[12].
+In a series of works [9, 10, 11, 13], Christodoulou established a proof of the weak cosmic censorship conjecture [^1] Generic asymptotically flat initial data have a maximal future development possessing a complete future null infinity[12].
 for the spherically symmetric Einstein massless scalar field system.
-It was further demonstrated by Li-Liu [16, 17] and An [1] under gravitational perturbations without symmetries.
-Recently, An-Tan [2] gave a proof for the system of Einstein charged massless scalar field in spherical symmetry as well.
-The instability mechanic of naked singularity in [13] was found out in the topology of bounded variation (BV).
-From the other perspective, the local existence theory and further the small data global existence in the BV topology were both verified ahead [10].
-As shown in [10, 13], the proofs of these (local or global) existence or instability statements were all built on a sharp criterion for $C^{1}$ extension.
-Along the line of [10], extension criterions were verified as well in the context of Einstein charged massless scalar field and Einstein massless Vlasov system in spherical symmetry [2, 20].
-We are now concerned with the Einstein Yang–Mills system, which is significant for various reasons.
-From a physical perspective, it triggers a wealth of fascinated gravitational collapses, as seen, for instance, [3, 6, 5] for several numerical results.
-On the other hand, we will show in this paper that the mathematical techniques in the treatment of Einstein Yang–Mills system will be significantly different from those for gravity coupled with massless matter fields [10, 20, 2].
-In this paper, we initiate the study of extension principle for the Einstein Yang–Mills system in spherical symmetry.
-We will first consider the characteristic problem for the spherically symmetric Einstein Yang–Mills (SSEYM) system where initial data are prescribed on a future light cone emanating from vertex at the center of symmetry.
-In this setting, we demonstrate the local existence theorem for data with a low-regularity norm ($H^{1}$ initial data).
+This result was further demonstrated by Li-Liu [16, 17] and An [1] under gravitational perturbations without symmetries.
+More recently, An-Tan [2] also verified the weak cosmic censorship conjecture for the Einstein charged massless scalar field system under spherical symmetry.
+As a core step in the proof, Christodoulou [13] introduced an instability mechanism for naked singularities within the bounded variation (BV) topology.
+This framwork is well motivated, since both the local existence theory and the subsequent small-data global existence result for the spherically symmetric Einstein massless scalar field system had already been established in the BV topology [10].
+In detailed proofs, one observes that the arguments for both (local or global) existence and instability rely crucially on a sharp criterion for $C^{1}$ extension [10, 13].
+This indicates that an extension principle in a suitable topology serves as the foundational element for addressing the weak cosmic censorship conjecture.
+Following the methodology of [10], such extension criteria have also been established in the context of the Einstein charged massless scalar field and the Einstein massless Vlasov system under spherical symmetry[2, 20].
+We are now concerned with the Einstein Yang–Mills system, which is of interest for several reasons.
+From a physical perspective, it triggers a rich variety of fascinated gravitational collapses, as evidenced by numerous numerical studies (see, for instance, [3, 6, 5]).
+On the mathematical side, we demonstrate in this paper that the techniques required for the Einstein Yang–Mills system differ substantially from those employed in the study of gravity coupled with massless matter fields [10, 20, 2].
+In this paper, we initiate the study of weak cosmic censorship conjecture for the Einstein Yang–Mills system in spherical symmetry.
+We will first consider the characteristic problem for the spherically symmetric Einstein Yang–Mills (SSEYM) system where the initial data are prescribed on a future light cone emanating from vertex at the center of symmetry.
+In this setting, we demonstrate the local existence theorem for data with a low-regularity norm ($H^{1}$ norm).
 We also deduce an extension principle for developments of $H^{1}$ data.
-Our second result is an extension theorem in a weaker topology.
-We show that if the interior is vacuum (and hence Minkowski), the spacetime is extendable in the topology of weighted $H^{1}$ space.
-Following theses extension theorems, the Cauchy stabilities are justified for the Einstein Yang–Mills system under spherically symmetric perturbations, which are initially small with respect to the above norms respectively.
-The weighted $H^{1}$ space allows the class of Hölder continuous data, or more precisely, data that are smooth except at one point and Hölder continuous there.
-It was conjectured by Rodnianski–Shlapentokh-Rothman [21] that well-posedness holds for Hölder continuous data with sufficiently many angular derivatives in vacuum (without symmetries) where the axis is included.
-Our extension theorem thus supports this conjecture.
-Previously, in the context of spherically symmetric massless matter field [10, 2], this conjecture is true in BV topology, where the instabilities of naked singularities are stated.
-However, as will be discussed later, the method used in massless case fails for Einstein–Yang–Mills scenario and hence a well-posed result in BV topology is not expected.
-The new $L^{2}$-based method we propose in this paper has a closer nature to the vacuum context, and may be relevant for future progress in vacuum.
+Following this extension theorem, the Cauchy stability is justified for the Einstein Yang–Mills system under spherically symmetric perturbations, which are initially small with respect to the $H^{1}$ norms.
+As will be discussed later, the method used for the Einstein massless scalar field fails in the Einstein Yang–Mills scenario and hence an extension principle within the BV topology is not expected.
+The new $L^{2}$-based method we propose in this paper has a closer nature to the vacuum setting, and may be relevant for any future progress in vacuum.
+Indeed, our method is intimately tied to the Hölder topology that we are currently working in progress.
 1.1. Statement of the main results
@@
 As shown in their definitions, $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$ is the domain enclosed by the axis $\Gamma$ and the hypersurfaces $C_{u_{0}}$, $\underline{C}_{\underline{u}_{\ast}}$, while $\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast})$ is the domain bounded by the axis $\Gamma$ and the hypersurfaces $C_{u_{0}}$, $C_{u_{1}}$, $\underline{C}_{\underline{u}_{\ast}}$.
-We formulate the SSEYM system with the gauge group $SU(2)$ in Section 2.1, where we assume the spherically symmetric purely magnetic ansatz for the Yang–Mills potential (2.2).
-The characteristic problem involves prescribing a $C^{0}$ function $\partial_{r}w|_{C_{u_{0}}}$ on an initial outgoing cone $C_{u_{0}}$, where $w$ is related to the Yang–Mills potential (2.2).
+We formulate the SSEYM system with the gauge group $SU(2)$ in Section 2.1, where the spherically symmetric purely magnetic ansatz for the Yang–Mills potential (2.2) is assumed.
+The characteristic problem of the SSEYM system involves prescribing a $C^{0}$ function $\partial_{r}w|_{C_{u_{0}}}$ on an initial outgoing cone $C_{u_{0}}$ that emanates from vertex at $\Gamma$.
+Here $w$ is related to the Yang–Mills potential (2.2).
 Throughout this paper, the $C^{0}$ Yang–Mills solution always refers to $\partial_{r}w\in C^{0}$.
@@
 $\frac{\hat{w}}{r}=\frac{1}{r}\int_{0}^{r}\partial_{r}w\,\mathrm{d}r^{\prime}|_{C_{u}}.$
-It suggests choosing the initial datum $w|_{C_{u_{0}}}$ can be replaced by prescribing $\partial_{r}w|_{C_{u_{0}}}$.
-Let us begin our statement with the $H^{1}$ case.
-1.1.1. Statement for the $H^{1}$ case
-Fixing a constant
-$0\leq b<\frac{\sqrt{89}-9}{4},$ (1.4)
+This suggests that instead of specifying the initial datum $w|_{C_{u_{0}}}$, one may equivalently prescribe $\partial_{r}w|_{C_{u_{0}}}$.
 We define the $H^{1}$ norm for $\partial_{r}w$
-$\|\partial_{r}w\|_{H^{1}}:=\int_{C_{u}}\frac{1}{r^{1+b}}\left(\frac{\hat{w}^{2}}{r^{2}}+(\partial_{r}w)^{2}+(r\partial_{r}(\partial_{r}w))^{2}\right)\mathrm{d}r.$ (1.5)
+$\|\partial_{r}w\|_{H^{1}}:=\int_{C_{u}}\frac{1}{r}\left(\frac{\hat{w}^{2}}{r^{2}}+(\partial_{r}w)^{2}+(r\partial_{r}(\partial_{r}w))^{2}\right)\mathrm{d}r.$ (1.4)
 Remark 1.2.
 It is remarkable that the norm (1.4) is equivalent to
-$\int_{0}^{r}\frac{1}{r^{1+b}}\sum_{0\leq k\leq 2}\left((r\partial_{r})^{k}\frac{\hat{w}}{r}\right)^{2}\mathrm{d}r.$ (1.6)
+$\int_{0}^{r}\frac{1}{r}\sum_{0\leq k\leq 2}\left((r\partial_{r})^{k}\frac{\hat{w}}{r}\right)^{2}\mathrm{d}r.$ (1.5)
 If we choose a new parameter $s:=\ln r$, (1.5) is alternatively written as
-$\int_{0}^{r_{1}}\frac{1}{r^{1+b}}\sum_{0\leq k\leq 2}\left((r\partial_{r})^{k}\frac{\hat{w}}{r}\right)^{2}\mathrm{d}r=\int_{-\infty}^{\ln r_{1}}e^{-bs}\sum_{0\leq k\leq 2}\left(\partial_{s}^{k}\frac{\hat{w}}{r}\right)^{2}\mathrm{d}s.$ (1.7)
-When $b=0$, (1.7) is the standard $H^{2}$ norm for $\frac{\hat{w}}{r}$.
+$\int_{0}^{r_{1}}\frac{1}{r}\sum_{0\leq k\leq 2}\left((r\partial_{r})^{k}\frac{\hat{w}}{r}\right)^{2}\mathrm{d}r=\int_{-\infty}^{\ln r_{1}}\sum_{0\leq k\leq 2}\left(\partial_{s}^{k}\frac{\hat{w}}{r}\right)^{2}\mathrm{d}s.$ (1.6)
+The norm (1.6) is the standard $H^{2}$ norm for $\frac{\hat{w}}{r}$.
 We abuse the notation a little bit and call (1.6) the $H^{2}$ norm of $\frac{\hat{w}}{r}$, or rather $H^{1}$ norm of $\partial_{r}w$.
-With respect to this $H^{1}$ space, the density theorem is proved in Appendix C.
 Remark 1.3.
-As a further remark, the standard energy for the Yang–Mills field does not account for the definition of energy (1.5), even if $b=0$.
-Actually, the standard first order energy for the Yang–Mills field should be roughly in the form of $\int_{C_{u}}\left(\frac{\hat{w}^{2}}{r^{2}}+(\partial_{r}w)^{2}\right)\mathrm{d}r$, where the weight $\frac{1}{r^{1+b}}$ is absent.
-The motivation for (1.5) indeed comes from the right hand side of (2.14a), since after integrating (2.14a) in $\underline{u}$, we expect to bound $\frac{m}{r}$ via the $H^{1}$ (in fact $H^{0}$) norm of $\partial_{r}w$.
-We present the $H^{1}$-extension theorem. One may refer to Theorem 3.1 for a comprehensive statement.
+As a further remark, the standard energy for the Yang–Mills field does not align with the definition of energy given in (1.4).
+In fact, the conventional first-order energy for the Yang-Mills field takes the form
+$\int_{C_{u}}\left(\frac{\hat{w}^{2}}{r^{2}}+(\partial_{r}w)^{2}\right)\mathrm{d}r,$
+which lacks the weight factor $\frac{1}{r}$.
+The motivation for (1.4) stems from the right hand side of (2.14a): after integrating (2.14a) in $\underline{u}$, we aim to bound $\frac{m}{r}$ via the $H^{1}$ (in fact, $H^{0}$) norm of $\partial_{r}w$.
+We now state the $H^{1}$-extension theorem. A complete and detailed formulation is provided in Theorem 3.1.
 Theorem 1.4.
-Suppose we have the initial data $\partial_{r}w|_{C_{u_{0}}}$ with bounded $H^{1}$ norm, and an $H^{1}$ solution (that is, $\partial_{r}w\in H^{1}$) to the SSEYM system on $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$.
+Suppose we have initial data $\partial_{r}w|_{C_{u_{0}}}$ with bounded $H^{1}$ norm, and let $w$ be an $H^{1}$ solution (that is, $\partial_{r}w\in H^{1}$) to the SSEYM system on the domain $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$.
 There is a constant $\varepsilon>0$ such that if
 $\lim_{u\rightarrow\underline{u}_{\ast}}\sup_{\mathcal{D}(u;\,\underline{u}_{\ast})}\frac{m}{r}<\varepsilon^{2},$
-the solution extends as an $H^{1}$ solution on $\mathcal{D}(u_{0};\,\underline{u}_{1})$ for some $\underline{u}_{1}>\underline{u}_{\ast}$.
-Following the extension theorem, there is the Cauchy stability statement, for which a more precise version is given in Theorem 3.9.
+the solution extends as an $H^{1}$ solution on the domain $\mathcal{D}(u_{0};\,\underline{u}_{1})$ for some $\underline{u}_{1}>\underline{u}_{\ast}$.
+Following the extension theorem, we state a Cauchy stability result. A precise version is presented in Theorem 3.7.
 Theorem 1.5.
-Suppose the data $\partial_{r}w|_{C_{u_{0}}}$ admit bounded $H^{1}$ norm that is small enough.
-Then there is a global and unique solution on $\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{0})$ with small $H^{1}$ norm, for any $\underline{u}_{0}\in(u_{0},+\infty)$ and $u_{1}\in(u_{0},\,\underline{u}_{0}]$.
-We next proceed to the weighted $H^{1}$ case.
-1.1.2. Statement for the weighted $H^{1}$ case
-We will consider a class of weighted energies in the domain $\mathcal{D}(u_{0},u;\underline{u}_{\ast}+\delta)$ with $u\in[u_{0},\,\underline{u}_{\ast})$ and $\delta>0$.
-The weighted energies are defined with the assistance of a non-negative function
-$\xi:=\frac{r-r_{\ast}}{r}\quad\text{with}\quad r_{\ast}(u):=r(u,\,\underline{u}_{\ast}),$ (1.8)
-for $\underline{u}\geq\underline{u}_{\ast}$. We fix a constant $\gamma\in(b,2]$ with $b$ given in (1.4) and define the following weighted $H^{1}$ norm for $\partial_{r}w$
-$\int_{C_{u}\cap\{\underline{u}\geq\underline{u}_{\ast}\}}\frac{1}{r^{1+b}}\left(\xi^{-3-\gamma}\frac{\hat{w}^{2}}{r^{2}}+\xi^{-1-\gamma}(\partial_{r}w)^{2}+\xi^{1-\gamma}\left(r\partial_{r}(\partial_{r}w)\right)^{2}\right)\mathrm{d}r.$ (1.9)
-Remark 1.6.
-The weighted $H^{1}$ norm (1.9) is equivalent to
-$\int_{\{r_{\ast}\leq r\}}\xi^{-3-\gamma}r^{-1-b}\sum_{0\leq k\leq 2}\left((\xi r\partial_{r})^{k}\frac{\hat{w}}{r}\right)^{2}\mathrm{d}r|_{C_{u}}.$ (1.10)
-For the density theorem in the weighted $H^{1}$ space, please refer to Appendix C for a proof.
-Theorem 1.7.
-Given datum $\partial_{r}w|_{C_{u_{0}}}$ that is supported in $(\underline{u}_{\ast},\,+\infty)$ (namely, $\partial_{r}w|_{C_{u_{0}}\cap\{\underline{u}\leq\underline{u}_{\ast}\}}\equiv 0$), we assume $|r-r_{\ast}|^{-\frac{\gamma}{2}}\partial_{r}w|_{C_{u_{0}}}$ is continuous at $\underline{u}=\underline{u}_{\ast}$ and the weighted $H^{1}$ norm of $\partial_{r}w|_{C_{u_{0}}}$ on $C_{u_{0}}\cap\{\underline{u}_{\ast}\leq\underline{u}\leq\underline{u}_{\ast}+\delta\}$ for some $\delta>0$ is bounded.
-If $\delta$ is small enough, the solution extends as a solution with bounded, weighted $H^{1}$ norm in the domain of $\mathcal{D}(u_{0},\,u;\,\underline{u}_{\ast}+\delta)$, where $u\in[u_{0},\,\underline{u}_{\ast})$.
-Remark 1.8.
-We remark that our data requirement allows the initially second derivative $\partial^{2}_{r}w|_{C_{u_{0}}}$ blows up [^2] The $C^{1}$ case ($\gamma=2$) is also included. like
-$(r-r_{\ast})^{-1+\frac{\gamma}{2}},\quad 0\leq b<\gamma\leq 2.$
-Thus, the data in Theorem 1.7 are, in fact, Hölder continuous.
-Roughly speaking, Theorem 1.7 states that if the interior $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$ is Minkowski, and the data are Hölder continuous, then the solution can be extended beyond the last incoming cone $\underline{C}_{\underline{u}_{\ast}}$.
-Furthermore, we will show that the extended solution is still Hölder continuous when it goes across $\underline{C}_{\underline{u}_{\ast}}$.
-Please refer to Theorem 4.1 for a detailed description.
-The next one is a Cauchy stability statement in the weighted norm space. One may refer to Theorem 4.10 for more details.
-Theorem 1.9.
-Given datum $\partial_{r}w|_{C_{u_{0}}}$ that is supported in $(\underline{u}_{\ast},\,+\infty)$, suppose $|r-r_{\ast}|^{-\frac{\gamma}{2}}\partial_{r}w|_{C_{u_{0}}}$ is continuous at $\underline{u}=\underline{u}_{\ast}$, and the weighted $H^{1}$ norm on $C_{u_{0}}\cap\{\underline{u}_{\ast}\leq\underline{u}\leq\underline{u}_{1}\}$, $\underline{u}_{1}<+\infty$, is small enough.
-Then there is a unique solution with bounded, weighted $H^{1}$ norm on $\mathcal{D}(u_{0},\,u;\,\underline{u}_{1})$, where $u\in[u_{0},\,\underline{u}_{\ast})$.
+Suppose the data $\partial_{r}w|_{C_{u_{0}}}$ admit a bounded $H^{1}$ norm that is small enough.
+Then, for any $\underline{u}_{0}\in(u_{0},+\infty)$ and $u_{1}\in(u_{0},\,\underline{u}_{0}]$, there exists a unique and global solution in $\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{0})$ whose $H^{1}$ norm remains small.
 Remark 1.6.
 In the subsection followed, we will sketch the main idea for the proofs.
-We propose the $L^{2}$ method, which is tailed for the Yang–Mills field, or rather, the wave type equation with a positive, singular potential.
-Indeed we can also adapt the $L^{2}$ scheme to a massless or massive scalar field so that all the preceding statements (Theorem 1.41.9) hold for the Einstein (massive or massless) scalar field system in spherical symmetry.
-Since our proof is independ of $C^{1}$ norm, but only depends on $H^{1}$ norm, we have in fact achieved an extension principle in a weaker topology ($H^{1}$ topology) than that in Christodoulou [10] ($C^{1}$ topology), and thus less regularity on the axis is obtained as well.
-To facilitate the extension theorems, we verify the associated local existence theory for the characteristic problem of SSEYM system in Appendix A.
-We also state an extension principle away from the axis in Appendix B for completeness.
-Thus, it confirms that the “first singularity” for the SSEYM system must emanate from the axis.
+We introduce the $L^{2}$-based method, tailored to the Yang–Mills field–or more precisely, to wave type equations with a positive, singular potential.
+This approach can also be adapted to both massless and massive scalar fields, ensuring that all the preceding statements (Theorems 1.41.5) remain valid for the spherically symmetric Einstein (massive or massless) scalar field system.
+Since our proof is independ of $C^{1}$ bounds, but only depends on $H^{1}$ estimates, we effectively establish an extension principle in a weaker topology–the $H^{1}$ topology, compared to the $C^{1}$ framework used by Christodoulou [10].
+Consequently, our results also require less regularity along the axis.
+To facilitate the extension theorems, we establish a local existence theory for the characteristic problem of the SSEYM system in Section 4.
+Additionally, for completeness, we present extension principle away from the axis in Appendix A.
+Thus, it confirms that the “first singularity” for the SSEYM system must orginate from the axis.
 1.2. Comments on the proof
-The purely magnetic Yang–Mills field obeys a wave-type equation with a singular potential $\frac{1}{r}$, which makes itself significantly different from a standard massless scalar field.
-To highlight the difference, it is useful to compare their formulations in Minkowski space.
-Notice that, the massless sclar field equation in Minkowski takes the form of
+The purely magnetic Yang–Mills field obeys a wave-type equation with a singular potential proportional to $\frac{1}{r}$, which distinguishes it fundamentally from a standard massless scalar field.
+To highlight the difference, it is instructive to compare their formulations in Minkowski spacetime.
+The massless scalar field equation in Minkowski takes the form
 $\partial_{u}\partial_{\underline{u}}(r\phi)=0,$
-while the Yang–Mills equation reads
+while the Yang–Mills equation is given by
 $\partial_{u}\partial_{\underline{u}}w+\frac{w(w^{2}-1)}{r^{2}}=0.$
-In terms of $\hat{w}$ (1.3), the Yang–Mills equation can be reformulated as follows
+In terms of the variable $\hat{w}$ defined in (1.3), the Yang–Mills equation can be reformulated as
 $\partial_{u}\partial_{\underline{u}}\hat{w}+\frac{2}{r}\frac{\hat{w}}{r}=\frac{\hat{w}^{2}}{r^{2}}\left(\hat{w}+3\right).$
-It manifests itself as a nonlinear wave equation with an additional linear, singular term
+This exhibits a nonlinear wave structure with an additional linear, singular term
 $\frac{2}{r}\frac{\hat{w}}{r},$
-compared to the scalar wave equation. As a consequence, the techniques for proving the extension principles are necessarily different.
+absent in the scalar wave case. As a result, the methods required to establish extension principles in the two settings necessarily differ.
 1.2.1. The $H^{1}$ case
@@
 $\partial_{u}(\partial_{r}\alpha_{s})={}$ $-2\frac{\partial_{u}r}{r}\frac{\frac{2m}{r}}{1-\frac{2m}{r}}\partial_{r}\alpha_{s}+3\frac{\partial_{u}r}{r^{2}}\frac{\frac{2m}{r}}{1-\frac{2m}{r}}r\partial_{r}\phi-\frac{\partial_{u}r}{r^{2}}\frac{1}{1-\frac{2m}{r}}(r\partial_{r}\phi)^{3}.$
-Remarkably, this equation contains no linear terms on the right hand side.
-Namely, terms that are linear on $\partial_{r}\alpha_{s}$, $\alpha_{s}$, $r\partial_{r}\phi$ and $\frac{2m}{r}$ are absent.
-It is exactly this feature that allows an usage of $L^{\infty}-L^{\infty}$ estimate in proving the extension theorem [10, Theorem 5.1].
-Without loss of generality, we take the first term $-2\frac{\partial_{u}r}{r}\frac{\frac{2m}{r}}{1-\frac{2m}{r}}\partial_{r}\alpha_{s}$ for instance to explain the main ingredient of this approach.
-In this quadratic term, the $\frac{2m}{r}$ offers smallness $\varepsilon$ so that the whole term is negligible and a type of inequality
+Remarkably, the right-hand side of this equation contains no linear terms in $\partial_{r}\alpha_{s}$, $\alpha_{s}$, $r\partial_{r}\phi$ and $\frac{2m}{r}$.
+It is exactly this structure that enables a usage of an $L^{\infty}-L^{\infty}$ estimate in the proof of the extension theorem [10, Theorem 5.1].
+To illustrate the key idea of this method, consider, without loss of generality, the term
+$2\frac{\partial_{u}r}{r}\frac{\frac{2m}{r}}{1-\frac{2m}{r}}\partial_{r}\alpha_{s}.$
+Its treatment exemplifies how nonlinear contributions are controlled within the $L^{\infty}$ framework.
+In this quadratic term, the $\frac{2m}{r}$ offers a smallness factor $\varepsilon$, rendering the entire term negligible.
+This allows an inequality of the form
 $\sup|\partial_{r}\alpha_{s}|\leq\text{data}+\varepsilon\sup|\partial_{r}\alpha_{s}|$
-yields the bound for $\partial_{r}\alpha_{s}$.
-In principle, this kind of $L^{\infty}-L^{\infty}$ estimate works well for ”generic” massless matter fields that share the common feature of vanishing linear terms.
-Along with the massless scalar field, as far as we know, the $L^{\infty}-L^{\infty}$ strategy truns out to be valid as well for the charged massless scalar field [2] and the massless Vlasov system [20].
-In contrast, for the Yang–Mills scenario, due to the additional potential, the main equation for $\partial_{r}\alpha$ (2.3), $\alpha:=\partial_{r}w$, involves linear terms that are linear on $\frac{\hat{w}}{r}$ and $\alpha$ respectively on the right hand side.
-Therefore, the strategy of $L^{\infty}-L^{\infty}$ estimate fails in the Yang–Mills case, since we are unable to gain extra smallness from these linear terms.
-Nevertheless, noticing the positive sign of the potential in the Yang–Mills equation, we propose the $L^{2}$ method, which allows us to treat the quadratic terms originated from the linear terms through integration by parts.
-After that, terms that have no additional smallness give rise to quadratic terms with definite signs.
-Since our energy is not the standard energy, it is in fact $r^{-1-b}$ weighted, these quadratic terms do not necessarily admit favorable signs.
-Specifically, we are unable to derive any energy bound using the first-order energy identities; see (3.1.4) and (3.1.4).
-To proceed, the idea is, after performing integration by parts on these linear-related terms, the second order energy identity generates positive or negative quadratic terms that are potential to compensate the corresponding terms with opposite signs in the first order energy identities, see (3.1.4).
-Then we play with these two energy identities by assigning different weights on them to render all the quadratic spacetime integrals positive.
-Eventually, we obtain the $H^{1}$ energy bounds, see Section 3.1.4, and the pointwise estimates follow from a Sobolev type inequality in the $H^{1}$ space.
-The scheme employed for the Yang–Mills case applies to the (massless or massive) scalar field so that an extension principle in a low regularity ($H^{1}$) can be established.
-Moreover, using our method, we establish the extension principle with less regularity on the axis: $r\partial_{r}\phi\sim r^{b}$ with $0\leq b<\frac{\sqrt{89}-9}{4}$, while in Christodoulou [10], $r\partial_{r}\phi$ is roughly $r$ near the axis.
-Besides, the $H^{1}$ energy estimate is crucial to promote a local existence theorem, Theorem A.1.
-1.2.2. The weighted $H^{1}$ case
-The proof in the weighted $H^{1}$ case heavily depends on the energy identities derived in the previous $H^{1}$ case and its modification through introducing a weight function $\xi$ (1.8).
-Remarkably, the weight function $\xi$ is designed in such a way that in the domain concerned,
-$0\leq\xi<1,\quad\partial_{u}\xi\geq 0,\quad\partial_{\underline{u}}\xi\geq 0$
-More specifically, noting the negative power on $\xi$ in the first order weighted energy, the sign of $\partial_{u}\xi$ and $\partial_{\underline{u}}\xi$ helps to produce positive spacetime integrals which will cancel the negative terms in the originally unweighted energy identities.
-Once the first order estimate is completed, the second order one follows.
+to yield a bound for $\partial_{r}\alpha_{s}$.
+In principle, such $L^{\infty}-L^{\infty}$ estimates are effective for “generic” massless matter fields that share the characteristic absence of linear terms.
+Besides the massless scalar field, this $L^{\infty}-L^{\infty}$ approach has also been shown to apply to the charged massless scalar field [2] and the massless Vlasov system [20], to the best of our knowledge.
+In contrast, for the Yang–Mills case, the presence of an additional potential leads to linear terms (linear in $\frac{\hat{w}}{r}$ and $\alpha$) in the main equation for $\partial_{r}\alpha$ (2.3), $\alpha:=\partial_{r}w$.
+Consequently, the $L^{\infty}-L^{\infty}$ estimate strategy fails here, as these linear terms do not provide the required smallness.
+Nevertheless, the positive sign of the potential in the Yang-Mills equation motivates the use of an $L^{2}$-based method.
+This approach allows us to handle quadratic terms arising from the linear contributions via integration by parts.
+After this manipulation, terms lacking inherent smallness give rise to quadratic expressions with definite signs.
+However, since our energy is weighted by $r^{-1}$ (and thus non-standard), these quadratic terms do not automatically exhibit favorable signs.
+In particular, the first-order energy identities (3.5) and (3.5) do not directly yield an energy bound.
+For this issue, we proceed to the second-order energy estimates and integrating by parts on the linear-derived terms.
+The second-order energy identity then produces quadratic terms with definite signs, which can potentially compensate for corresponding terms of opposite sign in the first-order identities (see (3.5)).
+By strategically weighting these two energy identities, we ensure that all quadratic spacetime integrals become positive.
+This ultimately yields $H^{1}$ energy bounds (as detailed in Section 3.5), and pointwise estimates follow via a Sobolev-type inequality in the $H^{1}$ space.
+The method developed for the Yang-Mills case can also be applied to both massless and massive scalar fields, enabling the establishment of an extension principle in a low-regularity ($H^{1}$) setting.
+Furthermore, our approach allows less regularity near the axis: we only require $r\partial_{r}\phi\sim 1$, while in Christodoulou [10], the setting $r\partial_{r}\phi\sim r$ near the axis is stronger.
+Additionally, the $H^{1}$ energy estimates play a crucial role in establishing the local existence result given in Theorem 4.1.
 1.3. Outline of this paper
@@
 Section 3 is devoted to establishing the extension theorem and Cauchy stability in $H^{1}$ space.
-The counterparts in the weighted $H^{1}$ space are proved in Section 4.
-We also demonstrate the local existence theorem in Appendix A, the extension principle away from the axis in Appendix B, and provide a density theorem in Appendix C.
+The proof of local existence theorem is presented in Section 4, while an extension principle away from the axis is provided in Appendix A.
 1.4. Acknowledgement
-J.L. is supported by National Key R&D Program of China (No. 2022YFA1005400) and NSFC (12326602, 12141106). J.W. is supported by NSFC (Grant No. 12271450).
+J.L. is supported by National Key R&D Program of China (No. 2022YFA1005400) and NSFC (No. 12326602, 12141106). J.W. is supported by NSFC (No. 12271450).
 2. Preliminaries
 2.1. Spherically symmetric Einstein Yang–Mills equations
-In this section, we shall formulate the SSEYM system with certain gauge assumptions.
-On a spherical symmetric spacetime $(\mathcal{M},g)$, there exists an isometric action of the rotational group $SO(3)$, where the orbits are $2$-dimensional spheres.
-Define the axis $\Gamma$ of $(\mathcal{M},g)$ to be the set of fixed points of the $SO(3)$ action.
-Let us introduce the double null coordinates $(u,\underline{u})$, and denote the level sets by $C_{u}$ and $\underline{C}_{\underline{u}}$, which are the incoming and outgoing null cones.
-The intersection of $C_{u}$ and $\underline{C}_{\underline{u}}$ is a $2$-sphere denoted by $S_{\underline{u},u}$.
-We define the area function $r$ by
+In this section, we formulate the SSEYM system with specific gauge assumption.
+On a spherical symmetric spacetime $(\mathcal{M},g)$, there exists an isometric action of the rotation group $SO(3)$, whose orbits are two-dimensional spheres.
+Define the axis $\Gamma$ of $(\mathcal{M},g)$ as the set of fixed points of the $SO(3)$ action.
+We introduce the double null coordinates $(u,\underline{u})$, and denote the level sets by $C_{u}$ and $\underline{C}_{\underline{u}}$, representing the incoming and outgoing null cones respectively.
+The intersection $S_{\underline{u},u}=C_{u}\cap\underline{C}_{\underline{u}}$ is a two-dimensional sphere.
+The area function $r$ is defined by
 $\text{Area}(S_{\underline{u},u})=4\pi r^{2}.$
-Then the axis is characterized by
+The axis is characterized by
 $r=0\quad\text{on}\,\,\Gamma.$
-The metric on $\mathcal{M}$ takes the form of
+The metric on $\mathcal{M}$ takes the form
 $g=-4\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}+r^{2}\mathrm{d}\sigma_{S^{2}},$
-where $\mathrm{d}\sigma_{S^{2}}$ is the standard metric on $S^{2}$ and $\Omega$ is the lapse function such that $-2\Omega^{2}=g(\partial_{u},\partial_{\underline{u}})$.
+where $\mathrm{d}\sigma_{S^{2}}$ is the standard metric on $S^{2}$ and the lapse function $\Omega$ satisfies $2\Omega^{2}=g(\partial_{u},\partial_{\underline{u}})$.
 We define
@@
 $\tau_{1}=\frac{i}{2}\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\,\,\tau_{2}=\frac{1}{2}\begin{pmatrix}0&-1\\ 1&0\end{pmatrix},\,\,\tau_{3}=\frac{i}{2}\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$,
-with $[\tau_{i},\,\tau_{j}]=\epsilon_{ijk}\tau_{k}$, $\epsilon_{ijk}$ being the Levi-Civita symbol, $\{i,j,k\}\subset\{1,2,3\}$.
+with $\tau_{i},\,\tau_{j}]=\epsilon_{ijk}\tau_{k}$, $\epsilon_{ijk}$ being the Levi-Civita symbol, $\{i,j,k\}\subset\{1,2,3\}$.
 We assume the spherically symmetric purely magnetic ansatz [4, 6] for the Yang–Mills potential,
@@
 $F={}$ $\mathrm{d}A+A\wedge A$
-[…]
-As followed from the ansatz (2.2), $F_{u\underline{u}}=0$, namely the electric charge vanishes; $F_{AB}=\frac{w^{2}-1}{r^{2}}\tau_{3}\varepsilon_{AB}$ [^3] The capital Latin letters $\{A,B\}$ refer to indices on $S^{2}$.
-and we define the magnetic charge within the sphere $S_{\underline{u},u}$ to be
+$={}$ $\mathrm{d}w\wedge\left(\tau_{1}\mathrm{d}\theta+\sin\theta\tau_{2}\mathrm{d}\phi\right)-\left(1-w^{2}\right)\tau_{3}\sin\theta\mathrm{d}\theta\wedge\mathrm{d}\phi.$
+As followed from the ansatz (2.2), $F_{u\underline{u}}=0$, indicating that the electric charge vanishes; The angular components are given by
+$F_{AB}=\frac{w^{2}-1}{r^{2}}\tau_{3}\varepsilon_{AB},$
+where capital Latin letters $\{A,B\}$ refer to indices on the 2-sphere $S^{2}$. We define the magnetic charge within the sphere $S_{\underline{u},u}$ as
 $Q:=w^{2}-1.$ (2.3)
-The vacuum of the Yang–Mills field are given by $w=\pm 1$.
-The Einstein Yang–Mills system takes the form of
+The vacua of the Yang–Mills field correspond to the values $w=\pm 1$.
+The Einstein Yang–Mills system takes the form
 $R_{\alpha\beta}-\frac{1}{2}Rg_{\alpha\beta}={}$ $2T_{\alpha\beta},$
 $\hat{D}^{\mu}F_{\mu\nu}={}$ $0,$
-[…]
+where $\hat{D}_{\mu}=D_{\mu}+[A_{\mu},\cdot\,]$ denotes the gauge covariant derivative, $\cdot$$\cdot\,$] is the Lie bracket on $su(2)$, and the energy momentum tensor of Yang–Mills field is given by
 $T_{\alpha\beta}=\left\langle F_{\alpha}{}^{\nu},\,F_{\beta\nu}\right\rangle-\frac{1}{4}g_{\alpha\beta}\left\langle F_{\sigma\rho},\,F^{\sigma\rho}\right\rangle.$
 For more detailed descriptions for the Einstein Yang–Mills system, please refer to [7] and [18, 19].
-We are now ready to write down the spherical symmetric Einstein Yang–Mills system (SSEYM).
-With the ansatz (2.2) and the double gauge, the SSEYM system consistes of the following equations for $(r,\Omega,w)$,
+We now present the spherical symmetric Einstein Yang–Mills system (SSEYM).
+With the ansatz (2.2) and the double gauge, the SSEYM system reduces to the following equations for the variables $(r,\Omega,w)$,
 $\partial_{\underline{u}}h={}$ $-2\Omega^{-2}\frac{(\partial_{\underline{u}}w)^{2}}{r},$ (2.4a)
@@
 $\partial_{u}m={}$ $\underline{h}\frac{Q^{2}}{2r^{2}}-h(\partial_{u}w)^{2}.$ (2.6b)
-A characteristic initial data set for the SSEYM system composes of a triplet $(r,\Omega,\partial_{r}w)|_{C_{u_{0}}}$ defined on an outgoing cone $C_{u_{0}}$, see Remark 1.1.
-In addition to the initial condition, the SSEYM system is incorporated with the following boundary condition
+A characteristic initial data set for the SSEYM system consists of a triplet $(r,\Omega,\partial_{r}w)|_{C_{u_{0}}}$ prescribed on an outgoing cone $C_{u_{0}}$, as illustrated in Remark 1.1.
+In addition to the initial condition, the SSEYM system is incorporated with the following boundary condition on the axis $\Gamma$
 $r|_{\Gamma}=0,\quad m|_{\Gamma}=0,\quad Q|_{\Gamma}=0.$ (2.7)
-Note that, $Q|_{\Gamma}=0$ ensures the regularity of (2.5) on the axis.
-We can reformulate the SSEYM system by eliminating $\Omega$ and replacing it by $m$, so that it is equivalent to the system of $(r,m,w)$ which is composed of
+The condition $Q|_{\Gamma}=0$ is essential to ensure the regularity of the Yang–Mills equation (2.5) on the axis.
+We can reformulate the SSEYM system by eliminating $\Omega$ and introducing the mass function $m$. The resulting equivalent system for the variables $(r,m,w)$ comprises the equations
 […]
-with the boundary condition (2.7).
+together with the boundary condition (2.7).
 2.2. Gauge choice
-We fix $\underline{u}$ on the initial outgoing cone $C_{u_{0}}$ by requiring that $\underline{u}=2r+u_{0}$ on $C_{u_{0}}$, and it suggests
+We fix the coordinate $\underline{u}$ on the initial outgoing cone $C_{u_{0}}$ by requiring that $\underline{u}=2r+u_{0}$ on $C_{u_{0}}$, which suggests the gauge condition
 $\Omega^{2}h|_{C_{u_{0}}}=\frac{1}{2}.$ (2.8)
@@
 $u=\underline{u}\quad\text{on}\,\,\Gamma.$
-As a consequence, in view of $r=0$ on $\Gamma$, it follows that
+Since $r=0$ on $\Gamma$, it follows that
 $\Omega^{2}h=-\underline{h}\quad\text{on}\,\,\Gamma.$ (2.9)
@@
 The transport equations for $m$ (2.6a)–(2.6b) help to deduce the subsequent equations for $\mu$,
-$\partial_{\underline{u}}\mu={}$ $-\Omega^{2}h\frac{\mu}{r}+\Omega^{2}h\frac{\Phi^{2}}{2r}+\left(1-\mu\right)\Omega^{2}h\frac{\alpha^{2}}{r},$ (2.14a)
-$\partial_{u}\mu={}$ $-\underline{h}\frac{\mu}{r}+\underline{h}\frac{\Phi^{2}}{2r}-h\frac{(\partial_{u}w)^{2}}{r}.$ (2.14b)
+$\partial_{\underline{u}}\mu={}$ $-\Omega^{2}h\frac{\mu}{r}+\Omega^{2}h\frac{\Phi^{2}}{r}+2\left(1-\mu\right)\Omega^{2}h\frac{\alpha^{2}}{r},$ (2.14a)
+$\partial_{u}\mu={}$ $-\underline{h}\frac{\mu}{r}+\underline{h}\frac{\Phi^{2}}{r}-2h\frac{(\partial_{u}w)^{2}}{r}.$ (2.14b)
 Due to the Yang–Mills equation (2.5), we have
@@
 the outgoing null vector parametrized by $r$.
-3. The $H^{1}$ case
-We aim to establish an extension theorem for the solution to the SSEYM system in the class of $\left(r,\,m,\,\alpha\right)\in C^{1}\times C^{1}\times H^{1}$.
-Recall (1.5) for the definition of the $H^{1}$ norm, where the fixed constant $b$ is given in (1.4).
-3.1. Extension theorem in $H^{1}$ space
+3. Extension theorem in $H^{1}$ space
+We aim to establish an extension theorem for solutions to the SSEYM system in the class of $\left(r,\,m,\,\alpha\right)\in C^{1}\times C^{0}\times H^{1}$.
+Recall (1.4) for the definition of $H^{1}$ norm.
 For any $u_{0}<\underline{u}_{\ast}\in\mathbb{R}$, the domain $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$ bounded by the axis $\Gamma$, $C_{u_{0}}$ and $\underline{C}_{\underline{u}_{\ast}}$, is defined in (1.1).
 Theorem 3.1.
-Given initial data $\alpha|_{C_{u_{0}}}$, assume the initial energy
-$\int_{C_{u_{0}}\cap\{u_{0}\leq\underline{u}\leq\underline{u}_{0}\}}\frac{1}{r^{1+b}}\left(\alpha^{2}+(\alpha^{\prime})^{2}+r^{-2}\hat{w}^{2}\right)\mathrm{d}r,\quad\text{for any}\,\,\,\underline{u}_{0}>u_{0}$
-is finite.
-Suppose we have in the domain $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$ a solution to the SSEYM system $(r,\,m,\,\alpha)\in C^{1}\times C^{1}\times H^{1}$ (if $b>0$) or $(r,\,m,\,\alpha)\in C^{1}\times C^{0}\times H^{1}$ (if $b=0$).
+Given initial data $\alpha|_{C_{u_{0}}}$, with finite initial energy
+$\int_{C_{u_{0}}\cap\{u_{0}\leq\underline{u}\leq\underline{u}_{0}\}}\frac{1}{r}\left(\alpha^{2}+(\alpha^{\prime})^{2}+r^{-2}\hat{w}^{2}\right)\mathrm{d}r,\quad\text{for any}\,\,\,\underline{u}_{0}>u_{0},$
+suppose we have on the domain $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$ a solution to the SSEYM system $(r,\,m,\,\alpha)\in C^{1}\times C^{0}\times H^{1}$.
 There is a constant $\varepsilon>0$ such that if
 $\lim_{u\rightarrow\underline{u}_{\ast}}\sup_{\mathcal{D}(u;\,\underline{u}_{\ast})}\mu<\varepsilon^{2},$
-then the following conclusions hold.
-Case I: $0<b<\frac{\sqrt{89}-9}{4}$.
-The solution $(r,\,m,\,\alpha)$ extends as a $C^{1}\times C^{1}\times H^{1}$ solution on $\mathcal{D}(u_{0};\,\underline{u}_{1})$, for some $\underline{u}_{1}>\underline{u}_{\ast}$.
-Moreover, the solution obeys the estimates $\mathcal{D}(u_{0};\,\underline{u}_{1})$
-$|r^{-1}\hat{w}|+|\alpha|+|\partial_{u}w|\lesssim{}$ $r^{b},$ (3.1)
-$|r^{-1}\hat{w}|+|\alpha|\lesssim{}$ $\varepsilon^{2},$ (3.2)
-and the energy bound holds
-$\int_{C_{u}\cap\mathcal{D}(u_{0};\,\underline{u}_{1})}\frac{1}{r^{1+b}}\left(\Phi^{2}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}r$
-$+\int_{\underline{C}_{\underline{u}}\cap\mathcal{D}(u_{0};\,\underline{u}_{1})}\frac{h^{-1}}{r^{1+b}}\left(\Phi^{2}+\alpha^{2}+h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u<C.$ (3.3)
-Case II: $b=0$.
-There exists $\underline{u}_{1}>\underline{u}_{\ast}$ such that the solution extends to $\mathcal{D}(u_{0};\,\underline{u}_{1})$ in the class of $(r,m,\alpha)\in C^{1}\times C^{0}\times H^{1}$, and the estimates (3.2)–(3.1) hold with $b=0$.
+then the solution extends to $\mathcal{D}(u_{0};\,\underline{u}_{1})$, for some $\underline{u}_{1}>\underline{u}_{\ast}$, with the same regularity $C^{1}\times C^{0}\times H^{1}$.
+Moreover, the solution obeys the estimates on $\mathcal{D}(u_{0};\,\underline{u}_{1})$
+$|r^{-1}\hat{w}|^{2}\lesssim\varepsilon^{2},\quad\text{and}\quad|\alpha|^{2}\lesssim\varepsilon,$ (3.1)
+and the energies remain bounded
+$\int_{C_{u}\cap\mathcal{D}(u_{0};\,\underline{u}_{1})}\frac{1}{r}\left(\Phi^{2}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}r$
+$+\int_{\underline{C}_{\underline{u}}\cap\mathcal{D}(u_{0};\,\underline{u}_{1})}\frac{h^{-1}}{r}\left(\Phi^{2}+\alpha^{2}+h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u<C.$ (3.2)
 In the rest of this section, we will prove Theorem 3.1.
+3.1. Conventions
 Let us introduce some conventions for this section.
-For any $u_{0}<u_{1}\leq\underline{u}_{\ast}$, we recall the definition of the domain $\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast})$ (1.2) that is bounded by $\Gamma$ and the hypersurfaces $C_{u_{0}}$, $C_{u_{1}}$, $\underline{C}_{\underline{u}_{\ast}}$.
+For any $u_{0}<u_{1}\leq\underline{u}_{\ast}$, we recall the definition of the domain $\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast})$ from (1.2), which is bounded by $\Gamma$ and the hypersurfaces $C_{u_{0}}$, $C_{u_{1}}$, $\underline{C}_{\underline{u}_{\ast}}$.
 We make the following simplification
 $\mathcal{D}:=\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast}+\zeta),\quad\text{with some constant}\,\,\zeta>0,$ (3.3)
-and denote the truncated null cones by
+and denote the truncated null cones as
 $C_{u}^{[\underline{u}_{0},\,\underline{u}_{1}]}:=C_{u}\cap\{\underline{u}_{0}\leq\underline{u}\leq\underline{u}_{1}\},\quad\underline{C}_{\underline{u}}^{[u_{0},\,u_{1}]}:=\underline{C}_{\underline{u}}\cap\{u_{0}\leq u\leq u_{1}\}.$
-Besides, we abbreviate the outgoing cone emanating from the vertex $C_{u}^{[u,\,\underline{u}]}$ by $C_{u}$, and $\underline{C}_{\underline{u}}^{[u_{0},\,u]}$ by $\underline{C}_{\underline{u}}$, whenever the interval of integration is clear in the context.
-For the fixed constant $b$ (1.4), we define the energy norms,
-$E_{0}(u,\,\underline{u}):={}$ $\int_{C_{u}}\frac{\Omega^{2}h}{r^{1+b}}\left(r^{-2}\hat{w}^{2}+\alpha^{2}\right)\mathrm{d}\underline{u}^{\prime},$
-$E_{1}(u,\,\underline{u}):={}$ $E_{0}(u,\,\underline{u})+\int_{C_{u}}\frac{\Omega^{2}h}{r^{1+b}}(\alpha^{\prime})^{2}\mathrm{d}\underline{u}^{\prime},$
-$\underline{E}_{0}(u,\,\underline{u}):={}$ $\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(r^{-2}\hat{w}^{2}+h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u^{\prime},$
-$\underline{E}_{0}^{\prime}(u,\,\underline{u}):={}$ $\underline{E}_{0}(u,\,\underline{u})+\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\alpha^{2}\mathrm{d}u^{\prime}.$ (3.5)
+Furthermore, we abbreviate the outgoing cone segment emanating from the vertex $C_{u}^{[u,\,\underline{u}]}$ as $C_{u}$, and $\underline{C}_{\underline{u}}^{[u_{0},\,u]}$ as $\underline{C}_{\underline{u}}$, when the interval of integration is clear in the context.
+We define the energy norms,
+$E_{0}(u,\,\underline{u}):={}$ $\int_{C_{u}}\frac{\Omega^{2}h}{r}\left(r^{-2}\hat{w}^{2}+\alpha^{2}\right)\mathrm{d}\underline{u}^{\prime},$
+$E_{1}(u,\,\underline{u}):={}$ $E_{0}(u,\,\underline{u})+\int_{C_{u}}\frac{\Omega^{2}h}{r}(\alpha^{\prime})^{2}\mathrm{d}\underline{u}^{\prime},$
+$\underline{E}_{0}(u,\,\underline{u}):={}$ $\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r}\left(r^{-2}\hat{w}^{2}+h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u^{\prime},$
+$\underline{E}_{0}^{\prime}(u,\,\underline{u}):={}$ $\underline{E}_{0}(u,\,\underline{u})+\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r}\alpha^{2}\mathrm{d}u^{\prime}.$ (3.4)
 If we use $r$ as the parameter on $C_{u}$, $\underline{C}_{\underline{u}}$ respectively, then the energies in (3.1) are rewritten as
-$E_{0}={}$ $\int_{0}^{r}\frac{r^{-2}\hat{w}^{2}+\alpha^{2}}{r^{1+b}}\mathrm{d}r|_{C_{u}},$
-$E_{1}={}$ $\int_{0}^{r}\frac{(\alpha^{\prime})^{2}+\alpha^{2}+r^{-2}\hat{w}^{2}}{r^{1+b}}\mathrm{d}r|_{C_{u}},$
-$\underline{E}_{0}={}$ $\int_{r}^{r_{0}}\frac{1}{1-\mu}\frac{r^{-2}\hat{w}^{2}+h^{2}(\partial_{u}w)^{2}}{r^{1+b}}\mathrm{d}r|_{\underline{C}_{\underline{u}}},$
-$\underline{E}_{0}^{\prime}={}$ $\underline{E}_{0}+\int_{r}^{r_{0}}\frac{1}{1-\mu}\frac{\alpha^{2}}{r^{1+b}}\mathrm{d}r|_{\underline{C}_{\underline{u}}},$
+$E_{0}={}$ $\int_{0}^{r}\frac{r^{-2}\hat{w}^{2}+\alpha^{2}}{r}\mathrm{d}r|_{C_{u}},$
+$E_{1}={}$ $\int_{0}^{r}\frac{(\alpha^{\prime})^{2}+\alpha^{2}+r^{-2}\hat{w}^{2}}{r}\mathrm{d}r|_{C_{u}},$
+$\underline{E}_{0}={}$ $\int_{r}^{r_{0}}\frac{1}{1-\mu}\frac{r^{-2}\hat{w}^{2}+h^{2}(\partial_{u}w)^{2}}{r}\mathrm{d}r|_{\underline{C}_{\underline{u}}},$
+$\underline{E}_{0}^{\prime}={}$ $\underline{E}_{0}+\int_{r}^{r_{0}}\frac{1}{1-\mu}\frac{\alpha^{2}}{r}\mathrm{d}r|_{\underline{C}_{\underline{u}}},$
 where $r_{0}:=r(u_{0},\underline{u})$.
-3.1.1. Energy estimates theorem
+3.2. Energy estimates theorem
 The proof of Theorem 3.1 follows from the energy estimates which we state as below.
@@
 $\mu\leq\varepsilon^{2},\quad\text{in}\,\,\mathcal{D},$
-the following statements are concluded.
-Case I: $0<b<\frac{\sqrt{89}-9}{4}$. The solution $(r,m,\alpha)\in C^{1}\times C^{1}\times H^{1}$ in $\mathcal{D}$ and it holds that
+the solution $(r,m,\alpha)\in C^{1}\times C^{0}\times H^{1}$ in $\mathcal{D}$. Moreover, it holds that
 $|r^{-1}\hat{w}|^{2}\leq C(I_{1})\varepsilon^{2},\quad|\alpha|^{2}\leq C(I_{1})\varepsilon,$ (3.5)
-and
-$|r^{-1}\hat{w}|^{2}+|\alpha|^{2}+|\partial_{u}w|^{2}\leq C(I_{1})r^{b},$ (3.7)
 and the energies are uniformly bounded
@@
 for some constant $C(I_{1})$ depending only on $I_{1}$.
-Case II: $b=0$. The solution $(r,m,\alpha)\in C^{1}\times C^{0}\times H^{1}$ in $\mathcal{D}$ and both of (3.6) and (3.8) hold with $b=0$.
-Remark 3.3.
-The energy arguments leading to Lemma 3.2 are closed in $H^{1}$ space. This fact enables us to prove a local well-posedness theorem with $H^{1}$ data, see Appendix A.
-3.1.2. Preliminary estimates
+Remark $3.$
+The energy arguments leading to Lemma 3.2 are closed in the $H^{1}$ space. This fact enables us to prove a local well-posedness theorem with $H^{1}$ data, as detailed in Section 4.
+3.3. Preliminary estimates
 As a preliminary step, we have
 $\underline{h}<0,\quad h>0\quad\text{in}\,\,\mathcal{D}.$ (3.7)
-The fact $\underline{h}<0$ follows from the monotonicity of $\Omega^{-2}\underline{h}$ in $u$ (see (2.4b)) and the initial requirement $\Omega^{-2}\underline{h}|_{u_{0}}<0$, and moreover, $h>0$ is due to the smallness of $\mu$ which indicates $01-\mu=-\underline{h}h>$ 0.
-3.1.3. Continuity argument
-We make the following bootstrap assumption. Let $\Lambda$ be a large constant that will be determined (depending on the initial energy). We assume
-$|r^{-1}\hat{w}|+|\alpha|\leq\varepsilon\Lambda,\quad\text{in}\,\,\mathcal{D}.$ (3.10)
-We will show that the same inequality with $\Lambda$ being replaced by $\frac{\Lambda}{2}$ holds.
+The fact $\underline{h}<0$ follows from the monotonicity of $\Omega^{-2}\underline{h}$ in $u$ (see (2.4b)) and the initial requirement $\Omega^{-2}\underline{h}|_{u_{0}}<0$, and moreover, $h>0$ is due to the smallness of $\mu$ which indicates $1-\mu=-\underline{h}h>0$.
+3.4. Continuity argument
+We make the following bootstrap assumption. Let $\Lambda$ be a large constant that will be determined (depending on the initial energy) later. We assume
+$|r^{-1}\hat{w}|^{2}+|\alpha|^{2}\leq\varepsilon\Lambda,\quad\text{in}\,\,\mathcal{D}.$ (3.8)
+and will show that the same inequality with $\Lambda$ being replaced by $\frac{\Lambda}{2}$ holds.
 Since $r$ is finite in $\mathcal{D}$, it is straightforward to know from (3.8) that
-$|\Phi|\lesssim\varepsilon\Lambda,\,\,|w^{2}-1|\lesssim\varepsilon\Lambda,\quad\text{in}\,\,\mathcal{D}.$ (3.11)
-3.1.4. Main energy estimates
+$|\Phi|^{2}\lesssim\varepsilon\Lambda,\,\,\text{and}\,\,|w^{2}-1|^{2}\lesssim\varepsilon\Lambda,\quad\text{in}\,\,\mathcal{D}.$ (3.9)
+3.5. Main energy estimates
 We will use the bootstrap assumption (3.8) to carry out the energy estimates in details.
-Let us start with the first order energy identities. We multiply $2\partial_{\underline{u}}w$ on the Yang–Mills equation (2.5) to obtain the identity
+Let us begin with the first order energy identities. We multiply $2\partial_{\underline{u}}w$ on the Yang–Mills equation (2.5) to obtain the identity
 $\partial_{u}(\partial_{\underline{u}}w)^{2}+\frac{\Omega^{2}}{2r^{2}}\partial_{\underline{u}}Q^{2}=0.$ (3.10)
-We further multiply $\Omega^{-2}h^{-1}r^{-1-b}$ on (3.12), then
-$\partial_{u}\left(r^{-1-b}\Omega^{-2}h^{-1}(\partial_{\underline{u}}w)^{2}\right)-\partial_{u}r^{-1-b}\Omega^{-2}h^{-1}(\partial_{\underline{u}}w)^{2}$
-$\qquad-\partial_{u}(\Omega^{-2}h^{-1})\frac{(\partial_{\underline{u}}w)^{2}}{r^{1+b}}+\partial_{\underline{u}}\left(\frac{h^{-1}}{2r^{3+b}}Q^{2}\right)-\partial_{\underline{u}}\left(\frac{h^{-1}}{2r^{3+b}}\right)Q^{2}=0.$
+We further multiply $\Omega^{-2}h^{-1}r^{-1}$ on (3.10), then
+$\partial_{u}\left(r^{-1}\Omega^{-2}h^{-1}(\partial_{\underline{u}}w)^{2}\right)-\partial_{u}r^{-1}\Omega^{-2}h^{-1}(\partial_{\underline{u}}w)^{2}$
+$\qquad-\partial_{u}(\Omega^{-2}h^{-1})\frac{(\partial_{\underline{u}}w)^{2}}{r}+\partial_{\underline{u}}\left(\frac{h^{-1}}{2r^{3}}Q^{2}\right)-\partial_{\underline{u}}\left(\frac{h^{-1}}{2r^{3}}\right)Q^{2}=0.$
 Using the equations (2.4a) and (2.4d), and integrating over the domain $\mathcal{D}$, we arrive at the following energy identity
-$\int_{C_{u}}\Omega^{2}h\frac{\alpha^{2}}{r^{1+b}}\mathrm{d}\underline{u}+\int_{\underline{C}_{\underline{u}}}h^{-1}\frac{\Phi^{2}}{2r^{1+b}}\mathrm{d}u$
-$+\iint_{\mathcal{D}}\left(\frac{3+b}{2}\frac{\Phi^{2}}{r^{2+b}}-(1+b)\frac{\alpha^{2}}{r^{2+b}}+b\mu\frac{\alpha^{2}}{r^{2+b}}\right)\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
-[…]
-Note that, among the spacetime integrals on the left hand side of (3.1.4), there is a negative term, which prevents us from deducing any energy bound at this stage.
+$\int_{C_{u}}\Omega^{2}h\frac{\alpha^{2}}{r}\mathrm{d}\underline{u}+\int_{\underline{C}_{\underline{u}}}h^{-1}\frac{\Phi^{2}}{2r}\mathrm{d}u$
+$+\iint_{\mathcal{D}}\left(\frac{3}{2}\frac{\Phi^{2}}{r^{2}}-\frac{\alpha^{2}}{r^{2}}\right)\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
+$={}$ $\int_{C_{u_{0}}}\Omega^{2}h\frac{\alpha^{2}}{r}\mathrm{d}\underline{u}.$ (3.11)
+Note that, among the spacetime integrals on the left-hand side of (3.5), there appears a negative term, which obstructs the derivation of energy bounds at this stage.
 On the other hand, multiplying $2\partial_{u}w$ on the Yang–Mills equation (2.5) yields
 $\partial_{\underline{u}}(\partial_{u}w)^{2}+\frac{\Omega^{2}}{2r^{2}}\partial_{u}Q^{2}=0.$ (3.12)
-Multiplied with $hr^{-1-b}$, (3.14) is turned into the following identity with the assistance of (2.4a) and (2.4d),
-$\partial_{\underline{u}}\left(h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\right)+(1+b)\Omega^{2}h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}+2\Omega^{-2}(\partial_{\underline{u}}w)^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}$
-$\quad+\partial_{u}\left(\frac{\Omega^{2}h}{2}\frac{Q^{2}}{r^{3+b}}\right)+\frac{3+b}{2}\Omega^{2}\underline{h}h\frac{Q^{2}}{r^{4+b}}+\frac{\Omega^{2}}{2}\left(\frac{\mu}{r}-\frac{Q^{2}}{r^{3}}\right)\frac{Q^{2}}{r^{3+b}}=0.$
+Multiplying by $hr^{-1}$ and using equations (2.4a), (2.4d), we transform (3.12) into the following identity,
+$\partial_{\underline{u}}\left(h\frac{(\partial_{u}w)^{2}}{r}\right)+\Omega^{2}h^{2}\frac{(\partial_{u}w)^{2}}{r^{2}}+2\Omega^{-2}(\partial_{\underline{u}}w)^{2}\frac{(\partial_{u}w)^{2}}{r^{2}}$
+$\quad+\partial_{u}\left(\frac{\Omega^{2}h}{2}\frac{Q^{2}}{r^{3}}\right)+\frac{3}{2}\Omega^{2}\underline{h}h\frac{Q^{2}}{r^{4}}+\frac{\Omega^{2}}{2}\left(\frac{\mu}{r}-\frac{Q^{2}}{r^{3}}\right)\frac{Q^{2}}{r^{3}}=0.$
 We then obtain an energy identity, after integrating over the domain $\mathcal{D}$,
-$\int_{\underline{C}_{\underline{u}}}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\mathrm{d}u+\int_{C_{u}}\frac{\Omega^{2}h}{2}\frac{\Phi^{2}}{r^{1+b}}\mathrm{d}\underline{u}$
-$+\iint_{\mathcal{D}}\left((1+b)+2\alpha^{2}\right)\Omega^{2}h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$
-$+\iint_{\mathcal{D}}\left(-\frac{3+b}{2}+\left(2+\frac{b}{2}\right)\mu\right)\Omega^{2}\frac{\Phi^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$
-[…]
-Analogous to (3.1.4), the above identity has a nonpositive spacetime integral on the left hand side and therefore fails to conclude any energy bound.
-What is more, cooperating the energy identity (3.1.4) with (3.1.4) does not give any bound for the first order energy neither.
-In what follows, we proceed to investigate the second order energy identities, which we expect to cancel those terms with unfavourable signs in (3.1.4) and (3.1.4).
-By virtue of (2.3), there is
-$\partial_{u}\alpha^{\prime}={}$ $\frac{\underline{h}}{r}\alpha^{\prime}-2w^{2}\frac{h^{-1}}{r}\alpha+2w\frac{h^{-1}}{r}\Phi$
+$\int_{\underline{C}_{\underline{u}}}h\frac{(\partial_{u}w)^{2}}{r}\mathrm{d}u+\int_{C_{u}}\Omega^{2}h\frac{\Phi^{2}}{2r}\mathrm{d}\underline{u}$
+$+\iint_{\mathcal{D}}\left(1+2\alpha^{2}\right)\Omega^{2}h^{2}\frac{(\partial_{u}w)^{2}}{r^{2}}\mathrm{d}u\mathrm{d}\underline{u}$
+$+\iint_{\mathcal{D}}\left(-\frac{3}{2}+2\mu\right)\Omega^{2}\frac{\Phi^{2}}{r^{2}}\mathrm{d}u\mathrm{d}\underline{u}$
+$={}$ $\int_{C_{u_{0}}}\Omega^{2}h\frac{\Phi^{2}}{2r}\mathrm{d}\underline{u}+\iint_{\mathcal{D}}\Omega^{2}\Phi^{2}\frac{\Phi^{2}}{2r^{2}}\mathrm{d}u\mathrm{d}\underline{u}.$ (3.13)
+Analogous to (3.5), the identity above contains a nonpositive spacetime integral on the left-hand side, which prevents it from yielding an energy bound.
+What is more, cooperating the energy identity (3.5) with (3.5) does not give any bound for the first order energies neither.
+To address this, we now derive the second order energy identities, which are designed to absorb the terms with unfavorable signs in (3.5) and (3.5).
+By virtue of (2.3), we have
+$\partial_{u}\alpha^{\prime}={}$ $\frac{\underline{h}}{r}\alpha^{\prime}-2w^{2}\frac{h^{-1}}{r}\alpha+2w\frac{h^{-1}}{r}\frac{Q}{r}$
 $+2\frac{h^{-1}}{r}\left(\mu-\Phi^{2}\right)\alpha^{\prime}-h^{-1}\Phi\alpha-6w\frac{h^{-1}}{r}\Phi\alpha^{2}$
 $+2\frac{h^{-1}}{r}\left(2\Phi^{2}-\mu\right)\alpha+2\frac{h^{-1}}{r}\left(1-\Phi^{2}\right)\alpha^{3},$ (3.14)
-where we denote the linear terms on the first line as follows
+where we mark the linear terms on the first line as below
 $\ell_{1}=\frac{\underline{h}}{r}\alpha^{\prime},\quad\ell_{2}=-2w^{2}\frac{h^{-1}}{r}\alpha,\quad\ell_{3}=2w\frac{h^{-1}}{r}\frac{Q}{r}.$
-Before going into the lengthy calculations for the second order energy estimates, here we sketch the main idea of proof. To obtain an energy identity from (3.1.4), we will multiply on this equation
-$r^{-1-b}\cdot\Omega^{2}h\alpha^{\prime}=r^{-b}\partial_{\underline{u}}\alpha$
-and integrate over $\iint_{\mathcal{D}}\mathrm{d}u\mathrm{d}\underline{u}$. Regarding to the first terms on the right hand side of the first line of (3.1.4),
-$\iint_{\mathcal{D}}\Omega^{2}h\frac{\alpha^{\prime}}{r^{1+b}}\cdot\ell_{1}\mathrm{d}u\mathrm{d}\underline{u}=-\iint_{\mathcal{D}}\left(1-\mu\right)\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$ (3.17)
-[…]
-$\iint_{\mathcal{D}}\Omega^{2}h\frac{\alpha^{\prime}}{r^{1+b}}\cdot\ell_{i}\,\mathrm{d}u\mathrm{d}\underline{u}$
-$\lesssim$ $\left(\iint_{\mathcal{D}}\Omega^{2}\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}\right)^{\frac{1}{2}}\cdot\left(\iint_{\mathcal{D}}\Omega^{2}\left(\frac{\alpha^{2}}{r^{2+b}}+\frac{\Phi^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}\right)^{\frac{1}{2}}$
-is not allowed, since we have no yet any improved bounds for the energies on the right hand side of the inequality above.
-The difficulty coming from these quadratic terms is resolved via applying integration by parts.
-In this way, we expect to eventually gain spacetime integrals, such as $\iint_{\mathcal{D}}\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$, $\iint_{\mathcal{D}}\Omega^{2}\frac{\Phi^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$, with favourable signs.
-Among all the terms on the right side of (3.1.4), the nonlinear terms in the second and third lines can be in principle estimated by means of (after multiplying $r^{-1-b}\cdot\Omega^{2}h\alpha^{\prime}$)
-$\lesssim\varepsilon\Lambda\iint_{\mathcal{D}}\left(\frac{\Phi^{2}}{r^{2+b}}+\frac{\alpha^{2}}{r^{2+b}}+\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\right)\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}.$ (3.18)
-The extra smallness of (3.18) suggests they could be absorbed, since in the preceding steps of treating the linear terms $\ell_{i}$, we have managed to extract a positive bulk terms as $\iint_{\mathcal{D}}\left(\frac{\Phi^{2}}{r^{2+b}}+\frac{\alpha^{2}}{r^{2+b}}+\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\right)\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$ on the left hand side of the energy inequality through integration by parts.
-We now go on with the detailed proof. First of all, let us turn to the left hand side of (3.1.4). After multiplied with $\Omega^{2}h\frac{\alpha^{\prime}}{r^{1+b}}$, it becomes,
-$\partial_{u}\alpha^{\prime}\cdot\Omega^{2}h\frac{\alpha^{\prime}}{r^{1+b}}$
-[…]
-[…]
-We notice that the last term
-$\frac{1+b}{2}\underline{h}h\Omega^{2}\frac{(\alpha^{\prime})^{2}}{r^{2+b}}=-\frac{1+b}{2}(1-\mu)\Omega^{2}\frac{(\alpha^{\prime})^{2}}{r^{2+b}}$
-which indeed lies on the left hand side the energy identity has a bad sign. Nevertheless, this term is compensated by (3.17) and their summation still contributes a positive term
-$\iint_{\mathcal{D}}\frac{1-b}{2}\left(1-\mu\right)\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u},\quad 0\leq b<\frac{\sqrt{89}-9}{4},$
-on the left side of the energy identity.
-[…]
-$r^{-1-b}\Omega^{2}h\alpha^{\prime}\cdot\ell_{2}=-w^{2}h^{-1}r^{-1-b}\partial_{\underline{u}}\alpha^{2}$
-[…]
-$+w^{2}r^{-1-b}\alpha^{2}\partial_{\underline{u}}h^{-1}+w^{2}\partial_{\underline{u}}r^{-1-b}\alpha^{2}h^{-1}$
-[…]
-$+2w^{2}\Omega^{2}r^{-2-b}\alpha^{4}-(1+b)w^{2}\Omega^{2}r^{-2-b}\alpha^{2},$
+Before proceeding to the detailed second order energy estimates, we outline the main idea of the proof. To obtain an energy identity from (3.5), we multiply the equation by
+$r^{-1}\cdot\Omega^{2}h\alpha^{\prime}=\partial_{\underline{u}}\alpha$
+and integrate over $\iint_{\mathcal{D}}\mathrm{d}u\mathrm{d}\underline{u}$. Regarding to the first terms on the right-hand side of the first line,
+$\iint_{\mathcal{D}}\Omega^{2}h\frac{\alpha^{\prime}}{r}\cdot\ell_{1}\mathrm{d}u\mathrm{d}\underline{u}=-\iint_{\mathcal{D}}\left(1-\mu\right)\frac{(\alpha^{\prime})^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$ (3.15)
+admits a favourable sign and can be moved to the left-hand side.
+In contrast, the remaining terms $\iint_{\mathcal{D}}r^{-1}\cdot\Omega^{2}h\alpha^{\prime}\cdot\ell_{i}\,\mathrm{d}u\mathrm{d}\underline{u}$, i=2, 3i=2,\,3, which are now quadratic, lack definite signs.
+A straightforward estimate such as
+$\iint_{\mathcal{D}}\Omega^{2}h\frac{\alpha^{\prime}}{r}\cdot\ell_{i}\,\mathrm{d}u\mathrm{d}\underline{u}$
+$\lesssim$ $\left(\iint_{\mathcal{D}}\Omega^{2}\frac{(\alpha^{\prime})^{2}}{r^{2}}\mathrm{d}u\mathrm{d}\underline{u}\right)^{\frac{1}{2}}\cdot\left(\iint_{\mathcal{D}}\Omega^{2}\left(\frac{\alpha^{2}}{r^{2}}+\frac{\Phi^{2}}{r^{2}}\right)\mathrm{d}u\mathrm{d}\underline{u}\right)^{\frac{1}{2}}$
+is not allowed at this stage, as we do not yet have improved bounds for the spacetime integrals on the right-hand side.
+The difficulty posed by these quadratic terms is resolved via applying integration by parts.
+This allows us to ultimately obtain spacetime integrals with favorable signs, such as $\iint_{\mathcal{D}}\Omega^{2}\frac{\alpha^{2}}{r^{2}}\mathrm{d}u\mathrm{d}\underline{u}$, $\iint_{\mathcal{D}}\Omega^{2}\frac{\Phi^{2}}{r^{2}}\mathrm{d}u\mathrm{d}\underline{u}$.
+Among all the terms on the right-hand side, the nonlinear terms in the second and third lines can, in principle, be estimated by means of (after multiplication by $r^{-1}\cdot\Omega^{2}h\alpha^{\prime}$)
+$\lesssim\varepsilon\Lambda\iint_{\mathcal{D}}\left(\frac{\Phi^{2}}{r^{2}}+\frac{\alpha^{2}}{r^{2}}+\frac{(\alpha^{\prime})^{2}}{r^{2}}\right)\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}.$ (3.16)
+Actually, the term (3.16) is expected to be absorbed, since during the treatment of the linear terms $\ell_{i}$, we have already managed to generate positive bulk terms of the form $\iint_{\mathcal{D}}\left(\frac{\Phi^{2}}{r^{2}}+\frac{\alpha^{2}}{r^{2}}+\frac{(\alpha^{\prime})^{2}}{r^{2}}\right)\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$ on the left-hand side of the energy inequality via integration by parts.
+We now go on with the detailed proof. First of all, we analyze the left-hand side of (3.5). After multiplied $\Omega^{2}h\frac{\alpha^{\prime}}{r}$, it becomes,
+$\partial_{u}\alpha^{\prime}\cdot\Omega^{2}h\frac{\alpha^{\prime}}{r}$
+$={}$ $\partial_{u}\left(\Omega^{2}h\frac{(\alpha^{\prime})^{2}}{2r}\right)-\partial_{u}\left(\Omega^{2}h\right)\frac{(\alpha^{\prime})^{2}}{2r}-\frac{1}{2}\partial_{u}\left(\frac{1}{r}\right)\Omega^{2}h(\alpha^{\prime})^{2}$
+$={}$ $\partial_{u}\left(\Omega^{2}h\frac{(\alpha^{\prime})^{2}}{2r}\right)+\Omega^{2}\left(\frac{\mu}{r}-\frac{Q^{2}}{r^{3}}\right)\frac{(\alpha^{\prime})^{2}}{2r}+\frac{1}{2}\underline{h}h\Omega^{2}\frac{(\alpha^{\prime})^{2}}{r^{2}}.$
+We observe that the last term
+$\frac{1}{2}\underline{h}h\Omega^{2}\frac{(\alpha^{\prime})^{2}}{r^{2}}=-\frac{1}{2}(1-\mu)\Omega^{2}\frac{(\alpha^{\prime})^{2}}{r^{2}}$
+on the left-hand side of the energy identity carries an unfavorable sign. Nevertheless, this term can be absorbed by the positive contribution from (3.15), and ultimately yields a net positive term
+$\iint_{\mathcal{D}}\frac{1}{2}\left(1-\mu\right)\frac{(\alpha^{\prime})^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
+on the left-hand side of the energy identity.
+To proceed, we focus on the two terms $\iint_{\mathcal{D}}r^{-1}\Omega^{2}h\alpha^{\prime}\cdot\ell_{i}\,\mathrm{d}u\mathrm{d}\underline{u}$, i=2, 3i=2,\,3. When $i=2$, it holds that
+$r^{-1}\Omega^{2}h\alpha^{\prime}\cdot\ell_{2}=-w^{2}h^{-1}r^{-1}\partial_{\underline{u}}\alpha^{2}$
+$={}$ $-\partial_{\underline{u}}\left(w^{2}h^{-1}r^{-1}\alpha^{2}\right)+\partial_{\underline{u}}w^{2}h^{-1}r^{-1}\alpha^{2}$
+$+w^{2}r^{-1}\alpha^{2}\partial_{\underline{u}}h^{-1}+w^{2}\partial_{\underline{u}}r^{-1}\alpha^{2}h^{-1}$
+$={}$ $-\partial_{\underline{u}}\left(w^{2}h^{-1}r^{-1}\alpha^{2}\right)+2\Omega^{2}wr^{-1}\alpha^{3}$
+$+2w^{2}\Omega^{2}r^{-2}\alpha^{4}-w^{2}\Omega^{2}r^{-2}\alpha^{2},$
 where we have made use of the equation (2.4a) in the last identity.
-In particular, we notice that the last term $-(1+b)w^{2}\Omega^{2}r^{-2-b}\alpha^{2}$ which lies on the right hand side the energy identity admits a favourable sign.
+In particular, we note that the last term $w^{2}\Omega^{2}r^{-2}\alpha^{2}$ on the right-hand side the energy identity has a favourable sign.
 When $i=3$, we have
-$r^{-1-b}\Omega^{2}h\alpha^{\prime}\cdot\ell_{3}=2h^{-1}r^{-2-b}(w^{2}-1)w\cdot\partial_{\underline{u}}\alpha$
-[…]
-$-\partial_{\underline{u}}w\cdot 2r^{-2-b}h^{-1}(w^{2}-1)\alpha$
-[…]
-$-4h^{-2}\Omega^{-2}\frac{(\partial_{\underline{u}}w)^{2}}{r^{3+b}}(w^{2}-1)w\alpha-4\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}-6(w^{2}-1)\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}.$
-Here the quadratic term $-4\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}$ admits a favourable sign, while the sign of $(4+2b)\Omega^{2}\frac{w^{2}-1}{r^{3+b}}w\alpha$ is not definite.
-Nevertheless, we can perform the same trick on this quadratic term as follows
-$(4+2b)\Omega^{2}\frac{w^{2}-1}{r^{3+b}}w\alpha=(2+b)h^{-1}\frac{w^{2}-1}{r^{3+b}}\partial_{\underline{u}}(w^{2}-1)$
-[…]
-$+(3+b)\left(1+\frac{b}{2}\right)\Omega^{2}\frac{(w^{2}-1)^{2}}{r^{4+b}}.$
+$r^{-1}\Omega^{2}h\alpha^{\prime}\cdot\ell_{3}=2h^{-1}r^{-2}(w^{2}-1)w\cdot\partial_{\underline{u}}\alpha$
+$={}$ $\partial_{\underline{u}}\left(2r^{-2}h^{-1}(w^{2}-1)w\alpha\right)-\partial_{\underline{u}}\left(2r^{-2}h^{-1}(w^{2}-1)\right)w\alpha$
+$-\partial_{\underline{u}}w\cdot\left(2r^{-2}h^{-1}(w^{2}-1)\alpha\right)$
+$={}$ $\partial_{\underline{u}}\left(2r^{-2}h^{-1}(w^{2}-1)w\alpha\right)+4\Omega^{2}r^{-3}(w^{2}-1)w\alpha$
+$-4h^{-2}\Omega^{-2}\frac{(\partial_{\underline{u}}w)^{2}}{r^{3}}(w^{2}-1)w\alpha-4\Omega^{2}\frac{\alpha^{2}}{r^{2}}-6(w^{2}-1)\Omega^{2}\frac{\alpha^{2}}{r^{2}}.$
+Here the quadratic term $4\Omega^{2}\frac{\alpha^{2}}{r^{2}}$ admits a favourable sign, whereas $4\Omega^{2}\frac{w^{2}-1}{r^{3}}w\alpha$ lacks a definite sign.
+Nevertheless, we can perform the same technique to this quadratic term to proceed
+$4\Omega^{2}\frac{w^{2}-1}{r^{3}}w\alpha=2h^{-1}\frac{w^{2}-1}{r^{3}}\partial_{\underline{u}}(w^{2}-1)$
+$={}$ $\partial_{\underline{u}}\left(h^{-1}\frac{(w^{2}-1)^{2}}{r^{3}}\right)-\partial_{\underline{u}}h^{-1}\frac{(w^{2}-1)^{2}}{r^{3}}+3\Omega^{2}\frac{(w^{2}-1)^{2}}{r^{4}}.$
 We then combine these calculations to obtain
-$r^{-1-b}\Omega^{2}h\alpha^{\prime}\cdot\ell_{3}={}$ $\partial_{\underline{u}}\left(2wh^{-1}\frac{\Phi\alpha}{r^{1+b}}\right)+\left(1+\frac{b}{2}\right)\partial_{\underline{u}}\left(h^{-1}\frac{\Phi^{2}}{r^{1+b}}\right)$
-$-4\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}+(3+b)\left(1+\frac{b}{2}\right)\Omega^{2}\frac{\Phi^{2}}{r^{2+b}}$
-$-2\left(1+\frac{b}{2}\right)\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}\Phi^{2}-4\Omega^{2}\frac{\Phi}{r^{2+b}}w\alpha^{3}-6\Omega^{2}\Phi\frac{\alpha^{2}}{r^{1+b}},$
-where in the second line, the quadratic term $-4\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}$ contributes a positive term in the energy identity, while $(3+b)\left(1+\frac{b}{2}\right)\Omega^{2}\frac{\Phi^{2}}{r^{2+b}}$ offers a bad sign.
-Nevertheless, noting the positive integral $\iint_{\mathcal{D}}\frac{\Phi^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$ on the left hand side of the first order energy identity (3.1.4), the bad-sign term $(3+b)\left(1+\frac{b}{2}\right)\Omega^{2}\frac{\Phi^{2}}{r^{2+b}}$ will be absorbed when (3.1.4) is added into the second order energy identity.
-Taking all the calculations into accounts, we integrate these identities over $\iint_{\mathcal{D}}\mathrm{d}u\mathrm{d}\underline{u}$ to derive the energy identity
-$\int_{C_{u}}\frac{(\alpha^{\prime})^{2}}{2r^{1+b}}\Omega^{2}h\mathrm{d}\underline{u}+\int_{\underline{C}_{\underline{u}}}\frac{1}{r^{1+b}}\left(\left(w\alpha-\Phi\right)^{2}-\left(2+\frac{b}{2}\right)\Phi^{2}\right)h^{-1}\mathrm{d}u$
-$+\iint_{\mathcal{D}}\left(\frac{1-b}{2}+\frac{3\Phi^{2}}{2}\right)\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
-$+\iint_{\mathcal{D}}\left((5+b)+(2+b)\Phi^{2}\right)\frac{\alpha^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
-$-\iint_{\mathcal{D}}(3+b)\left(1+\frac{b}{2}\right)\frac{\Phi^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
-[…]
-$+\iint_{\mathcal{D}}\left(2(1-\Phi^{2})\alpha^{2}+4\Phi^{2}-2\mu-6w\Phi\alpha-r\Phi\right)\frac{\alpha\alpha^{\prime}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
-$+\iint_{\mathcal{D}}\left(2\left(wr\alpha+w^{2}\alpha^{2}\right)-4w\Phi\alpha-(7+b)r\Phi\right)\frac{\alpha^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}.$ (3.19)
-Now in view of the identities (3.1.4), (3.1.4) and (3.1.4), to eliminate the negative term $-\iint_{\mathcal{D}}(3+b)\left(1+\frac{b}{2}\right)\Omega^{2}\frac{\Phi^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$ on the left hand side of (3.1.4), and obtain coercive energy on $\underline{C}_{\underline{u}}$, we consider the linear combination
+$r^{-1}\Omega^{2}h\alpha^{\prime}\cdot\ell_{3}$
+$={}$ $\partial_{\underline{u}}\left(2wh^{-1}\frac{\Phi\alpha}{r}\right)+\partial_{\underline{u}}\left(h^{-1}\frac{\Phi^{2}}{r}\right)$
+$-4\Omega^{2}\frac{\alpha^{2}}{r^{2}}+3\Omega^{2}\frac{\Phi^{2}}{r^{2}}-2\Omega^{2}\frac{\alpha^{2}}{r^{2}}\Phi^{2}-4\Omega^{2}\frac{\Phi}{r^{2}}w\alpha^{3}-6\Omega^{2}\Phi\frac{\alpha^{2}}{r}.$
+In the third line, the quadratic term $4\Omega^{2}\frac{\alpha^{2}}{r^{2}}$ contributes positively to the energy identity, while $3\Omega^{2}\frac{\Phi^{2}}{r^{2}}$ has an unfavorable sign.
+Nevertheless, noting the presence of the positive integral $\iint_{\mathcal{D}}\frac{\Phi^{2}}{r^{2}}\mathrm{d}u\mathrm{d}\underline{u}$ on the left-hand side of the first order energy identity (3.5), the term $3\Omega^{2}\frac{\Phi^{2}}{r^{2}}$ can be absorbed when (3.5) is taken into accounts.
+Integrating over $\iint_{\mathcal{D}}\mathrm{d}u\mathrm{d}\underline{u}$ leads to the energy identity
+$\int_{C_{u}}\frac{(\alpha^{\prime})^{2}}{2r}\Omega^{2}h\mathrm{d}\underline{u}+\int_{\underline{C}_{\underline{u}}}\frac{1}{r}\left(\left(w\alpha-\Phi\right)^{2}-2\Phi^{2}\right)h^{-1}\mathrm{d}u$
+$+\iint_{\mathcal{D}}\left(\frac{1}{2}+\frac{3\Phi^{2}}{2}\right)\frac{(\alpha^{\prime})^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
+$+\iint_{\mathcal{D}}\left(5+2\Phi^{2}\right)\frac{\alpha^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}-\iint_{\mathcal{D}}3\frac{\Phi^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
+$={}$ $\int_{C_{u_{0}}}\frac{(\alpha^{\prime})^{2}}{2r}\Omega^{2}h\mathrm{d}\underline{u}+\iint_{\mathcal{D}}2\mu\frac{(\alpha^{\prime})^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
+$+\iint_{\mathcal{D}}\left(2(1-\Phi^{2})\alpha^{2}+4\Phi^{2}-2\mu-6w\Phi\alpha-r\Phi\right)\frac{\alpha\alpha^{\prime}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
+$+\iint_{\mathcal{D}}\left(2\left(wr\alpha+w^{2}\alpha^{2}\right)-4w\Phi\alpha-7r\Phi\right)\frac{\alpha^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}.$ (3.17)
+Now in view of the identities (3.5), (3.5) and (3.5), we aim to eliminate the negative term $\iint_{\mathcal{D}}3\Omega^{2}\frac{\Phi^{2}}{r^{2}}\mathrm{d}u\mathrm{d}\underline{u}$ on the left-hand side of (3.5), and meanwhile obtain a coercive energy on $\underline{C}_{\underline{u}}$.
+To achieve this, we form the linear combination
 […]
 which leads to
-$\int_{C_{u}}\frac{\Omega^{2}h}{r^{1+b}}\left((9+2b)\alpha^{2}+(\alpha^{\prime})^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}$
-$+\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(2\left(w\alpha-\Phi\right)^{2}+\frac{\Phi^{2}}{2}+2h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u$
-$+\iint_{\mathcal{D}}\left((1-b)+3\Phi^{2}\right)\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
-$+\iint_{\mathcal{D}}\left((1-9b-2b^{2})+2(2+b)\Phi^{2}+(9b+2b^{2})\mu\right)\frac{\alpha^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
-$+\iint_{\mathcal{D}}\left(\frac{3(3+b)}{2}+(4+b)\mu\right)\frac{\Phi^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
-$+\iint_{\mathcal{D}}\left(2(1+b)+4\alpha^{2}\right)\frac{h^{2}(\partial_{u}w)^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
-[…]
-$+\iint_{\mathcal{D}}\left(\mu+r|\Phi|+\Phi^{2}+\alpha^{2}\right)\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
-$+\iint_{\mathcal{D}}\left(\mu+r|\Phi|+\Phi^{2}+\alpha^{2}+r|\alpha|\right)\frac{\alpha^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}.$ (3.20)
-We note that, since $0\leq b<\frac{\sqrt{89}-9}{4}$ and hence $01-9b-2b^{2}>$ 0, terms on the left hand side of (3.1.4) are all positive.
+$\int_{C_{u}}\frac{\Omega^{2}h}{r}\left(9\alpha^{2}+(\alpha^{\prime})^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}$
+$+\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r}\left(2\left(w\alpha-\Phi\right)^{2}+\frac{\Phi^{2}}{2}+2h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u$
+$+\iint_{\mathcal{D}}\left(1+3\Phi^{2}\right)\frac{(\alpha^{\prime})^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}+\iint_{\mathcal{D}}\left(1+4\Phi^{2}\right)\frac{\alpha^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
+$+\iint_{\mathcal{D}}\left(\frac{9}{2}+4\mu\right)\frac{\Phi^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}+\iint_{\mathcal{D}}\left(2+4\alpha^{2}\right)\frac{h^{2}(\partial_{u}w)^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
+$={}$ $\int_{C_{u_{0}}}\frac{\Omega^{2}h}{r}\left(9\alpha^{2}+(\alpha^{\prime})^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}+\iint_{\mathcal{D}}\Phi^{2}\frac{\Phi^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
+$+\iint_{\mathcal{D}}2\left(2(1-\Phi^{2})\alpha^{2}+4\Phi^{2}-2\mu-6w\Phi\alpha-r\Phi\right)\frac{\alpha\alpha^{\prime}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
+$+\iint_{\mathcal{D}}2\left(2\left(wr\alpha+w^{2}\alpha^{2}\right)-4w\Phi\alpha-7r\Phi\right)\frac{\alpha^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}.$ (3.18)
+We note that, terms on the left-hand side of (3.5) are all positive.
 In particular, the energy on $\underline{C}_{\underline{u}}$ is coercive (the energy on $C_{u}$ is obviously coercive), and there is
-$\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(\Phi^{2}+\alpha^{2}+h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u$
-$\lesssim$ $\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(2\left(w\alpha-\Phi\right)^{2}+\frac{\Phi^{2}}{2}+2h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u.$
-In view of the bootstrap assumption (3.10) and the fact that $\mu$ is small, and $r$ is finite in $\mathcal{D}$, the spacetime integrals on the right hand side of (3.1.4) can be absorbed by the bulk integrals on the left hand side.
+$\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r}\left(\Phi^{2}+\alpha^{2}+h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u$
+$\lesssim$ $\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r}\left(2\left(w\alpha-\Phi\right)^{2}+\frac{\Phi^{2}}{2}+2h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u.$
+In view of the bootstrap assumption (3.8), the smallness of $\mu$, and the boundedness of $r$ in $\mathcal{D}$, the spacetime integrals on the right-hand side of (3.5) can be absorbed by the bulk integrals on the left-hand side.
 We thus achieve the energy bound
-$\int_{C_{u}}\frac{\Omega^{2}h}{r^{1+b}}\left(\Phi^{2}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}\underline{u}$
-$+\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(\Phi^{2}+\alpha^{2}+h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u\lesssim I_{1},$ (3.21)
-and then (3.8) follows.
-3.1.5. Pointwise estimates
-With the help of the energy estimates obtained in the previous step, we will deduce various $L^{\infty}$ estimates to improve the bootstrap assumption.
+$\int_{C_{u}}\frac{\Omega^{2}h}{r}\left(\alpha^{2}+(\alpha^{\prime})^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}$
+$+\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r}\left(\Phi^{2}+\alpha^{2}+h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u\lesssim I_{1}.$ (3.19)
+As a result, the energy bound (3.6) follows.
+3.6. Pointwise estimates
+With the help of energy estimates derived earlier, we now proceed to establish various $L^{\infty}$ estimates in order to improve the bootstrap assumption.
 Lemma 3.4.
 We have the enhanced estimates for $\mu$
-$\mu\lesssim r^{b}\quad\text{and}\quad\lim_{u\rightarrow\underline{u}_{\ast}}\sup_{\mathcal{D}(u,\,\underline{u}_{\ast})}\frac{\mu}{r^{b}}=0.$
-In particular, since $b\geq 0$, there is
-$\mu|_{\Gamma}=0.$
+$\lim_{u\rightarrow\underline{u}_{\ast}}\sup_{\mathcal{D}(u,\,\underline{u}_{\ast})}\mu=0,\quad\text{and}\quad\mu|_{\Gamma}=0.$
 Proof.
@@
 $m\sim\int_{0}^{r}\left(\alpha^{2}+\Phi^{2}\right)\mathrm{d}r^{\prime}|_{C_{u}}.$ (3.20)
-Since $0\leq r^{\prime}\leq r$, and hence $\frac{1}{r}\leq\frac{1}{r^{\prime}}$, the energy bound (3.1.4) leads to
-$\frac{m}{r^{1+b}}\leq\int_{0}^{r}\frac{\left(\left(1-\mu\right)\alpha^{2}+\Phi^{2}\right)}{r^{\prime 1+b}}\mathrm{d}r^{\prime}|_{C_{u}}\leq E_{0}\lesssim I_{0}.$
+Since $0\leq r^{\prime}\leq r$, and hence $\frac{1}{r}\leq\frac{1}{r^{\prime}}$, the energy bound (3.5) leads to
+$\frac{m}{r}\leq\int_{0}^{r}\frac{\left(\left(1-\mu\right)\alpha^{2}+\Phi^{2}\right)}{r^{\prime}}\mathrm{d}r^{\prime}|_{C_{u}}\leq E_{0}\lesssim I_{0}.$
 This lemma is concluded.
-As a sequel, we will improve the pointwise estimates for $\alpha$ and $r^{-1}\hat{w}$.
+As a next step, we will improve the pointwise estimates for $\alpha$ and $r^{-1}\hat{w}$.
 Lemma 3.5.
 The estimates for $\alpha$ and $r^{-1}\hat{w}$ are enhanced as follows
-$\alpha^{2}\lesssim{}$ $r^{b}\quad\text{and}\quad\alpha^{2}\lesssim\varepsilon,$
-$r^{-2}\hat{w}^{2}\lesssim{}$ $r^{b}\quad\text{and}\quad r^{-2}\hat{w}^{2}\lesssim\varepsilon^{2}.$
-Remark 3.6.
-If $b>0$, Lemma 3.5 suggests
-$\alpha|_{\Gamma}=0\quad\text{and}\quad r^{-1}\hat{w}|_{\Gamma}=0.$
+$r^{-2}\hat{w}^{2}\lesssim\varepsilon^{2}\quad\text{and}\quad\alpha^{2}\lesssim\varepsilon,$
+and
+$r^{-2}\hat{w}^{2}|_{\Gamma}=0\quad\text{and}\quad\alpha^{2}|_{\Gamma}=0.$
 Proof of Lemma 3.5.
@@
 $+\int_{\underline{C}_{\underline{u}}}2h^{-1}\left(\mu+\Phi^{2}\right)\frac{\alpha^{2}}{r}\mathrm{d}u$
-[…]
-where the bootstrap assumption (3.10) and the energy bound (3.1.4) are used. That is, we have derived a preliminary estimate for $\alpha$,
+$\lesssim{}$ $\underline{E}_{0}^{\prime}\lesssim I_{1},$
+where the bootstrap assumption (3.8) and the energy bound (3.5) are used. That is, we have derived a preliminary estimate for $\alpha$,
 $|\alpha|^{2}\lesssim\alpha^{2}|_{C_{u_{0}}}+I_{1}.$
@@
 $r\alpha^{2}={}$ $\int_{0}^{r}\left(\alpha^{2}+2\alpha\alpha^{\prime}\right)\mathrm{d}r^{\prime}|_{C_{u}}$
-[…]
-[…]
-Here in the last inequality, we have access to the equivalence (3.22) and the energy bound (3.1.4). As a result,
-$\alpha^{2}\lesssim\mu+\mu^{\frac{1}{2}}r^{\frac{b}{2}}.$
-Combined with the smallness of $\mu$ and Lemma 3.4, the first part of Lemma 3.5 is justified.
+$\leq{}$ $\int_{0}^{r}\alpha^{2}\mathrm{d}r^{\prime}|_{C_{u}}+\left(\int_{0}^{r}\alpha^{2}\mathrm{d}r^{\prime}|_{C_{u}}\right)^{\frac{1}{2}}\left(\int_{0}^{r}(\alpha^{\prime})^{2}\mathrm{d}r^{\prime}|_{C_{u}}\right)^{\frac{1}{2}}$
+$\lesssim{}$ $m+m^{\frac{1}{2}}r^{\frac{1}{2}}I_{1}^{\frac{1}{2}}.$
+Here in the last inequality, we have access to the equivalence (3.20) and the energy bound (3.5). As a result,
+$\alpha^{2}\lesssim\mu+\mu^{\frac{1}{2}}.$
+Combined with Lemma 3.4, the first part of Lemma 3.5 is justified.
 The estimate for $r^{-1}\hat{w}$ (or $\Phi$) follows in an analogous fashion.
@@
 $\frac{\hat{w}^{2}}{r^{2}}\Big{|}_{u_{0}}^{u}={}$ $\int_{\underline{C}_{\underline{u}}}\left(\partial_{u}w\frac{2\hat{w}}{r^{2}}-2\underline{h}\frac{\hat{w}^{2}}{r^{3}}\right)\mathrm{d}u$
-[…]
+$\lesssim{}$ $\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r}\left(h^{2}(\partial_{u}w)^{2}+\frac{\hat{w}^{2}}{r^{2}}\right)\mathrm{d}u\lesssim\underline{E}_{0}\lesssim I_{0}.$
 Therefore, $r^{-2}\hat{w}^{2}$ is bounded as
 $r^{-2}\hat{w}^{2}\lesssim r^{-2}\hat{w}^{2}\Big{|}_{C_{u_{0}}}+I_{0},$
-which entails $r\cdot\left(r^{-2}\hat{w}^{2}\right)|_{\Gamma}=0$.
+which further entails $r\cdot\left(r^{-2}\hat{w}^{2}\right)|_{\Gamma}=0$.
 It enables us to integrate $\partial_{r}\left(r\cdot\left(r^{-2}\hat{w}^{2}\right)\right)$ in $r$ on each $C_{u}$ and then an application of the Cauchy-Schwarz inequality leads to
 $r\frac{\hat{w}^{2}}{r^{2}}={}$ $\int_{0}^{r}\left(\frac{2\alpha\hat{w}}{r}-\frac{\hat{w}^{2}}{r^{2}}\right)\mathrm{d}r^{\prime}|_{C_{u}}$
-[…]
+$\lesssim{}$ $\int_{0}^{r}\left(2\alpha^{2}+r^{-2}\hat{w}^{2}\right)\mathrm{d}r^{\prime}|_{C_{u}}\lesssim m.$
 We arrive at
@@
 and thus obtain the concluding estimate for $r^{-1}\hat{w}$ with the help of Lemma 3.4. ∎
-With Lemma 3.5 concluded, we have improved the pointwise estimates for $r^{-1}\hat{w}$ and $\alpha$ and hence close the bootstrap argument.
+With Lemma 3.5 concluded, we have improved the pointwise estimates for $r^{-1}\hat{w}$ and $\alpha$, thereby closing the bootstrap argument.
 In the subsequent analysis, we will derive more precise estimates to provide a detailed description for the solution.
-3.1.6. Estimates for the geometry
+3.7. Estimates for the geometry
 Lemma 3.6.
@@
 Proof.
-Before going into the pointwise estimates, we will deduce some integral bounds for $\mu$.
+Before going into the pointwise estimates, we deduce some integral bounds for $\mu$.
 Denote $r_{0}=r(u_{0},\,\underline{u})$. Based on (2.14b), an integration in $u$ along each $\underline{C}_{\underline{u}}$ yields
 $\mu|_{r}^{r_{0}}+\int_{r}^{r_{0}}\frac{\mu}{r^{\prime}}\mathrm{d}r^{\prime}|_{\underline{C}_{\underline{u}}}$
-[…]
+$={}$ $\int_{\underline{C}_{\underline{u}}}\left(2h^{2}\frac{(\partial_{u}w)^{2}}{r}+\left(1-\mu\right)\frac{\Phi^{2}}{r}\right)h^{-1}\mathrm{d}u\lesssim I_{0},$ (3.21)
 where the energy bound (3.6) is employed in the last inequality. Since $\mu\leq\varepsilon^{2}$, we thus obtain
@@
 $\mu|_{0}^{r}+\int_{0}^{r}\frac{\mu}{r^{\prime}}\mathrm{d}r^{\prime}|_{C_{u}}$
-[…]
+$={}$ $\int_{0}^{r}\left(2\left(1-\mu\right)\frac{\alpha^{2}}{r^{\prime}}+\frac{\Phi^{2}}{r^{\prime}}\right)\mathrm{d}r^{\prime}|_{C_{u}}\lesssim I_{0},$ (3.23)
 which leads to
@@
 $\partial_{\underline{u}}\ln h=-2\Omega^{2}h\frac{\alpha^{2}}{r},$
-and as a consequence of the energy bound (3.8),
+and as a consequence of the energy bound, there is (3.6),
 $|\ln h|_{\Gamma}-\ln h|\lesssim\int_{0}^{r}\frac{\alpha^{2}}{r^{\prime}}\mathrm{d}r^{\prime}|_{C_{u}}\lesssim I_{0}^{2}.$
-In view of Lemma 3.4, $01+h\underline{h}|_{\Gamma}=\mu|_{\Gamma}=$ 0, and hence $h|_{\Gamma}=-\underline{h}^{-1}|_{\Gamma}=(\Omega^{2}h)^{-1}|_{\Gamma}$, the estimate for $\ln h$ is proved.
-In particular, if $b>0$, we can further show the solution $w\in C^{1}$.
-Lemma 3.8.
-In the case of $0<b<\frac{\sqrt{89}-9}{4}$, we have the following estimate for $\partial_{u}w$,
-$|\partial_{u}w|^{2}\lesssim r^{b}.$
-Proof.
-The fact $u=\underline{u}$ on $\Gamma$ and the boundary condition $w|_{\Gamma}=\pm 1$, plus Remark 3.6 entail
-$\partial_{u}w|_{\Gamma}=-\partial_{\underline{u}}w|_{\Gamma}=0.$
-Now we can integrate the YM equation (2.5) from the axis and use the $L^{\infty}$ estimate for $\Phi$ (Lemma 3.5) to conclude
-$|\partial_{u}w|\lesssim{}$ $|\partial_{u}w|_{\Gamma}+\int_{0}^{r}h^{-1}|w|\frac{|\Phi|}{r^{\prime}}\mathrm{d}r^{\prime}|_{C_{u}}$
-[…]
-3.1.7. Proof of Theorem 3.1
+In view of Lemma 3.4, $1+h\underline{h}|_{\Gamma}=\mu|_{\Gamma}=0$, and hence $h|_{\Gamma}=-\underline{h}^{-1}|_{\Gamma}=(\Omega^{2}h)^{-1}|_{\Gamma}$, the estimate for $\ln h$ is proved. ∎
+3.8. Proof of Theorem 3.1
 For the SSEYM system, the first singularity must emanate from the axis.
-One may refer to Theorem B.2 for more details.
-Therefore, it suffices to focus on a small neighbourhood of the tip $u=\underline{u}=\underline{u}_{\ast}$, and we may assume $u_{0}$ is close to $\underline{u}_{\ast}$ and
+For further details, we refer to Theorem A.2.
+It therefore suffices to focus on a small neighbourhood of the tip $u=\underline{u}=\underline{u}_{\ast}$.
+We may assume $u_{0}$ is close to $\underline{u}_{\ast}$ and that
 $\mu\leq\varepsilon,\quad\text{on}\,\,\mathcal{D}(u_{0};\,\underline{u}_{\ast})$
-As suggested by Lemma 3.2, the solution ($\alpha$) extends as an $H^{1}$ solution on the closure of the domain $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$.
-To complete the proof Theorem 3.1, we have to show that the solution can be extended beyond $\underline{C}_{\underline{u}_{\ast}}$.
-More explicitly, we shall verify that the solution can be extended to $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\,\underline{u}_{\ast}+\zeta)$ if $\zeta>0$ is small enough.
+As indicated by Lemma 3.2, the solution ($\alpha$) extends as an $H^{1}$ function to the closure of domain $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$.
+To complete the proof Theorem 3.1, it remains to show that the solution can be extended beyond $\underline{C}_{\underline{u}_{\ast}}$.
+Specifically, we must demonstrate that the solution can be extended to $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\,\underline{u}_{\ast}+\zeta)$ for some sufficiently small $\zeta>0$.
 We assume that
-$\mu\leq\frac{\varepsilon_{1}+\varepsilon}{2}\,\,\text{on}\,\,\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast}+\zeta),\quad\text{for any}\,\,u_{1}\in(u_{0},\,\underline{u}_{\ast}],$ (3.27)
+$\mu\leq\frac{\varepsilon_{1}+\varepsilon}{2}\,\,\text{in}\,\,\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast}+\zeta),\quad\text{for any}\,\,u_{1}\in(u_{0},\,\underline{u}_{\ast}],$ (3.25)
 where $\varepsilon_{1}>\varepsilon$ is a small constant.
 Define
-$u_{2}=\sup\,\{u_{1}\in(u_{0},\,\underline{u}_{\ast}]:\,\,\sup_{\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast}+\zeta)}\mu\leq\frac{\varepsilon_{1}+\varepsilon}{2}\}.$
+$u_{2}=\sup\,\left\{u_{1}\in(u_{0},\,\underline{u}_{\ast}]:\,\,\sup_{\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast}+\zeta)}\mu\leq\frac{\varepsilon_{1}+\varepsilon}{2}\right\}.$
 Then either $u_{2}=\underline{u}_{\ast}$ or $u_{2}<\underline{u}_{\ast}$ and $\sup_{\mathcal{D}(u_{0},\,u_{2};\,\underline{u}_{\ast}+\zeta)}\mu=\frac{\varepsilon_{1}+\varepsilon}{2}$.
@@
 This will be proved by showing that the bootstrap assumption (3.25) can be improved as
-$\mu<\frac{\varepsilon_{1}+\varepsilon}{2},\quad\text{on}\,\,\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast}+\zeta),\quad\text{for any}\,\,u_{1}\in(u_{0},\,\underline{u}_{\ast}]$
+$\mu<\frac{\varepsilon_{1}+\varepsilon}{2},\quad\text{in}\,\,\mathcal{D}(u_{0},\,u_{1};\,\underline{u}_{\ast}+\zeta),\quad\text{for any}\,\,u_{1}\in(u_{0},\,\underline{u}_{\ast}],$
 if $\zeta$ is small enough.
-Preliminarily, by virtue of the local existence theorem (Appendix A) and Lemma 3.2, the bootstrap assumption (3.27) leads to an $H^{1}$ solution on $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\,\underline{u}_{\ast}+\zeta)$.
+Preliminarily, by virtue of the local existence theorem (Section 4) and Lemma 3.2, the bootstrap assumption (3.25) leads to an $H^{1}$ solution in $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\,\underline{u}_{\ast}+\zeta)$.
 In particular, the $H^{1}$ norm
@@
 is uniformly bounded.
-On any $C_{u}$, $u\in[u_{0},\,\underline{u}_{\ast}]$, an integration in $\underline{u}^{\prime}\in[\underline{u}_{\ast},\,\underline{u}]$ as (3.1.6) yields the identity
-$\mu\Big{|}_{\underline{u}_{\ast}}^{\underline{u}}+\int_{r(u,\,\underline{u}_{\ast})}^{r(u,\,\underline{u})}\frac{\mu}{r}\mathrm{d}r|_{C_{u}}=\int_{\underline{u}_{\ast}}^{\underline{u}}\frac{\Omega^{2}h}{r}\left(\left(1-\mu\right)\alpha^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}^{\prime},$
+On any $C_{u}$, $u\in[u_{0},\,\underline{u}_{\ast}]$, an integration in $\underline{u}^{\prime}\in[\underline{u}_{\ast},\,\underline{u}]$ as (3.7) yields the identity
+$\mu\Big{|}_{\underline{u}_{\ast}}^{\underline{u}}+\int_{r(u,\,\underline{u}_{\ast})}^{r(u,\,\underline{u})}\frac{\mu}{r}\mathrm{d}r|_{C_{u}}=\int_{\underline{u}_{\ast}}^{\underline{u}}\frac{\Omega^{2}h}{r}\left(2\left(1-\mu\right)\alpha^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}^{\prime},$
 where $\underline{u}\in[\underline{u}_{\ast},\,\underline{u}_{\ast}+\zeta]$. Then $\mu$ is bounded as
-$\mu(u,\,\underline{u})<\mu(u,\,\underline{u}_{\ast})+\int_{\underline{u}_{\ast}}^{\underline{u}}\frac{\Omega^{2}h}{r}\left(\alpha^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}^{\prime},\quad\underline{u}\in[\underline{u}_{\ast},\,\underline{u}_{\ast}+\zeta].$
+$\mu(u,\,\underline{u})<\mu(u,\,\underline{u}_{\ast})+\int_{\underline{u}_{\ast}}^{\underline{u}}\frac{\Omega^{2}h}{r}\left(2\alpha^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}^{\prime},\quad\underline{u}\in[\underline{u}_{\ast},\,\underline{u}_{\ast}+\zeta].$
 Due to the uniform boundedness of the $H^{1}$ norm (3.26), we can take $\zeta$ small enough, such that
-$\int_{\underline{u}_{\ast}}^{\underline{u}}\frac{\Omega^{2}h}{r}\left(\alpha^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}^{\prime}<\frac{\varepsilon_{1}-\varepsilon}{2},\quad\forall\,\underline{u}\in[\underline{u}_{\ast},\,\underline{u}_{\ast}+\zeta].$
+$\int_{\underline{u}_{\ast}}^{\underline{u}}\frac{\Omega^{2}h}{r}\left(\alpha^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}^{\prime}<\frac{\varepsilon_{1}-\varepsilon}{2},\quad\underline{u}\in[\underline{u}_{\ast},\,\underline{u}_{\ast}+\zeta].$
 Thus, making use of the bootstrap assumption (3.25) improves the upper bound for $\mu$
@@
 In summary, we have proved (3.25) holds with $u_{1}=\underline{u}_{\ast}$.
-As a result, by the proof leading to Lemma 3.2, we know that the solution extends as an $H^{1}$ solution on $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\,\underline{u}_{\ast}+\zeta)$.
-Combined with the local existence theorem (Appendix A) for $H^{1}$ data, it further allows us to extend the solution to $\mathcal{D}(\underline{u}_{\ast};\,\underline{u}_{\ast}+\zeta^{\prime})$ for some $\zeta^{\prime}\in(0,\,\zeta]$.
-We complete the proof of Theorem 3.1.
+Consequently, following the argument leading to Lemma 3.2, the solution extends as an $H^{1}$ solution to the domian $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\,\underline{u}_{\ast}+\zeta)$.
+Together with the local existence theorem (Section 4) for $H^{1}$ data, it allows us to further extend the solution to $\mathcal{D}(\underline{u}_{\ast};\,\underline{u}_{\ast}+\zeta^{\prime})$ for some $\zeta^{\prime}\in(0,\,\zeta]$.
+This completes the proof of Theorem 3.1.
 3.9. Cauchy stability in $H^{1}$ space
-Having established the extension theorem 3.1, we will justify the Cauchy stability within the same topology.
-This amounts to proving a small data global existence in the finite domain $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\underline{u}_{\ast})$, where $\underline{u}_{\ast}\in(u_{0},\,+\infty)$.
-Recalling that the energies $E_{i}$, $i=0,1$, and $\underline{E}_{0}$, $\underline{E}_{0}^{\prime}$, are defined as in (3.1), we now state the Cauchy stability theorem.
+Having established the extension theorem (Theorem 3.1), we now turn to Cauchy stability within the same topology.
+This amounts to proving a small-data global existence on the finite domain $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\underline{u}_{\ast})$, where $\underline{u}_{\ast}\in(u_{0},\,+\infty)$.
+Recall that the energies $E_{i}$, $i=0,1$, and $\underline{E}_{0}$, $\underline{E}_{0}^{\prime}$, are defined as in (3.1).
+We now state the Cauchy stability theorem.
 Theorem 3.7.
@@
 $E_{1}(u_{0},\,\underline{u}_{\ast})<\varepsilon^{2},$
-with $\varepsilon$ being small enough.
-Case I: $0<b<\frac{\sqrt{89}-9}{4}$. There is a unique solution $\left(r,m,\alpha\right)\in C^{1}\times C^{1}\times H^{1}$ on $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\underline{u}_{\ast})$ such that
-$r^{-2}|\hat{w}|^{2}+|\alpha|^{2}+|\partial_{u}w|^{2}+\mu\lesssim\varepsilon^{2}r^{b},$
-and for $u_{0}<u\leq\underline{u}_{\ast}$, $u_{0}<\underline{u}\leq\underline{u}_{\ast}$,
+with $\varepsilon$ sufficiently small.
+Then there exists a unique and global solution $\left(r,m,\alpha\right)\in C^{1}\times C^{0}\times H^{1}$ in the domian $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\underline{u}_{\ast})$ such that
+$r^{-2}|\hat{w}|^{2}+|\alpha|^{2}+\mu\lesssim\varepsilon^{2},$
+and for $u_{0}<u\leq\underline{u}_{\ast}$, $u_{0}<\underline{u}\leq\underline{u}_{\ast}$, the solution satisfies the following unifom energy bound
 $E_{1}(u,\,\underline{u})+\underline{E}^{\prime}_{0}(u,\,\underline{u})\lesssim\varepsilon^{2}.$ (3.27)
-Case II: $b=0$. There is a unique solution $\left(r,m,\alpha\right)\in C^{1}\times C^{0}\times H^{1}$ on $\mathcal{D}(u_{0},\,\underline{u}_{\ast};\underline{u}_{\ast})$ such that
-$r^{-2}|\hat{w}|^{2}+|\alpha|^{2}+\mu\lesssim\varepsilon^{2},$
-and (3.29) holds as well.
 Proof.
-The proof goes through in line with the one leading to Theorem 3.1 and thus we omit the details.
-4. The weighted $H^{1}$ case
-Given two constants $u_{0}$ and $\underline{u}_{\ast}$ with $u_{0}<\underline{u}_{\ast}$, for any $\delta>0$, we will consider in the domain $\mathcal{D}(u_{0},\,u;\underline{u}_{\ast}+\delta)$ with $u\in[u_{0},\,\underline{u}_{\ast})$, a class of weighted energies.
-The weighted norm is defined through a scale invariant function $\xi$ (1.8).
-Indeed, we will prove later in Proposition 4.4 that in the domain concerned,
-$10\leq\xi<$ 1 .
-Before the statement of the main theorem, let us introduce some additional notations for this section. We denote the strip
-$\mathcal{S}(u,\,\underline{u}):=[u_{0},\,u]\times[\underline{u}_{\ast},\,\underline{u}],\quad u\in(u_{0},\,\underline{u}_{\ast}),\,\,\underline{u}\in(\underline{u}_{\ast},\,\underline{u}_{\ast}+\delta],$ (4.1)
-and $\mathcal{S}(u,\,\underline{u})$ will be simplified as $\mathcal{S}$.
-Besides, we abbreviate $C_{u}^{[\underline{u}_{\ast},\,\underline{u}]}$ by $C_{u}$, and $\underline{C}_{\underline{u}}^{[u_{0},\,u]}$ by $\underline{C}_{\underline{u}}$.
-We also denote the value of $m$ at $\underline{u}=\underline{u}_{\ast}$ by $m_{\ast}$.
-Similar conventions apply to $\underline{h}_{\ast},\,\mu_{\ast},\cdots$.
-Fix $b$ as in (1.4), and a constant $\gamma\in(b,2]$. We define the $\xi$ weighted energy norms
-$\|\alpha\|^{2}_{H^{0}_{w}(C_{u})}:={}$ $\int_{C_{u}}\frac{\Omega^{2}h}{r^{1+b}}\left(\frac{r^{-2}\hat{w}^{2}}{\xi^{3+\gamma}}+\frac{\alpha^{2}}{\xi^{1+\gamma}}\right)\mathrm{d}\underline{u},$ (4.2a)
-$\|\hat{w}\|^{2}_{H^{1}_{w}(\underline{C}_{\underline{u}})}:={}$ $\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(\frac{(\partial_{u}w)^{2}}{\xi^{3+\gamma}}+\frac{r^{-2}\hat{w}^{2}}{\xi^{1+\gamma}}\right)\mathrm{d}u,$ (4.2b)
-$\|\alpha\|^{2}_{H^{1}_{w}(C_{u})}:={}$ $\|\alpha\|^{2}_{H^{0}_{w}(C_{u})}+\int_{C_{u}}\xi^{1-\gamma}\frac{(\alpha^{\prime})^{2}}{r^{1+b}}\Omega^{2}h\mathrm{d}\underline{u},$ (4.2c)
-$\|\hat{w}\|^{2}_{H^{1^{\prime}}_{w}(\underline{C}_{\underline{u}})}:={}$ $\|\hat{w}\|^{2}_{H^{1}_{w}(\underline{C}_{\underline{u}})}+\int_{\underline{C}_{\underline{u}}}\xi^{1-\gamma}\frac{\alpha^{2}}{r^{1+b}}h^{-1}\mathrm{d}u.$ (4.2d)
-4.1. Extension theorem in the weighted $H^{1}$ space
+The proof goes through in line with the one leading to Theorem 3.1, and we therefore omit the details here.
+4. Local existence theorem
+For the characteristic initial value problem with data prescribed on a future light cone emanating from the center of symmetry, the local existence theorem for classical solutions to the spherically symmetric Einstein scalar field system was proved in [8].
+Regarding to the proof, Christodoulou reduced the whole system in Bondi coordinates to a scalar equation and utilized the contracting mapping principle to obtain the existence theorem.
+In this section, however, we aim to present a proof based on energy methods.
+In more details, we work in the double null gauge and demonstrate a local existence theory for the SSEYM system in $H^{1}$, using an iteration method.
+The Yang–Mills equation can be alternatively written in terms of $\hat{w}$ (1.3)
+$\partial_{u}\partial_{\underline{u}}\hat{w}+2\frac{\Omega^{2}}{r^{2}}\hat{w}=\Omega^{2}\frac{\hat{w}^{2}}{r^{2}}\left(\hat{w}+3\right).$ (4.1)
 Theorem 4.1.
-Fix a number $b$ as in (1.4), and a constant $\gamma\in(b,2]$.
-Given datum $\alpha|_{C_{u_{0}}}$ that is supported in $(\underline{u}_{\ast},\,+\infty)$ (namely, $\alpha|_{C_{u_{0}}\cap\{\underline{u}\leq\underline{u}_{\ast}\}}=0$), we assume the initial energy
-$\|\alpha\|^{2}_{H^{1}_{w}(C^{[\underline{u}_{\ast},\,\underline{u}_{\ast}+\delta]}_{u_{0}})}$
-is bounded for some $\delta>0$, and $|r-r_{\ast}|^{-\frac{\gamma}{2}}\alpha|_{C_{u_{0}}}$ is continuous at $\underline{u}=\underline{u}_{\ast}$. If $\delta$ is small enough, we might suppose
-$\|\alpha\|^{2}_{H^{1}_{w}(C^{[\underline{u}_{\ast},\,\underline{u}_{\ast}+\delta]}_{u_{0}})}\leq\varepsilon^{2},$
-where $\varepsilon$ is small enough (depending on $\delta$).
-Case I: $b>0$.
-If $\delta$ is small enough, the solution $(r,m,\alpha)$ extends as a $C^{1}\times C^{1}\times H^{1}_{w}$ solution in the domain of $\mathcal{D}(u_{0},\,u;\,\underline{u}_{\ast}+\delta)$ with $u\in[u_{0},\,\underline{u}_{\ast})$.
-Moreover, for each $u\in[u_{0},\,\underline{u}_{\ast})$, $\underline{u}\in[\underline{u}_{\ast},\,\underline{u}_{\ast}+\delta]$, the solution admits the energy bound
-$\|\hat{w}\|^{2}_{H^{1\prime}_{w}(\underline{C}_{\underline{u}})}+\|\alpha\|^{2}_{H^{1}_{w}(C_{u})}\lesssim\varepsilon^{2}.$
-In the strip $\mathcal{S}(u,\,\underline{u}_{\ast}+\delta)$, $u\in[u_{0},\,\underline{u}_{\ast})$, there is the pointwise estimate
-$\xi^{-2-\gamma}r^{-2}\hat{w}^{2}+\xi^{-\gamma}\alpha^{2}+\xi^{-3-\gamma}|\partial_{u}w|^{2}+\xi^{-1-\gamma}\mu\lesssim\varepsilon^{2}r^{b}.$
-Moreover, if $|(r-r_{\ast})^{1-\frac{\gamma}{2}}\alpha^{\prime}|_{L^{\infty}(C^{[\underline{u}_{\ast},\,\underline{u}_{\ast}+\delta]}_{u_{0}})}$ is bounded, then
-$\xi^{2-\gamma}(\alpha^{\prime})^{2}\lesssim r^{b},\quad\text{in}\,\,\mathcal{S}(u,\,\underline{u}_{\ast}+\delta),\,\,u\in[u_{0},\,\underline{u}_{\ast}).$ (4.3)
-Case II: $b=0$.
-If $\delta$ is small enough, the solution $(r,m,\alpha)$ extends as a $C^{1}\times C^{0}\times H^{1}_{w}$ solution in the domain of $\mathcal{D}(u_{0},\,u;\,\underline{u}_{\ast}+\delta)$ with $u\in[u_{0},\,\underline{u}_{\ast})$.
-The estimates as in Case I, except the $L^{\infty}$ estimate for $\partial_{u}w$, hold as well with $b=0$.
-Remark 4.2.
-Since the data $\alpha|_{C_{u_{0}}}\equiv 0$ in $\underline{u}\leq\underline{u}_{\ast}$, the domain $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$ is flat and hence
-$m=0,\quad\text{on}\,\,\mathcal{D}(u_{0};\,\underline{u}_{\ast}).$
-Therefore, it suffices to establish the energy estimates in the strip $\mathcal{S}(u,\,\underline{u}_{\ast}+\delta)$ with $u\in[u_{0},\,\underline{u}_{\ast})$.
-In the proof, we have made the following continuity assumptions on the flat boundary cone $\underline{C}_{\underline{u}_{\ast}}$.
-Since the domain $\mathcal{D}(u_{0};\,\underline{u}_{\ast})$ is flat and hence
-$\lim_{\underline{u}\rightarrow\underline{u}_{\ast}^{-}}\int_{\underline{C}_{\underline{u}}}\left(|\xi|^{-1-\gamma}h^{-1}\frac{\Phi^{2}}{r^{1+b}}+|\xi|^{-3-\gamma}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\right)\mathrm{d}u=0,$
-$\lim_{\underline{u}\rightarrow\underline{u}_{\ast}^{-}}|\xi|^{-1-\gamma}\frac{\mu}{r^{b}}+\xi^{-2-\gamma}\frac{\Phi^{2}}{r^{b}}+\xi^{-3-\gamma}\frac{(\partial_{u}w)^{2}}{r^{b}}=0.$
-We assume that these quantities are continuous at $\underline{u}=\underline{u}_{\ast}$. That is, we assume
-$\lim_{\underline{u}\rightarrow\underline{u}_{\ast}^{+}}\int_{\underline{C}_{\underline{u}}}\left(|\xi|^{-1-\gamma}h^{-1}\frac{\Phi^{2}}{r^{1+b}}+|\xi|^{-3-\gamma}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\right)\mathrm{d}u=0,$
-$\lim_{\underline{u}\rightarrow\underline{u}_{\ast}^{+}}|\xi|^{-1-\gamma}\frac{\mu}{r^{b}}+\xi^{-2-\gamma}\frac{\Phi^{2}}{r^{b}}+\xi^{-3-\gamma}\frac{(\partial_{u}w)^{2}}{r^{b}}=0.$
-Remark 4.3.
-The estimate (4.3) indicates the singularity emanating from the initial sphere $S_{\underline{u}_{\ast},u_{0}}$ (with singular $\alpha^{\prime}$) is propagated along the incoming cone $\underline{C}_{\underline{u}_{\ast}}$.
-We intend to carry out the proof of Theorem 4.1 in the remainder of this section.
-4.1.1. Bootstrap assumptions
-Let $\Lambda$ be a large constant that will be determined (depending on the initial energy). We suppose in $\mathcal{S}$ there are
-$|r^{-1}\hat{w}|^{2}+|\alpha|^{2}+\mu\leq{}$ $\varepsilon^{2}\Lambda^{2},$ (4.4a)
-$|rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}|\leq{}$ $\varepsilon^{2}\Lambda^{2}.$ (4.4b)
-It follows from the bootstrap assumption for $\hat{w}$ (4.4a) that
-$w^{2}\lesssim 1+\varepsilon\Lambda,\quad\text{in}\,\,\mathcal{S}.$ (4.5)
-Using (2.4b) and the bootstrap assumption for $\mu$, we have, as stated in (3.9), that
-$\underline{h}<0,\quad h>0,\quad\text{in}\,\,\mathcal{S}.$
-In regard to the weight $\xi$, we list some basic facts followed from the bootstrap assumptions.
-Proposition 4.4.
-We have the identities concerning the derivatives of $\xi$
-$r\partial_{\underline{u}}\xi=(1-\xi)\Omega^{2}h,$
-$rh\frac{\partial_{u}\xi}{\xi}+\underline{h}h=rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}.$
-As a result, under the bootstrap assumption (4.4a), there is
-$0\leq\xi<1,\quad\partial_{u}\xi\geq 0,\quad\partial_{\underline{u}}\xi>0,\quad\text{in}\,\,\mathcal{S}.$
+Let $\underline{u}_{1}\in(u_{0},\,+\infty)$ be a finite constant. Suppose on the initial slice $C_{u_{0}}$, the energy is bounded
+$\int_{C_{u_{0}}}r^{-1}\left(\frac{\hat{w}^{2}}{r^{2}}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}r<I_{1}.$
+i) There exists $u_{\ast}>u_{0}$ such that the SSEYM system admits a unique solution $\left(r,\,m,\,\alpha\right)\in C^{1}\times C^{0}\times H^{1}$ in the domain $\mathcal{D}(u_{0},u_{\ast};\underline{u}_{1})$, satisfying the bound,
+$\int_{C_{u}}r^{-1}\left(\frac{\hat{w}^{2}}{r^{2}}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}r\lesssim I_{1},$
+$\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r}\left(\frac{\hat{w}^{2}}{r^{2}}+\alpha^{2}+h^{2}(\partial_{u}\hat{w})^{2}\right)\mathrm{d}u\lesssim I_{1},$
+$|r^{-1}\hat{w}|^{2}+|\alpha|^{2}+\mu\lesssim I_{1},$
+$|\ln\left(\Omega^{2}h\right)|,\,|\ln\left(-\underline{h}\right)|,\,|\ln h|\lesssim I_{1}.$
+ii) The solution depends continuously on the initial data in the following sense.
+Suppose $\left(r_{n},\,m_{n},\,\hat{w}_{n}\right)$ and $\left(r_{m},\,m_{m},\,\hat{w}_{m}\right)$ are solutions with data $\hat{w}_{0n}$, $\hat{w}_{0m}$ respectively, then
+$\int_{C_{u}}\left(\frac{|\hat{w}_{n}-\hat{w}_{m}|^{2}}{r^{3}_{n}}+\frac{|\alpha_{n}-\alpha_{m}|^{2}}{r_{n}}+\frac{|\alpha^{\prime}_{n}-\alpha^{\prime}_{m}|^{2}}{r_{n}}\right)\mathrm{d}r_{n}$
+$+\int_{\underline{C}_{\underline{u}}}\frac{h_{n}^{-1}}{r_{n}}\left(\frac{\left(\hat{w}_{n}-\hat{w}_{m}\right)^{2}}{r_{n}^{2}}+\left(\alpha_{n}-\alpha_{m}\right)^{2}+h_{n}^{2}(\partial_{u}\hat{w}_{n}-\partial_{u}\hat{w}_{m})^{2}\right)\mathrm{d}u$
+$\lesssim{}$ $\int_{C_{u_{0}}}\left(\frac{|\hat{w}_{0n}-\hat{w}_{0m}|^{2}}{r^{3}_{n}}+\frac{|\alpha_{0n}-\alpha_{0m}|^{2}}{r_{n}}+\frac{|\alpha^{\prime}_{0n}-\alpha^{\prime}_{0m}|^{2}}{r_{n}}\right)\mathrm{d}r_{n}.$
 Proof.
-For any $u\in[u_{0},\,\underline{u}_{\ast})$, $\underline{u}\in[\underline{u}_{\ast},\,\underline{u}_{\ast}+\delta]$, there is always, in the strip $\mathcal{S}$, $r\geq r_{\ast}>0$ since $h>0$, and thus
-$10\leq\xi<$ 1 .
-Moreover, it follows that
-$\partial_{\underline{u}}\xi=\frac{\Omega^{2}hr_{\ast}}{r^{2}}=(1-\xi)\frac{\Omega^{2}h}{r}>0.$
-A computation yields $\partial_{u}\xi=\frac{\underline{h}-\underline{h}_{\ast}}{r}-\frac{\underline{h}\xi}{r}$, and hence
-$\frac{\partial_{u}\xi}{\xi}=\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}-\frac{\underline{h}}{r}.$
-Alternatively, we have the identity
-$\partial_{u}\xi=\frac{-\underline{h}_{\ast}}{r}+\frac{\underline{h}r_{\ast}}{r^{2}}=-\frac{r_{\ast}}{r}\left(\frac{\underline{h}_{\ast}}{r_{\ast}}-\frac{\underline{h}}{r}\right).$
-It is suggested from this formulation that $\partial_{u}\xi$ will be nonnegative if we can show that $\frac{\underline{h}}{r}$ is increasing in $\underline{u}$.
-In fact, using the equation (2.4c) and the bootstrap assumption (4.4a),
-$\partial_{\underline{u}}\frac{\underline{h}}{r}=\frac{\partial_{\underline{u}}\underline{h}}{r}-\frac{\Omega^{2}h\underline{h}}{r^{2}}=-\left(\mu-\Phi^{2}\right)\frac{\Omega^{2}}{r^{2}}+(1-\mu)\frac{\Omega^{2}}{r^{2}}>0,$
-and we finish the proof. ∎
-4.1.2. Weighted energy estimates
-To begin with, we multiply $\Omega^{-2}h^{-1}r^{-1-b}\xi^{-1-\gamma}$ on (3.12) and follow the procedures leading to (3.1.4).
-Consequently, we obtain an identity in the strip $\mathcal{S}$ for any $u\in[u_{0},\,\underline{u}_{\ast})$, $\underline{u}\in[\underline{u}_{\ast},\,\underline{u}_{\ast}+\delta]$,
-$\int_{C_{u}}\xi^{-1-\gamma}\Omega^{2}h\frac{\alpha^{2}}{r^{1+b}}\mathrm{d}\underline{u}+\int_{\underline{C}_{\underline{u}}}\xi^{-1-\gamma}h^{-1}\frac{\Phi^{2}}{2r^{1+b}}\mathrm{d}u$
-$+\iint_{\mathcal{S}}\xi^{-1-\gamma}\Omega^{2}\left(\frac{3+b}{2}\frac{\Phi^{2}}{r^{2+b}}-(1+b)\frac{\alpha^{2}}{r^{2+b}}+b\mu\frac{\alpha^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}$
-[…]
-$+\iint_{\mathcal{S}}\left(\partial_{u}\xi^{-1-\gamma}\Omega^{2}h\frac{\alpha^{2}}{r^{1+b}}+\partial_{\underline{u}}\xi^{-1-\gamma}h^{-1}\frac{\Phi^{2}}{2r^{1+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}.$
-Note that, the energy on the initial cone $\underline{C}_{\underline{u}_{\ast}}$, $\int_{\underline{C}_{\underline{u}_{\ast}}}\xi^{-1-\gamma}h^{-1}\frac{\Phi^{2}}{2r^{1+b}}\mathrm{d}u$, vanishes due to the continuity assumption on $\underline{C}_{\underline{u}_{\ast}}$, as shown in Remark 4.2.
-With the help of Proposition 4.4, the above energy identity is further calculated as
-$\int_{C_{u}}\xi^{-1-\gamma}\Omega^{2}h\frac{\alpha^{2}}{r^{1+b}}\mathrm{d}\underline{u}+\int_{\underline{C}_{\underline{u}}}\xi^{-1-\gamma}h^{-1}\frac{\Phi^{2}}{2r^{1+b}}\mathrm{d}u$
-$+\iint_{\mathcal{S}}\xi^{-1-\gamma}\Omega^{2}\left(\frac{2+b-\gamma}{2}\frac{\Phi^{2}}{r^{2+b}}+(\gamma-b)\frac{\alpha^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}$
-$+\iint_{\mathcal{S}}(1+\gamma)\Omega^{2}\xi^{-2-\gamma}\frac{\Phi^{2}}{2r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$
-[…]
-$-\iint_{\mathcal{S}}(1+\gamma)\xi^{-1-\gamma}\Omega^{2}\left(rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\right)\frac{\alpha^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}.$
-Since $\gamma\in(b,2]$, terms on the left hand side are all positive.
-At the same time, due to the bootstrap assumptions (4.4a)–(4.4b), the spacetime integrals on the right hand side have extra smallness and hence can be absorbed by the positive term $\iint_{\mathcal{S}}(\gamma-b)\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$ on the left side.
-Therefore, we conclude
-$\int_{C_{u}}\xi^{-1-\gamma}\Omega^{2}h\frac{\alpha^{2}}{r^{1+b}}\mathrm{d}\underline{u}+\int_{\underline{C}_{\underline{u}}}\xi^{-1-\gamma}h^{-1}\frac{\Phi^{2}}{2r^{1+b}}\mathrm{d}u$
-$+\iint_{\mathcal{S}}\xi^{-1-\gamma}\Omega^{2}\left(\frac{2+b-\gamma}{2}\frac{\Phi^{2}}{r^{2+b}}+\frac{\gamma-b}{2}\frac{\alpha^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}$
-$+\iint_{\mathcal{S}}(1+\gamma)\xi^{-2-\gamma}\Omega^{2}\frac{\Phi^{2}}{2r^{2+b}}\lesssim\int_{C_{u_{0}}}\xi^{-1-\gamma}\Omega^{2}h\frac{\alpha^{2}}{r^{1+b}}\mathrm{d}\underline{u}.$ (4.6)
-Meanwhile, multiplying $\xi^{-3-\gamma}hr^{-1-b}$ on (3.14) and following the proof leading to (3.1.4), we derive another energy identity
-$\int_{\underline{C}_{\underline{u}}}\xi^{-3-\gamma}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\mathrm{d}u+\int_{C_{u}}\xi^{-3-\gamma}\Omega^{2}h\frac{\Phi^{2}}{2r^{1+b}}\mathrm{d}\underline{u}$
-$+\iint_{\mathcal{S}}\xi^{-3-\gamma}\Omega^{2}\left((1+b)h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}-\frac{3+b}{2}\frac{\Phi^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}$
-$+\iint_{\mathcal{S}}\xi^{-3-\gamma}\Omega^{2}\left(2\alpha^{2}h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}+\left(2+\frac{b}{2}\right)\mu\frac{\Phi^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}$
-[…]
-$+\iint_{\mathcal{S}}\left(\partial_{\underline{u}}\xi^{-3-\gamma}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}+\partial_{u}\xi^{-3-\gamma}\Omega^{2}h\frac{\Phi^{2}}{2r^{1+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}.$
-Note that, we have used the fact that the energy $\int_{\underline{C}_{\underline{u}_{\ast}}}\xi^{-3-\gamma}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\mathrm{d}u$ vanishes, in view of Remark 4.2.
-Then substituting the calculations in Proposition 4.4 into the last line of the above identity, we have
-$\int_{\underline{C}_{\underline{u}}}\xi^{-3-\gamma}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\mathrm{d}u+\int_{C_{u}}\xi^{-3-\gamma}\Omega^{2}h\frac{\Phi^{2}}{2r^{1+b}}\mathrm{d}\underline{u}$
-$+\iint_{\mathcal{S}}\xi^{-3-\gamma}\Omega^{2}\left((1+b)h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}+\frac{\gamma-b}{2}\frac{\Phi^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}$
-$+\iint_{\mathcal{S}}(3+\gamma)\xi^{-4-\gamma}\left(1-\xi\right)\Omega^{2}h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$
-$+\iint_{\mathcal{S}}2\xi^{-3-\gamma}\Omega^{2}\alpha^{2}h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$
-[…]
-$+\iint_{\mathcal{S}}(3+\gamma)\xi^{-3-\gamma}\Omega^{2}\left(\mu+\Phi^{2}+\big{|}rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\big{|}\right)\frac{\Phi^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}.$
-By the bootstrap assumptions (4.4a)–(4.4b), the spacetime integrals on the right hand side are in fact small error terms and can be absorbed by $\iint_{\mathcal{D}}\frac{\gamma-b}{2}\Omega^{2}\frac{\Phi^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$ on the left side.
-Consequently, we achieve the following energy bound
-$\int_{\underline{C}_{\underline{u}}}\xi^{-3-\gamma}h\frac{(\partial_{u}w)^{2}}{r^{1+b}}\mathrm{d}u+\int_{C_{u}}\xi^{-3-\gamma}\Omega^{2}h\frac{\Phi^{2}}{r^{1+b}}\mathrm{d}\underline{u}$
-$+\iint_{\mathcal{S}}\xi^{-3-\gamma}\Omega^{2}\left((1+b)h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}+\frac{\gamma-b}{4}\frac{\Phi^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}$
-$+\iint_{\mathcal{S}}(3+\gamma)\xi^{-4-\gamma}\left(1-\xi\right)\Omega^{2}h^{2}\frac{(\partial_{u}w)^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$
-[…]
-In particular, the energy estimates (4.1.2) and (4.1.2) indicates that
-$\|\alpha\|^{2}_{H^{0}_{w}(C_{u})}+\|\hat{w}\|^{2}_{H^{1}_{w}(\underline{C}_{\underline{u}})}\lesssim\|\alpha\|^{2}_{H^{0}_{w}(C_{u_{0}})}\leq\varepsilon^{2}.$
-This completes the first order weighted energy estimate.
-The second order weighted energy estimate follows from multiplying $2\xi^{1-\gamma}r^{-1-b}\cdot\Omega^{2}h\alpha^{\prime}$ on (3.1.4), and modifying the proof leading to (3.1.4).
-Moreover, using Proposition 4.4 and the bootstrap assumptions (4.4a)–(4.4b), we obtain
-$\int_{C_{u}}\xi^{1-\gamma}\frac{\Omega^{2}h}{r^{1+b}}\left(\Phi^{2}+(9+2b)\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}\underline{u}$
-$+\int_{\underline{C}_{\underline{u}}}\xi^{1-\gamma}\frac{h^{-1}}{r^{1+b}}\left(2\left(w\alpha-\Phi\right)^{2}+\frac{\Phi^{2}}{2}+2h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u$
-$+\iint_{\mathcal{S}}(\gamma-b)\xi^{1-\gamma}\frac{\Omega^{2}}{r^{2+b}}(\alpha^{\prime})^{2}\mathrm{d}u\mathrm{d}\underline{u}$
-[…]
-$+\iint_{\mathcal{S}}\varepsilon\Lambda\xi^{1-\gamma}\frac{\Omega^{2}}{r^{2+b}}(\alpha^{\prime})^{2}\mathrm{d}u\mathrm{d}\underline{u}+\iint_{\mathcal{S}}\varepsilon\Lambda\xi^{1-\gamma}\frac{\Omega^{2}}{r^{2+b}}\left(\alpha^{2}+\Phi^{2}\right)\mathrm{d}u\mathrm{d}\underline{u}$
-$+\iint_{\mathcal{S}}\xi^{-\gamma}\frac{\Omega^{2}}{r^{2+b}}\left(\xi\alpha^{2}+\left(w\alpha-\Phi\right)^{2}+\Phi^{2}+2h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u\mathrm{d}\underline{u}.$
-Due to the choice $\gamma>b$, the second term on the right hand side having the small constant $\varepsilon\Lambda$ can be absorbed by the spacetime integral on the left hand side.
-To bound the other spacetime integrals on the right hand side, we appeal to the first order estimates (4.1.2) and (4.1.2) so that we are able to deduce the energy bound
-$\int_{C_{u}}\xi^{1-\gamma}\frac{\Omega^{2}h}{r^{1+b}}\left(\Phi^{2}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}\underline{u}$
-$+\int_{\underline{C}_{\underline{u}}}\xi^{1-\gamma}\frac{h^{-1}}{r^{1+b}}\left(\left(w\alpha-\Phi\right)^{2}+\Phi^{2}+h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u$
-$+\iint_{\mathcal{S}}\frac{(\gamma-b)}{2}\xi^{1-\gamma}\frac{\Omega^{2}}{r^{2+b}}(\alpha^{\prime})^{2}\mathrm{d}u\mathrm{d}\underline{u}$
-[…]
-Once the energy estimates are completed, we are able to improve the pointwise estimates.
-4.1.3. Pointwise estimates
-In the following lemmas 4.54.7, we will improve the $L^{\infty}$ estimates for $r^{-1}\hat{w},\,\alpha$, $\mu$ and $\big{|}rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\big{|}$ to close the bootstrap argument.
-Lemma 4.5.
-We have the refine estimate
-$\xi^{-1-\gamma}\mu\lesssim\varepsilon^{2}r^{b},\quad\text{in}\,\,\mathcal{S}.$
-Proof.
-The proof is analogous to that of Lemma 3.4. By the mass formula,
-$\frac{m}{r^{1+b}\xi^{1+\gamma}}={}$ $\frac{1}{r^{1+b}\xi^{1+\gamma}}\int_{r_{\ast}}^{r}\left((1-\mu)\alpha^{2}+\Phi^{2}\right)\mathrm{d}r^{\prime}|_{C_{u}},$
-[…]
-[…]
-where in the second line we have used the fact $\partial_{r}\xi\geq 0$ along $C_{u}\cap\mathcal{S}$. ∎
-Lemma 4.6.
-We have in $\mathcal{S}$
-$\xi^{-\gamma}\alpha^{2}\lesssim\varepsilon^{2}r^{b},\quad\xi^{-2-\gamma}r^{-2}\hat{w}^{2}\lesssim\varepsilon^{2}r^{b}.$
-Proof.
-We will first show that $\|\alpha\|^{2}_{H^{1}_{w}(C_{u_{0}})}\lesssim\varepsilon^{2}$ implies
-$\|\xi^{-\gamma}r^{-b}\alpha^{2}\|_{L^{\infty}(C_{u_{0}}\cap\mathcal{S})}\lesssim\varepsilon^{2}.$ (4.9)
-Integrating along $\partial_{r}$ on $C_{u_{0}}$, we obtain the following identity
-$\xi^{-\gamma}r^{-b}\alpha^{2}\big{|}_{C_{u_{0}}}(\underline{u})-\xi^{-\gamma}r^{-b}\alpha^{2}\big{|}_{C_{u_{0}}}(\underline{u}_{\ast})$
-[…]
-Noting that $\xi^{-\gamma}r^{-b}\alpha^{2}\big{|}_{C_{u_{0}}}(\underline{u}_{\ast})=0$, we have
-$\xi^{-\gamma}r^{-b}\alpha^{2}\big{|}_{C_{u_{0}}}(\underline{u})+\int_{C_{u_{0}}}\gamma\frac{\alpha^{2}}{\xi^{1+\gamma}r^{1+b}}\mathrm{d}r$
-[…]
-[…]
-which justifies (4.9).
-Next, we prove pointwise estimate for $\alpha$, using (4.9).
-Multiplying $2\xi^{-\gamma}r^{-b}\alpha$ on (2.15) and integrating along $\partial_{u}$ on each $\underline{C}_{\underline{u}}$, noting that $\gamma>b$, we have
-$\big{|}\xi^{-\gamma}r^{-b}\alpha^{2}-\xi^{-\gamma}r^{-b}\alpha^{2}|_{u_{0}}\big{|}+\int_{\underline{C}_{\underline{u}}}\left(\gamma-b\right)(1-\mu)\xi^{-\gamma}\frac{\alpha^{2}}{r^{1+b}}h^{-1}\mathrm{d}u$
-[…]
-$+\int_{\underline{C}_{\underline{u}}}\gamma\Big{|}r\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\Big{|}\xi^{-\gamma}\frac{\alpha^{2}}{r^{1+b}}\mathrm{d}u$
-[…]
-$+\int_{\underline{C}_{\underline{u}}}\left(2\mu+\gamma\big{|}rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\big{|}\right)\xi^{-\gamma}\frac{\alpha^{2}}{r^{1+b}}h^{-1}\mathrm{d}u.$
-By the bootstrap assumptions, $\big{|}rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\big{|}$ and $\mu$ are both small enough, then
-$(\gamma-b)(1-\mu)>\mu+\gamma\big{|}rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\big{|}.$
-After absorbing the error terms on the right hand side, we conclude
-$\xi^{-\gamma}r^{-b}\alpha^{2}+\int_{\underline{C}_{\underline{u}}}\frac{\gamma-b}{2}\xi^{-\gamma}\frac{\alpha^{2}}{r^{1+b}}h^{-1}\mathrm{d}u$
-[…]
-where in the last inequality we have appealed to (4.9) and the energy estimates (4.1.2) and (4.1.2).
-Meanwhile, integrating $\xi^{-2-\gamma}r^{-b-2}\hat{w}^{2}$ along $\partial_{\underline{u}}$ on each $C_{u}$ gives rise to
-$\xi^{-2-\gamma}r^{-b-2}\hat{w}^{2}{}$ $+(2+\gamma)\int_{C_{u}}\xi^{-3-\gamma}(1-\xi)\frac{r^{-2}\hat{w}^{2}}{r^{1+b}}\Omega^{2}h\mathrm{d}\underline{u}$
-$+(2+b)\int_{C_{u}}\xi^{-2-\gamma}\frac{r^{-2}\hat{w}^{2}}{r^{1+b}}\Omega^{2}h\mathrm{d}\underline{u}$
-$\lesssim$ $\int_{C_{u}}\left(\xi^{-1-\gamma}\frac{\alpha^{2}}{r^{1+b}}+\xi^{-3-\gamma}\frac{r^{-2}\hat{w}^{2}}{r^{1+b}}\right)\Omega^{2}h\mathrm{d}\underline{u}\lesssim\varepsilon^{2}.$
-We complete the proof of this lemma. ∎
-Lemma 4.7.
-We have the improvement,
-$\Big{|}rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}\Big{|}\lesssim\varepsilon^{2},\quad\text{in}\,\,\mathcal{S}.$
-Proof.
-By means of a straightforward calculation, using the mean value theorem, we find that
-$rh\frac{\underline{h}-\underline{h}_{\ast}}{r-r_{\ast}}={}$ $\frac{rh\underline{h}-r_{\ast}h_{\ast}\underline{h}_{\ast}+r_{\ast}h_{\ast}\underline{h}_{\ast}-rh\underline{h}_{\ast}}{r-r_{\ast}}$
-[…]
-[…]
-[…]
-which is small in view of Lemma 4.6 and the fact
-$\big{|}\ln\frac{h|_{r_{2}}}{h_{\ast}}\big{|}\lesssim\int^{r_{2}}_{r_{\ast}}\frac{\alpha^{2}}{r}\mathrm{d}r\lesssim\varepsilon^{2}.$
-What is more, along the lines of the subsection 3.1.6, we are able to prove the boundedness of the following geometric quantities,
-$|\ln(\Omega^{2}h)|+|\ln(-\underline{h})|+|\ln h|.$
-In the end, we present the pointwise estimates for $\partial_{u}w$ and $\alpha^{\prime}$.
-Lemma 4.8.
-In the case of $0<b<\frac{\sqrt{89}-9}{4}$, there is
-$\xi^{-3-\gamma}|\partial_{u}w|^{2}\lesssim\varepsilon^{2}r^{b}.$
-The proof resembles the one in the second part of Lemma 4.6.
-Lemma 4.9.
-Suppose $\xi^{2-\gamma}r^{-b}(\alpha^{\prime})^{2}\big{|}_{C_{u_{0}}}$ is bounded, then we have the following $L^{\infty}$ estimate for $\alpha^{\prime}$,
-$\xi^{2-\gamma}(\alpha^{\prime})^{2}\lesssim r^{b}.$
-Proof.
-Multiply $2\xi^{2-\gamma}r^{-b}\alpha^{\prime}$ on the second order equation (3.1.4) and integrate along $\partial_{u}$. Then
-$\xi^{2-\gamma}r^{-b}(\alpha^{\prime})^{2}+\int_{\underline{C}_{\underline{u}}}(\gamma-b)\xi^{2-\gamma}\frac{(\alpha^{\prime})^{2}}{r^{1+b}}h^{-1}\mathrm{d}u$
-[…]
-$+\int_{\underline{C}_{\underline{u}}}\left(c+\varepsilon\right)\xi^{2-\gamma}\frac{(\alpha^{\prime})^{2}}{r^{1+b}}h^{-1}\mathrm{d}u,$
-where we take the constant $c>0$ to be small, so that $c+\varepsilon$ can be absorbed by the positive integral on the left hand side, noting $02\geq\gamma>b>$ 0. Thus this lemma is concluded.
-4.2. Cauchy stability in weighted $H^{1}$ space
-The proof of Theorem 4.1 itself leads to the Cauchy stability theorem in $H^{1}_{w}$ space.
-Theorem 4.10.
-[…]
-$\int_{C^{[\underline{u}_{\ast},\underline{u}_{1}]}_{u_{0}}}\frac{\Omega^{2}h}{r^{1+b}}\left(\frac{r^{-2}\hat{w}^{2}}{\xi^{3+\gamma}}+\frac{\alpha^{2}}{\xi^{1+\gamma}}+\xi^{1-\gamma}\frac{(\alpha^{\prime})^{2}}{2}\right)\mathrm{d}\underline{u}<\varepsilon^{2},$
-and assume $|r-r_{\ast}|^{-\frac{\gamma}{2}}\alpha|_{C_{u_{0}}}$ is continuous at $\underline{u}=\underline{u}_{\ast}$.
-There exists an $\varepsilon_{0}>0$ such that if $\varepsilon^{2}\leq\varepsilon^{2}_{0}$, the following conclusions hold.
-Case I: $b>0$. There is a unique solution
-$(r,\,m,\,\alpha)\in C^{1}\times C^{1}\times H^{1}_{w}\quad\text{in}\,\,\mathcal{D}(u_{0},\,u;\,\underline{u}_{1}),\,\,u\in[u_{0},\,\underline{u}_{\ast}).$
-Moreover, for any $(u,\,\underline{u})\in[u_{0},\,\underline{u}_{\ast})\times[\underline{u}_{\ast},\,\underline{u}_{1}]$, there is
-$\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(\frac{(\partial_{u}w)^{2}}{\xi^{3+\gamma}}+\frac{r^{-2}\hat{w}^{2}}{\xi^{1+\gamma}}+\xi^{1-\gamma}\alpha^{2}\right)\mathrm{d}u$
-$+\int_{C_{u}}\frac{\Omega^{2}h}{r^{1+b}}\left(\frac{r^{-2}\hat{w}^{2}}{\xi^{3+\gamma}}+\frac{\alpha^{2}}{\xi^{1+\gamma}}+\xi^{1-\gamma}(\alpha^{\prime})^{2}\right)\mathrm{d}\underline{u}\lesssim\varepsilon^{2},$
-and we have the following pointwise estimates
-$\xi^{-2-\gamma}r^{-2}\hat{w}^{2}+\xi^{-\gamma}\alpha^{2}+\xi^{-3-\gamma}|\partial_{u}w|^{2}+\xi^{-1-\gamma}\mu\lesssim\varepsilon^{2}r^{b}.$
-Case II: $b=0$. There is a unique solution
-$(r,\,m,\,\alpha)\in C^{1}\times C^{0}\times H^{1}_{w},\quad\text{in}\,\,\mathcal{D}(u_{0},\,u;\,\underline{u}_{1}),\,\,u\in[u_{0},\,\underline{u}_{\ast}).$
-The estimates in Case I hold (with $b=0$) except the $L^{\infty}$ estimate for $\partial_{u}w$.
-Appendix A Local existence theorem
-For the characteristic problem where initial data are prescribed on a future light cone with vertex at the center of symmetry, the local existence theorem for the classical solution to the spherically symmetric Einstein scalar field system was proved in [8].
-Regarding to the proof, Christodoulou reduced the whole system in Bondi coordinates to a scalar equation and utilized the contracting mapping principle to obtain the existence theorem.
-However, in this section, we mean to provide an alternative proof based on the energy method.
-To be more precisely, we work within the double null gauge and demonstrate a local existence theory for the SSEYM system in $H^{1}$ (and $H^{1}_{w}$) space, using an iteration method.
-Theorem A.1.
-Let $\underline{u}_{1}\in(u_{0},\,+\infty)$ be a finite constant. Suppose on the initial slice $C_{u_{0}}$, the following energy is bounded
-$\int_{C_{u_{0}}}r^{-1-b}\left(\frac{\hat{w}^{2}}{r^{2}}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}r<I_{1}.$
-Case I: $0<b<\frac{\sqrt{89}-9}{4}$.
-i) There exists $u_{\ast}>u_{0}$ such that the SSEYM system has a unique solution $\left(r,\,m,\,\alpha\right)\in C^{1}\times C^{1}\times H^{1}$ on the domain $\mathcal{D}(u_{0},u_{\ast};\underline{u}_{1})$, which enjoys the subsequent bound,
-$\int_{C_{u}}r^{-1-b}\left(\frac{\hat{w}^{2}}{r^{2}}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}r\lesssim I_{1},$
-$\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(\frac{\hat{w}^{2}}{r^{2}}+\alpha^{2}+h^{2}(\partial_{u}\hat{w})^{2}\right)\mathrm{d}u\lesssim I_{1},$
-$|r^{-1}\hat{w}|^{2}+|\partial_{u}\hat{w}|+|\alpha|^{2}+\mu\lesssim I_{1}r^{b},$
-$|\ln\left(\Omega^{2}h\right)|,\,|\ln\left(-\underline{h}\right)|,\,|\ln h|\lesssim I_{1}.$
-ii) The solution depends continuously on data in the following sense.
-Suppose $\left(r_{n},\,m_{n},\,\hat{w}_{n}\right)$ and $\left(r_{m},\,m_{m},\,\hat{w}_{m}\right)$ are solutions with data $\hat{w}_{0n}$, $\hat{w}_{0m}$ respectively, then
-$\int_{C_{u}}\left(\frac{|\hat{w}_{n}-\hat{w}_{m}|^{2}}{r^{3+b}_{n}}+\frac{|\alpha_{n}-\alpha_{m}|^{2}}{r^{1+b}_{n}}+\frac{|\alpha^{\prime}_{n}-\alpha^{\prime}_{m}|^{2}}{r^{1+b}_{n}}\right)\mathrm{d}r_{n}$
-$+\int_{\underline{C}_{\underline{u}}}\frac{h_{n}^{-1}}{r^{1+b}_{n}}\left(\frac{\left(\hat{w}_{n}-\hat{w}_{m}\right)^{2}}{r_{n}^{2}}+\left(\alpha_{n}-\alpha_{m}\right)^{2}+h_{n}^{2}(\partial_{u}\hat{w}_{n}-\partial_{u}\hat{w}_{m})^{2}\right)\mathrm{d}u$
-[…]
-Case II: $b=0$.
-The statement in Case I holds with $b=0$, except that the pointwise estimate for $|\partial_{u}w|$ is absent and hence in this case the solution $\left(r,\,m,\,\alpha\right)\in C^{1}\times C^{0}\times H^{1}$.
-Remark A.2.
-We can prove the local existence theorem in $H^{1}_{w}$ space (or rather the space of data in Theorem 4.1) with $0\leq b<\frac{\sqrt{89}-9}{4}$ and $b<\gamma\leq 2$ straightforwardly in the same manner as Theorem A.1.
-Thus, we omit the details.
-Alternatively, since $H^{1}$ is dense in $H^{1}_{w}$ (actually, smooth functions with compact support are dense in $H^{1}_{w}$, see Appendix C), we can make use of the property ii) that the solution continuously depends on data, to verify the local existence in $H^{1}_{w}$ space.
-Proof.
-For the sake of simplicity, we only carry out the proof for the case of $b=0$ in details. The proof for the other case $0<b<\frac{\sqrt{89}-9}{4}$ goes through in an analogous fashion.
-It suffices to prove the local existence theorem near the axis, which essentially turns the local existence problem into the one with small data.
-Away from the axis, the local existence follows in a standard way.
-Choose $\underline{u}_{1}>u_{0}$ and $\underline{u}_{1}-u_{0}$ sufficiently small so that the initial energy is sufficiently small.
-We denote the size of data by $\varepsilon^{2}I_{1}$, where $\varepsilon^{2}$ is small enough, namely,
-$\int_{C_{u_{0}}^{[u_{0},\,\underline{u}_{1}]}}r^{-1-b}\left(\frac{\hat{w}^{2}}{r^{2}}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}r\leq\varepsilon^{2}I_{1}.$
-We shall find a local solution for the SSEYM system using the method of Picard iteration. Notice that, the Yang–Mills equation can be alternatively written in terms of $\hat{w}$ (1.3)
-$\partial_{u}\partial_{\underline{u}}\hat{w}+2\frac{\Omega^{2}}{r^{2}}\hat{w}=\Omega^{2}\frac{\hat{w}^{2}}{r^{2}}\left(\hat{w}+3\right).$ (A.1)
-[…]
-$\partial_{u}\partial_{\underline{u}}\hat{w}_{k+1}+{}$ $2\Omega_{k}^{2}\frac{\hat{w}_{k+1}}{r_{k}^{2}}=\Omega_{k}^{2}\frac{\hat{w}_{k}^{2}}{r_{k}^{2}}\left(\hat{w}_{k}+3\right),$ (A.2)
-$\partial_{\underline{u}}m_{k+1}={}$ $\Omega^{2}_{k}h_{k}\left(\frac{\Phi_{k+1}^{2}}{2}-h_{k}\underline{h}_{k}\alpha_{k+1}^{2}\right),$ (A.3)
-$\partial_{u}m_{k+1}={}$ $\underline{h}_{k}\left(\frac{\Phi_{k+1}^{2}}{2}-\left(h_{k}\underline{h}_{k}\right)^{-1}h^{2}_{k}(\partial_{u}\hat{w}_{k+1})^{2}\right),$ (A.4)
-$\partial_{u}\left(\Omega^{2}_{k+1}h_{k+1}\right)={}$ $\Omega^{2}_{k+1}\underline{h}^{-1}_{k+1}\underline{h}_{k}\left(-\frac{2m_{k+1}}{r^{2}_{k}}+\frac{\Phi_{k+1}^{2}}{r_{k}}\right),$ (A.5)
-$\partial_{\underline{u}}\left(\underline{h}_{k+1}\right)={}$ $\Omega^{2}_{k}h_{k}h_{k+1}^{-1}\left(-\frac{2m_{k+1}}{r^{2}_{k}}+\frac{\Phi_{k+1}^{2}}{r_{k}}\right),$ (A.6)
-with initial data prescribed in Theorem A.1 and the boundary conditions
+It suffices to prove the local existence theorem near the axis, which essentially reduces the problem to a small-data regime. Away from the axis, local existence follows in a standard way.
+Choose $\underline{u}_{1}>u_{0}$ such that $\underline{u}_{1}-u_{0}$ is sufficiently small, ensuring that the initial energy remains sufficiently small.
+We denote the size of the initial data by $\varepsilon^{2}I_{1}$, where $\varepsilon^{2}$ is small enough, namely,
+$\int_{C_{u_{0}}^{[u_{0},\,\underline{u}_{1}]}}r^{-1}\left(\frac{\hat{w}^{2}}{r^{2}}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}r\leq\varepsilon^{2}I_{1}.$
+We shall find a local solution for the SSEYM system using the method of Picard iteration.
+Linearized system.
+We let $\left(r_{0},\,\hat{w}_{0},\,m_{0}\right)=\left(r|_{C_{u_{0}}}+\frac{u_{0}-u}{2},\,0,\,0\right)$ and then define $\left(r_{k+1},\,m_{k+1},\,\hat{w}_{k+1}\right)$ iteratively for $k\geq 0$ as the unique solution to the linearized system:
+$\partial_{u}\partial_{\underline{u}}\hat{w}_{k+1}+{}$ $2\Omega_{k}^{2}\frac{\hat{w}_{k+1}}{r_{k}^{2}}=\Omega_{k}^{2}\frac{\hat{w}_{k}^{2}}{r_{k}^{2}}\left(\hat{w}_{k}+3\right),$ (4.2)
+$\partial_{\underline{u}}m_{k+1}={}$ $\Omega^{2}_{k}h_{k}\left(\frac{\Phi_{k+1}^{2}}{2}-h_{k}\underline{h}_{k}\alpha_{k+1}^{2}\right),$ (4.3)
+$\partial_{u}m_{k+1}={}$ $\underline{h}_{k}\left(\frac{\Phi_{k+1}^{2}}{2}-\left(h_{k}\underline{h}_{k}\right)^{-1}h^{2}_{k}(\partial_{u}\hat{w}_{k+1})^{2}\right),$ (4.4)
+$\partial_{u}\left(\Omega^{2}_{k+1}h_{k+1}\right)={}$ $\Omega^{2}_{k+1}\underline{h}^{-1}_{k+1}\underline{h}_{k}\left(-\frac{2m_{k+1}}{r^{2}_{k}}+\frac{\Phi_{k+1}^{2}}{r_{k}}\right),$ (4.5)
+$\partial_{\underline{u}}\left(\underline{h}_{k+1}\right)={}$ $\Omega^{2}_{k}h_{k}h_{k+1}^{-1}\left(-\frac{2m_{k+1}}{r^{2}_{k}}+\frac{\Phi_{k+1}^{2}}{r_{k}}\right),$ (4.6)
+with initial data prescribed in Theorem 4.1 and subject to the boundary conditions
 $m_{k+1}|_{\Gamma}=0,\quad r_{k+1}|_{\Gamma}=0,\quad\hat{w}_{k+1}|_{\Gamma}=0.$
@@
 We insist the choice of gauge in Section 2.2 so that
-$\Omega_{k+1}^{2}h_{k+1}|_{C_{u_{0}}}=\frac{1}{2}.$ (A.7)
+$\Omega_{k+1}^{2}h_{k+1}|_{C_{u_{0}}}=\frac{1}{2}.$ (4.7)
 Besides, our gauge choice
@@
 We remark that, for the first generation, we have $\underline{h}_{0}=-\frac{1}{2}$, and $h_{0}=2$, $\Omega^{2}_{0}=\frac{1}{4}$. In other words, the first generation is the vacuum spacetime.
-In the linearized system (A.2)-(A.6), we first use (A.2) to solve $\hat{w}_{k+1}$.
-Next, substituting the $\hat{w}_{k+1}$ into (A.3) and (A.4), we can further solve $m_{k+1}$.
-After that, substituting both of $\hat{w}_{k+1}$ and $m_{k+1}$ into (A.5) and (A.6), we are able to solve $\Omega^{2}_{k+1}h_{k+1}$ and $\underline{h}_{k+1}$, and hence obtain $r_{k+1}$.
-Auxiliary equations. Along with the main equations, we can derive the following two equations from (A.3) and (A.6), (A.4) and (A.5).
-$\partial_{\underline{u}}h_{k+1}={}$ $2\Omega^{2}_{k}h_{k}(1-\mu_{k})\underline{h}_{k+1}^{-1}\frac{(\alpha_{k+1})^{2}}{r_{k}},$ (A.8)
-$\partial_{u}(\Omega^{-2}_{k+1}\underline{h}_{k+1})={}$ $-2\Omega^{-2}_{k+1}h^{-1}_{k+1}h_{k}\frac{(\partial_{u}\hat{w}_{k+1})^{2}}{r_{k}}.$ (A.9)
+In the linearized system (4.2)-(4.6), we first solve (4.2) to determine $\hat{w}_{k+1}$.
+Substituting the $\hat{w}_{k+1}$ into (4.3) and (4.4), we then solve $m_{k+1}$.
+Finally, substituting both of $\hat{w}_{k+1}$ and $m_{k+1}$ into (4.5) and (4.6), we are able to solve $\Omega^{2}_{k+1}h_{k+1}$ and $\underline{h}_{k+1}$, thereby obtaining $r_{k+1}$.
+Auxiliary equations. Along with the main equations, we can derive the following two equations from (4.3) and (4.6), (4.4) and (4.5).
+$\partial_{\underline{u}}h_{k+1}={}$ $2\Omega^{2}_{k}h_{k}(1-\mu_{k})\underline{h}_{k+1}^{-1}\frac{(\alpha_{k+1})^{2}}{r_{k}},$ (4.8)
+$\partial_{u}(\Omega^{-2}_{k+1}\underline{h}_{k+1})={}$ $-2\Omega^{-2}_{k+1}h^{-1}_{k+1}h_{k}\frac{(\partial_{u}\hat{w}_{k+1})^{2}}{r_{k}}.$ (4.9)
 With the above setup, we now continue the proof with $H^{1}$ estimates.
-Uniform bound. We claim that if $\varepsilon$ (or rather $\underline{u}_{1}-u_{0}$) is small enough, there is a constant $M<+\infty$ such that we have the uniform bounds, for all $i\geq 0$,
+Uniform bound. We assert that if $\varepsilon$ (or equivalently, $\underline{u}_{1}-u_{0}$) is small enough, there is a constant $M<+\infty$ such that the uniform bounds,
 $A_{i+1}(u,\underline{u}):=$ $\int_{C_{u}}\Omega^{2}_{i}h_{i}\left(\frac{(\alpha^{\prime}_{i+1})^{2}}{r_{i}}+\frac{\alpha_{i+1}^{2}}{r_{i}}+\frac{\hat{w}_{i+1}^{2}}{r^{3}_{i}}\right)\mathrm{d}u$
 $+\int_{\underline{C}_{\underline{u}}}h_{i}^{-1}\left(\frac{\alpha_{i+1}^{2}}{r_{i}}+h_{i}^{2}\frac{(\partial_{u}\hat{w}_{i+1})^{2}}{r_{i}}+\frac{\hat{w}_{i+1}^{2}}{r^{3}_{i}}\right)\mathrm{d}u$
-[…]
-When $i=0$, an energy argument (in line with the proof leading to (3.1.4) in the subsection 3.1.4) for the linearized equation
+$\leq{}$ $\varepsilon^{2}M,\quad u\in[u_{0},\,u_{\ast}],\,\,\underline{u}\in[u,\,\underline{u}_{1}],$ (4.10)
+hold for all $i\geq 0$,
+When $i=0$, an energy argument (in line with the proof leading to (3.5) in the subsection 3.5) for the linearized equation
 $\partial_{u}\partial_{\underline{u}}\hat{w}_{1}+2\Omega^{2}_{0}\frac{\hat{w}_{1}}{r_{0}^{2}}=0,$
@@
 $+\iint_{\mathcal{D}}\frac{\Omega^{2}_{0}}{r^{2}_{0}}\left(18r^{-2}_{0}\hat{w}_{1}^{2}+2h_{0}^{2}(\partial_{u}\hat{w}_{1})^{2}+\alpha^{2}_{1}+(\alpha^{\prime}_{1})^{2}\right)\mathrm{d}u\mathrm{d}\underline{u}$
-[…]
-Thus, (A) holds for $i=0$, which further implies
+$={}$ $\int_{C_{u_{0}}}\frac{\Omega^{2}_{0}h_{0}}{r_{0}}\left((\alpha^{\prime}_{1})^{2}+9\alpha_{1}^{2}+4r_{0}^{-2}\hat{w}_{1}^{2}\right)\mathrm{d}\underline{u}.$
+Thus, (4) holds for $i=0$, which further implies
 $\mu_{1}+\alpha^{2}_{1}+\frac{\hat{w}^{2}_{1}}{r_{0}^{2}}\lesssim\varepsilon^{2}I_{1},$
 $|\ln(\Omega_{1}h_{1})|,\,|\ln(-\underline{h}_{1})|,\,|\ln h_{1}|\lesssim I_{1}.$
-To proceed, let us suppose the estimate (A) holds for $i\leq k-1$ with the constant $M$ to be determined later. Then we can derive the pointwise estimates, as in the subsections 3.1.5 and 3.1.6
+To proceed, let us suppose the estimate (4) holds for $i\leq k-1$ with the constant $M$ to be determined later. Then we can derive the pointwise estimates, as in the subsections 3.6 and 3.7
 $\mu_{i+1}+\alpha^{2}_{i+1}+\frac{\hat{w}^{2}_{i+1}}{r_{i}^{2}}\lesssim\varepsilon^{2}M,\quad i\leq k-1,$
 $|\ln(\Omega_{i+1}h_{i+1})|,\,|\ln(-\underline{h}_{i+1})|,\,|\ln h_{i+1}|\lesssim I_{1},\quad i\leq k-1.$
-Moreover, using the equation of (A.5) with $k$ replaced by $k-1$ or $k-2$, and noting the gauge choice (A.7), we have
+Moreover, using the equation of (4.5) with $k$ replaced by k1k-1 or k2k-2, and noting the gauge choice (4.7), we have
 $\Big{|}\ln\frac{\Omega_{k-1}^{2}h_{k-1}}{\Omega_{k}^{2}h_{k}}\Big{|}\lesssim\varepsilon^{2}M,\quad\Big{|}\ln\frac{\Omega_{k-2}^{2}h_{k-2}}{\Omega_{k-1}^{2}h_{k-1}}\Big{|}\lesssim\varepsilon^{2}M.$
@@
 $\frac{1}{2}<\big{|}\frac{r_{k}}{r_{k-1}}\big{|},\,\,\big{|}\frac{r_{k-1}}{r_{k-2}}\big{|}<\frac{3}{2}.$
-In addition, taking the equation of (A.6) (with $k$ replaced by $k-1$) and the gauge condition of $\underline{h}_{k}|_{\Gamma}=-\Omega^{2}_{k}h_{k}|_{\Gamma}$ into account, we find that
+In addition, taking the equation of (4.6) (with $k$ replaced by k1k-1) and the gauge condition of $\underline{h}_{k}|_{\Gamma}=-\Omega^{2}_{k}h_{k}|_{\Gamma}$ into account, we find that
 $\Big{|}\ln\frac{\underline{h}_{k-1}}{\underline{h}_{k}}\Big{|}\lesssim\varepsilon^{2}M,$
@@
 follows, since $\frac{h_{k-1}}{h_{k}}=\frac{h_{k-1}\underline{h}_{k-1}}{h_{k}\underline{h}_{k}}\cdot\frac{\underline{h}_{k}}{\underline{h}_{k-1}}$.
-We shall then prove (A) holds for $i=k$. Thanks to the preceding estimates and an energy argument in line with the subsection 3.1.4, we are able to derive
+We shall then prove (4) holds for $i=k$. Thanks to the preceding estimates and an energy argument analogous to that in Subsection 3.5, we are able to derive
 $\int_{C_{u}}\Omega^{2}_{k}h_{k}\left(\frac{(\alpha^{\prime}_{k+1})^{2}}{r_{k}}+9\frac{\alpha_{k+1}^{2}}{r_{k}}+4\frac{\hat{w}_{k+1}^{2}}{r_{k}}\right)\mathrm{d}\underline{u}$
@@
 $+\iint_{\mathcal{D}}\frac{1}{2}\Omega^{2}_{k}r^{-2}_{k}\left(18r^{-2}_{k}\hat{w}_{k+1}^{2}+2h_{k}^{2}(\partial_{u}\hat{w}_{k+1})^{2}+\alpha^{2}_{k+1}+(\alpha^{\prime}_{k+1})^{2}\right)\mathrm{d}u\mathrm{d}\underline{u}$
-[…]
-where the universal constant $C$ is independent of $k$. If we choose $M>CI_{1}$, then (A) with $i=k$ is verified.
+$\leq{}$ $\varepsilon^{2}I_{1}\cdot C,$ (4.11)
+where the universal constant $C$ is independent of $k$. If we choose $M>CI_{1}$, then (4) with $i=k$ is verified.
 Convergence.
-In the compact domain concerned, since $\{r_{k}\}$ is equi-continuous, there is a subsequence, which we still denote by $\{r_{k}\}$, such that $r_{k}\rightarrow r$ on the compact neighbourhood.
-In what follows, we always use this converging subsequence to replace the original one.
-Based on (A.2), we derive the following linear equation,
+In the compact domain concerned, the sequence $\{r_{k}\}$ is equicontinuous.
+Therefore, there is a subsequence (still denoted by $\{r_{k}\}$) that converges uniformly to a function $r$, $r_{k}\rightarrow r$, on the compact neighbourhood.
+In the subsequent analysis, we will always work with this convergent subsequence in place of the original sequence.
+Based on (4.2), we derive the following linear equation,
 $\partial_{u}\partial_{\underline{u}}(\hat{w}_{k+1}-\hat{w}_{k})+2\frac{\Omega_{k-1}^{2}}{r_{k-1}^{2}}(\hat{w}_{k+1}-\hat{w}_{k})$
-[…]
+$={}$ $f\left(\frac{\hat{w}_{k+1}}{r_{k}},\,\frac{\hat{w}_{k}}{r_{k}},\,\frac{\hat{w}_{k-1}}{r_{k-1}}\right)\cdot\frac{(U_{k}-U_{k-1})}{r_{k}},$ (4.12)
 where $U_{k}-U_{k-1}$ involves
@@
 We will prove the convergence by induction. Suppose
-$\|U_{k}-U_{k-1}\|_{L^{\infty}}\leq M2^{-k+1},\quad\|\mu_{k}-\mu_{k-1}\|_{L^{\infty}}\leq M2^{-k+1},$ (A.13)
-$\Big{|}\int_{C_{u}}\frac{m_{k}}{r_{k-1}^{2}}\mathrm{d}r_{k-1}|_{C_{u}}-\int_{C_{u}}\frac{m_{k-1}}{r_{k-2}^{2}}\mathrm{d}r_{k-2}|_{C_{u}}\Big{|}\leq M2^{-k+1},$ (A.14)
-$\Big{|}\int_{\underline{C}_{\underline{u}}}\frac{m_{k}}{r_{k-1}^{2}}\mathrm{d}r_{k-1}|_{\underline{C}_{\underline{u}}}-\int_{\underline{C}_{\underline{u}}}\frac{m_{k-1}}{r_{k-2}^{2}}\mathrm{d}r_{k-2}|_{\underline{C}_{\underline{u}}}\Big{|}\leq M2^{-k+1},$ (A.15)
-for a large constant $M$ to be determined. When $k=1$, (A.13)-(A.15) hold by the uniform boundedness in the previous step. We will prove the same inequalities hold with $k$ replaced by $k+1$.
-As is obvious from the expression, the linear equation (A) shares the same structure as (A.2). Thus, we can establish an analogous energy identity, noting that data vanishes now
+$\|U_{k}-U_{k-1}\|_{L^{\infty}}\leq M2^{-k+1},\quad\|\mu_{k}-\mu_{k-1}\|_{L^{\infty}}\leq M2^{-k+1},$ (4.13)
+$\Big{|}\int_{C_{u}}\frac{m_{k}}{r_{k-1}^{2}}\mathrm{d}r_{k-1}|_{C_{u}}-\int_{C_{u}}\frac{m_{k-1}}{r_{k-2}^{2}}\mathrm{d}r_{k-2}|_{C_{u}}\Big{|}\leq M2^{-k+1},$ (4.14)
+$\Big{|}\int_{\underline{C}_{\underline{u}}}\frac{m_{k}}{r_{k-1}^{2}}\mathrm{d}r_{k-1}|_{\underline{C}_{\underline{u}}}-\int_{\underline{C}_{\underline{u}}}\frac{m_{k-1}}{r_{k-2}^{2}}\mathrm{d}r_{k-2}|_{\underline{C}_{\underline{u}}}\Big{|}\leq M2^{-k+1},$ (4.15)
+for a large constant $M$ to be determined. When $k=1$, (4.13)-(4.15) hold by the uniform boundedness in the previous step. We will prove the same inequalities hold with $k$ replaced by $k+1$.
+As is evident from its expression, the linear equation (4) shares the same structure as (4.2).
+Thus, we can establish an analogous energy identity, taking into account that the initial data now vanish,
 $\int_{C_{u}}h_{k-1}\Omega^{2}_{k-1}\frac{|X_{k+1}-X_{k}|^{2}}{r_{k-1}}\mathrm{d}\underline{u}+\int_{\underline{C}_{\underline{u}}}h^{-1}_{k-1}\frac{|Y_{k+1}-Y_{k}|^{2}}{r_{k-1}}\mathrm{d}u$
 $+\iint_{\mathcal{D}}\Omega^{2}_{k-1}r^{-2}_{k-1}|V_{k+1}-V_{k}|^{2}\mathrm{d}u\mathrm{d}\underline{u}$
-[…]
-where, for simplicity, we denote
+$\lesssim{}$ $\iint_{\mathcal{D}}\left(\big{|}\frac{\hat{w}_{k+1}}{r_{k}}\big{|}+\big{|}\frac{\hat{w}_{k}}{r_{k-1}}\big{|}+\big{|}\frac{\hat{w}_{k-1}}{r_{k-2}}\big{|}\right)\Omega_{k-1}^{2}\frac{|U_{k}-U_{k-1}||V_{k+1}-V_{k}|}{r_{k-1}r_{k-2}}\mathrm{d}u\mathrm{d}\underline{u}.$
+Here, for simplicity, we denote
 $X_{k+1}-X_{k}={}$ $\left(r_{k-1}\partial_{r_{k-1}}^{2}(\hat{w}_{k+1}-\hat{w}_{k}),\,\partial_{r_{k-1}}(\hat{w}_{k+1}-\hat{w}_{k}),\,r_{k-1}^{-1}(\hat{w}_{k+1}-\hat{w}_{k})\right),$
@@
 Then we estimate the right hand side of the inequality as follows,
-[…]
+$\lesssim{}$ $\|U_{k}-U_{k-1}\|_{L^{\infty}}\left(\iint_{\mathcal{D}}\frac{\Omega_{k-1}^{2}}{r^{2}_{k-2}}\left(\big{|}\frac{\hat{w}_{k+1}}{r_{k}}\big{|}^{2}+\big{|}\frac{\hat{w}_{k}}{r_{k-1}}\big{|}^{2}+\big{|}\frac{\hat{w}_{k-1}}{r_{k-2}}\big{|}^{2}\right)\mathrm{d}u\mathrm{d}\underline{u}\right)^{\frac{1}{2}}$
 $\qquad\cdot\left(\iint_{\mathcal{D}}\Omega_{k-1}^{2}\frac{|V_{k+1}-V_{k}|^{2}}{r^{2}_{k-1}}\mathrm{d}u\mathrm{d}\underline{u}\right)^{\frac{1}{2}}$
-[…]
+$\lesssim{}$ $\varepsilon\|U_{k}-U_{k-1}\|_{L^{\infty}}\left(\iint_{\mathcal{D}}\Omega_{k-1}^{2}\frac{|V_{k+1}-V_{k}|^{2}}{r^{2}_{k-1}}\mathrm{d}u\mathrm{d}\underline{u}\right)^{\frac{1}{2}}.$
 As a result, an application of the Cauchy-Schwarz leads to
@@
 $+\iint_{\mathcal{D}}\frac{1}{2}\Omega^{2}_{k-1}r^{-2}_{k-1}|V_{k+1}-V_{k}|^{2}\mathrm{d}u\mathrm{d}\underline{u}$
-[…]
-For notational convenience, in what follows, we will denote
+$\lesssim{}$ $\varepsilon\|U_{k}-U_{k-1}\|^{2}_{L^{\infty}}.$ (4.16)
+For notational simplification, in what follows, we will denote
 $\|\hat{w}_{k+1}-\hat{w}_{k}\|^{2}_{H^{1}(C_{u})}$
-[…]
+$:={}$ $\int_{C_{u}}\left(|\partial_{r_{k-1}}\left(\hat{w}_{k+1}-\hat{w}_{k}\right)|^{2}+\frac{|\hat{w}_{k+1}-\hat{w}_{k}|^{2}}{r_{k-1}^{4}}\right)\Omega_{k-1}^{2}h_{k-1}\mathrm{d}\underline{u},$
 $\|\hat{w}_{k+1}-\hat{w}_{k}\|^{2}_{H^{2}(C_{u})}$
-[…]
-Convergence of $\{\hat{w}_{k}\}$ uniformly in the energy norm. By (A.13) and the fact $\varepsilon$ is small, the estimate (A) tells that
+$:={}$ $\int_{C_{u}}|r_{k-1}\partial^{2}_{r_{k-1}}\left(\hat{w}_{k+1}-\hat{w}_{k}\right)|^{2}\Omega_{k-1}^{2}h_{k-1}\mathrm{d}\underline{u}+\|\hat{w}_{k+1}-\hat{w}_{k}\|^{2}_{H^{1}(C_{u})}.$
+Convergence of $\{\hat{w}_{k}\}$ uniformly in the energy norm. By (4.13) and the fact $\varepsilon$ is small, the estimate (4) tells that
 $\|\hat{w}_{k+1}-\hat{w}_{k}\|^{2}_{H^{2}}(C_{u})+\int_{\underline{C}_{\underline{u}}}h^{-1}_{k-1}\frac{|Y_{k+1}-Y_{k}|^{2}}{r_{k-1}}\mathrm{d}u$
 $\qquad+\iint_{\mathcal{D}}\Omega^{2}_{k-1}r^{-2}_{k-1}|V_{k+1}-V_{k}|^{2}\mathrm{d}u\mathrm{d}\underline{u}$
-[…]
-Moreover, by virtue of the fact $\frac{|\hat{w}_{k+1}-\hat{w}_{k}|^{2}}{r_{k-1}^{2}}\lesssim\|\hat{w}_{k+1}-\hat{w}_{k}\|^{2}_{H^{1}(C_{u})}$, the previous $L^{\infty}$ estimates and (A.13), it follows that
+$\leq{}$ $\varepsilon CM^{2}2^{-2k+2}.$ (4.17)
+Moreover, by virtue of the fact $\frac{|\hat{w}_{k+1}-\hat{w}_{k}|^{2}}{r_{k-1}^{2}}\lesssim\|\hat{w}_{k+1}-\hat{w}_{k}\|^{2}_{H^{1}(C_{u})}$, the previous $L^{\infty}$ estimates and (4.13), it follows that
 $\Big{|}\frac{\hat{w}_{k+1}}{r_{k}}-\frac{\hat{w}_{k}}{r_{k-1}}\Big{|}\lesssim{}$ $\Big{|}\frac{\hat{w}_{k+1}}{r_{k}}\Big{|}\Big{|}1-\frac{r_{k}}{r_{k-1}}\Big{|}+\|\hat{w}_{k+1}-\hat{w}_{k}\|_{H^{1}}$
-[…]
+$\leq{}$ $\sqrt{\varepsilon}CM2^{-k+1}<M2^{-k},$ (4.18)
 if we choose $\varepsilon$ small enough such that $2\sqrt{\varepsilon}C<\frac{1}{2}$ and $M>1$.
 Convergence of $\mu_{k}$ (uniformly in $C^{0}$ norm). Noting that
-$\frac{m_{k+1}}{r_{k}}-\frac{m_{k}}{r_{k-1}}={}$ $\frac{m_{k+1}-m_{k}}{r_{k}}+\frac{m_{k}}{r_{k}}\left(1-\frac{r_{k}}{r_{k-1}}\right),$ (A.19)
-and using (A.3), we obtain
+$\frac{m_{k+1}}{r_{k}}-\frac{m_{k}}{r_{k-1}}={}$ $\frac{m_{k+1}-m_{k}}{r_{k}}+\frac{m_{k}}{r_{k}}\left(1-\frac{r_{k}}{r_{k-1}}\right),$ (4.19)
+and using (4.3), we obtain
 $\frac{m_{k+1}-m_{k}}{r_{k}}\lesssim{}$ $\|\hat{w}_{k+1}\|^{2}_{H^{1}}\left(|U_{k}-U_{k-1}|+|\mu_{k}-\mu_{k-1}|\right)$
 $+\left(\|\hat{w}_{k+1}\|_{H^{1}}+\|\hat{w}_{k}\|_{H^{1}}\right)\|\hat{w}_{k+1}-\hat{w}_{k}\|_{H^{1}}.$
-Then it follows from the inductive assumptions (A.13) and (A) that
-$|\mu_{k+1}-\mu_{k}|\leq\varepsilon CM2^{-k+1}.$ (A.20)
+Then it follows from the inductive assumptions (4.13) and (4) that
+$|\mu_{k+1}-\mu_{k}|\leq\varepsilon CM2^{-k+1}.$ (4.20)
 We conclude through choosing $\varepsilon$ and $M$ properly that
 $|\mu_{k+1}-\mu_{k}|\leq M2^{-k}.$
-Furthermore, in view of the formulae (3.1.6) and (3.1.6), it follows from (A) and (A.20) that (A.14)–(A.15) hold with $k$ replaced by $k+1$.
-Convergence of $\Omega^{2}_{k}h_{k}$, $\underline{h}_{k}$, $h_{k}$, $m_{k}$ (uniformly in $C^{0}$ norm). Similarly, using the equations (A.5) and (A.6), we prove that
+Furthermore, in view of the formulae (3.7) and (3.7), it follows from (4) and (4.20) that (4.14)–(4.15) hold with $k$ replaced by $k+1$.
+Convergence of $\Omega^{2}_{k}h_{k}$, $\underline{h}_{k}$, $h_{k}$, $m_{k}$ (uniformly in $C^{0}$ norm). Similarly, using the equations (4.5) and (4.6), we prove that
 $\Big{|}\ln\frac{\Omega^{2}_{k+1}h_{k+1}}{\Omega^{2}_{k}h_{k}}\Big{|}\leq\varepsilon CM2^{-k+1}\leq M2^{-k},$
 $\Big{|}\ln\frac{\underline{h}_{k+1}}{\underline{h}_{k}}\Big{|}\leq\varepsilon CM2^{-k+1}\leq M2^{-k},$
-which together with (A.20) implies $|h_{k+1}^{-1}-h^{-1}_{k}|\leq\varepsilon M2^{-k}$. The inductive proof is thus complete.
-Finally, substituting all the previous results into the equation of mass (A.3), we obtain
+which together with (4.20) implies $|h_{k+1}^{-1}-h^{-1}_{k}|\leq\varepsilon M2^{-k}$. The inductive proof is thus complete.
+Finally, substituting all the previous results into the equation of mass (4.3), we obtain
 $m_{k}-m_{k-1}=O(2^{-k}).$
@@
 $\mu=\frac{m}{r}.$
-Along with the formula of $\mu$, $m$ and $r$, the relation $h_{k}=\frac{\mu_{k}-1}{\underline{h}_{k}}$ gives rise to an identity
+Besides, the identity $h_{k}=\frac{\mu_{k}-1}{\underline{h}_{k}}$ gives rise to
 $h=\frac{\mu-1}{\underline{h}}.$
@@
 $\mu\in C^{0},\,\,\ln\Omega^{2}h\in C^{0},\,\,\ln(-\underline{h})\in C^{0},\,\,\ln h\in C^{0},\,\,\text{and}\,\,r\in C^{1}.$
-In addition, (A) suggests
+In addition, (4) suggests
 $\int_{C_{u}}\Omega^{2}hr^{-1}\left(\frac{\left(\hat{w}_{k+1}-\hat{w}_{k}\right)^{2}}{r^{2}}+\left(\partial_{r}\left(\hat{w}_{k+1}-\hat{w}_{k}\right)\right)^{2}+\left(r\partial^{2}_{r}\left(\hat{w}_{k+1}-\hat{w}_{k}\right)\right)^{2}\right)\mathrm{d}\underline{u}$
 $+\int_{\underline{C}_{\underline{u}}}h^{-1}r^{-1}\left(\frac{\left(\hat{w}_{k+1}-\hat{w}_{k}\right)^{2}}{r^{2}}+\left(\partial_{r}\left(\hat{w}_{k+1}-\hat{w}_{k}\right)\right)^{2}+h^{2}\left(\partial_{u}\left(\hat{w}_{k+1}-\hat{w}_{k}\right)\right)^{2}\right)\mathrm{d}u$
-[…]
+$={}$ $O(2^{-k}).$
 Hence, there exists $\hat{w}\in C^{0}$ such that
@@
 $w:=\hat{w}+1,$
-we conclude $(r,\,m,\,w)$ satisfies the SSEYM equations.
-However, since $\partial_{u}\partial_{r}w$ is not necessarily bounded on the axis, the Yang–Mills equation (and also the whole SSEYM system) holds in the sense of distribution.
-Continuous dependence on data. Suppose $\hat{w}_{n}$ and $\hat{w}_{m}$ are solutions with data $\hat{w}_{0n}$, $\hat{w}_{0m}$ respectively. The following equation holds
+we conclude that (r,m,w)(r,\,m,\,w) satisfies the SSEYM equations.
+However, since $\partial_{u}\partial_{r}w$ is not necessarily bounded on the axis, the Yang–Mills equation, and indeed the entire SSEYM system, holds in the sense of distribution.
+Continuous dependence on the data. Suppose $\hat{w}_{n}$ and $\hat{w}_{m}$ are solutions with data $\hat{w}_{0n}$, $\hat{w}_{0m}$ respectively. The following equation holds
 $\partial_{u}\partial_{\underline{u}}(\hat{w}_{n}-\hat{w}_{m})+2\frac{\Omega_{n}^{2}}{r_{n}^{2}}(\hat{w}_{n}-\hat{w}_{m})=f\left(\frac{\hat{w}_{n}}{r_{n}},\,\frac{\hat{w}_{m}}{r_{m}}\right)\cdot\frac{(U_{n}-U_{m})}{r_{n}},$
@@
 In the end, the uniqueness follows in a similar way.
-Appendix B An extension principle away from the axis
-The extension principle away from the axis had been confirmed for systems, such as Einstein-Vlasov and Einstein charged scalar field in spherical symmetry [14, 15].
-A prior to an extension theorem, a local existence theorem away from the axis is needed and it was addressed in [14] as well.
+Appendix A An extension principle away from the axis
+The extension principle away from the axis had been established for several systems, including the spherically symmetric Einstein-Vlasov and Einstein charged scalar field models [14, 15].
+As a prerequisite for such an extension theorem, a local existence theorem away from the axis is needed, which was also provided in [14] as well.
 The proof in [14] was based on the fixed point method and the double null gauge.
-This scheme indeed works well when the Einstein equations are coupled with other matter fields.
-In this section, we follow the approaches in [14, 15] to develop the local theory associated to the SSEYM system.
+This approach remains valid for the Einstein equations coupled with various matter fields.
+In this section, we adapt the methods of [14, 15] to develop a local theory to the SSEYM system away from the axis.
 We rewrite the SSEYM system (2.4a)–(2.4e) and (2.5) in terms of $(r,\Omega,w)$ as follows,
-$r\partial_{u}\partial_{\underline{u}}r={}$ $-\Omega^{2}-\partial_{u}r\partial_{\underline{u}}r+\Omega^{2}r^{-2}Q^{2},$ (B.1a)
-$r^{2}\partial_{u}\partial_{\underline{u}}\ln\Omega={}$ $\Omega^{2}+\partial_{u}r\partial_{\underline{u}}r-2\Omega^{2}r^{-2}Q^{2},$ (B.1b)
-$\partial_{u}\left(\Omega^{-2}\partial_{u}r\right)={}$ $-2\Omega^{-2}r^{-2}\left(\partial_{u}w\right)^{2},$ (B.1c)
-$\partial_{\underline{u}}\left(\Omega^{-2}\partial_{\underline{u}}r\right)={}$ $-2\Omega^{-2}r^{-2}\left(\partial_{\underline{u}}w\right)^{2},$ (B.1d)
-$\partial_{u}\partial_{\underline{u}}w+{}$ $\Omega^{2}r^{-2}wQ=0,$ (B.1e)
-where $Q=w^{2}-1$. In view of the structures of the system (B.1a)-(B.1e), a local existence theorem away from the axis follows in a standard way [14].
-Theorem B.1.
+$r\partial_{u}\partial_{\underline{u}}r={}$ $-\Omega^{2}-\partial_{u}r\partial_{\underline{u}}r+\Omega^{2}r^{-2}Q^{2},$ (A.1a)
+$r^{2}\partial_{u}\partial_{\underline{u}}\ln\Omega={}$ $\Omega^{2}+\partial_{u}r\partial_{\underline{u}}r-2\Omega^{2}r^{-2}Q^{2},$ (A.1b)
+$\partial_{u}\left(\Omega^{-2}\partial_{u}r\right)={}$ $-2\Omega^{-2}r^{-2}\left(\partial_{u}w\right)^{2},$ (A.1c)
+$\partial_{\underline{u}}\left(\Omega^{-2}\partial_{\underline{u}}r\right)={}$ $-2\Omega^{-2}r^{-2}\left(\partial_{\underline{u}}w\right)^{2},$ (A.1d)
+$\partial_{u}\partial_{\underline{u}}w+{}$ $\Omega^{2}r^{-2}wQ=0,$ (A.1e)
+where $Q=w^{2}-1$. In view of the structures of the system (A.1a)-(A.1e), a local existence theorem away from the axis follows in a standard way [14].
+Theorem A.1.
 Let $k\geq 0$ and $\Sigma:=[0,d]\times\{0\}\cup\{0\}\times[0,d]$.
 On $\Sigma$, let $r,\Omega$ be positive functions such that $r\in C^{k+2}(\Sigma)$ and $\Omega\in C^{k+1}(\Sigma)$, and let $w\in C^{k+1}(\Sigma)$.
-Suppose (B.1c)–(B.1d) hold initially on $[0,d]\times\{0\}$ and $\{0\}\times[0,d]$, respectively.
+Suppose (A.1c)–(A.1d) hold initially on $0,d]\times\{0\}$ and $\{0\}\times[0,d]$, respectively.
 Define
@@
 $N:={}$ $\sup\{N_{u},\,N_{\underline{u}}\}.$
-Then there exists a $\delta>0$ depending only on $N$ such that there are functions $(r,\Omega,w)\in C^{k+2}\times C^{k+1}\times C^{k+1}$ satisfying equations (B.1a)–(B.1e) in $[0,\delta^{\ast}]\times[0,\delta^{\ast}]$, where $\delta^{\ast}=\min\{d,\,\delta\}$, with the above prescribed data on $\Sigma$.
-With respect to the above local theorem, we can prove a generalized extension principle away from the axis (i.e. “strongly tame” in the language of [15]) following [15].
-Theorem B.2.
+Then there exists a $\delta>0$ depending only on NN such that there are functions $(r,\Omega,w)\in C^{k+2}\times C^{k+1}\times C^{k+1}$ satisfying equations (A.1a)–(A.1e) in $0,\delta^{\ast}]\times[0,\delta^{\ast}]$, where $\delta^{\ast}=\min\{d,\,\delta\}$, with the above prescribed data on $\Sigma$.
+Building upon the local existence theorem established above, we can prove a generalized extension principle away from the axis (referred to as “strongly tame” in the terminology of [15]) by adapting the methods of [15].
+Theorem A.2.
 Let $\left(\mathcal{Q}^{+},g_{\mu\nu},w\right)$ be the maximal future development of smooth spherical symmetric initial data for the SSEYM system.
@@
 then $p\in\mathcal{Q}^{+}$.
-The proof relies on the dominant energy condition and the special structure of Einstein Yang–Mills system. This structure manifests itself in that to bound all the variables, it suffices to bound
+The proof relies on the dominant energy condition and the special structure of the Einstein Yang–Mills system.
+This structure is reflected in the fact that bounding all variables reduces to estimating the integrals
 $\iint_{\mathcal{D}}r^{2}T_{u\underline{u}}\mathrm{d}u\mathrm{d}\underline{u}\quad\text{and}\quad\iint_{\mathcal{D}}\Omega^{2}T_{AB}g^{AB}\mathrm{d}u\mathrm{d}\underline{u},$
-which in the SSEYM case takes the form of
+which in the SSEYM case takes the form
 $\iint_{\mathcal{D}}\Omega^{2}\frac{Q^{2}}{r^{2}}\mathrm{d}u\mathrm{d}\underline{u}.$
-Nevertheless, this quantity can be bounded straightforwardly through integrating (B.1a), please refer to [15] for more details.
-Theorem B.2 shows that the “first singularity” in the maximal future development of spherically symmetric asymptotically flat data for the SSEYM system must emanate from the axis.
-Appendix C Density theorem
-As explained in Remark 1.6 (or Remark 1.2), regarding to the energy spaces in the extension theorems, we consider their equivalence, the weighted Sobolev space $H^{\prime s}_{w}$ with finite $H^{\prime s}_{w}$ norm [^4] If we take $r_{\ast}=0$ and $s=2$, the $H^{\prime s}_{w}$ space reduces to the $H^{1}$ space (1.5).
-, where
-$\|\phi\|_{H^{\prime s}_{w}}:=\sum_{k\leq s}\int_{\{r_{\ast}<r\}}\xi^{-3-\gamma}r^{-1-b}\left((\xi r\partial_{r})^{k}\phi\right)^{2}\mathrm{d}r,\quad s\in\mathbb{Z}_{\geq 0}.$
-We denote
-$\rho:=\xi^{-1}r^{-1},\quad\text{that is}\,\,\rho=(r-r_{\ast})^{-1}.$
-If $r_{\ast}=0$, $\xi=1$, and
-$\xi^{-3-\gamma}r^{-1-b}=\rho^{1+b};$
-If $r_{\ast}>0$, there is
-$\xi^{-3-\gamma}r^{-1-b}\sim\rho^{3+\gamma}+\rho^{1+b}.$
-Thus the $H^{\prime s}_{w}$ space is equivalent to the Sobolev space $H^{s}_{w}$ which has finite $H^{s}_{w}$ norm with
-$\|\phi\|^{2}_{H^{s}_{w}}:=\sum_{k\leq s}\int_{\{r_{\ast}<r\}}f(\rho)\left((\rho^{-1}\partial_{r})^{k}\phi\right)^{2}\mathrm{d}r.$
-Here $f(\rho)=(\rho^{3+\gamma}+\rho^{1+b})$ if $r_{\ast}>0$, and $f(\rho)=\rho^{1+b}$ if $r_{\ast}=0$.
-One can refer to [22, Section 3.2.4.] for more details of the $H^{s}_{w}$ space.
-In the remainder of this section, we aim to establish the density theorem in the weighted Sobolev space $H^{s}_{w}$.
-Theorem C.1.
-Smooth functions with compact support in $(r_{\ast},+\infty)$ is dense in $H^{s}_{w}\left((r_{\ast},+\infty)\right)$, $s\in\mathbb{Z}_{\geq 0}$.
-Remark C.2.
-If we take $r_{\ast}=0$, Theorem C.1 amounts to stating the density theorem in the $H^{1}$ space (1.5).
-Proof.
-The domain $(r_{\ast},+\infty)$ has an open covering
-$(r_{\ast},+\infty)=\cup_{i\geq 0}\mathcal{D}_{1i},$
-where we define
-$\mathcal{D}_{1i}:=\{x\in(r_{\ast},+\infty)|\quad 2^{i}<\rho(x)<2^{i+3}\},\quad i\geq 1,$
-and
-$\mathcal{D}_{10}:=\{x\in(r_{\ast},+\infty)|\quad\rho(x)<2^{3}\}.$
-On the other hand, letting
-$\mathcal{D}_{2i}:=\{x\in(r_{\ast},+\infty)|\quad 2^{-i+1}<\rho(x)<2^{-i+1+3}\},\quad i\geq 1,$
-we have
-$\mathcal{D}_{10}\subset\cup_{i=1}^{+\infty}\mathcal{D}_{2i}.$
-Therefore, $(r_{\ast},+\infty)$ is also covered by
-$(r_{\ast},+\infty)=\cup_{i\geq 1}\left(\mathcal{D}_{1i}\cup\mathcal{D}_{2i}\right).$
-Suppose the open covering $\cup_{i\geq 1}\left(\mathcal{D}_{1i}\cup\mathcal{D}_{2i}\right)$ which is locally finite, is associated to the partition of unity $\{\psi_{1i},\,\psi_{2i}\}_{i\geq 1}$.
-Namely,
-$0\leq\psi_{ki}\leq 1,\,\psi_{ki}\in C_{0}^{\infty}(\mathcal{D}_{ki}),\,\,k=1,2,$
-$\sum_{i\geq 1}\psi_{1i}+\psi_{2i}=1,$
-and for all $m\in\mathbb{Z}_{>0}$,
-[…]
-The existence of such partition of unity is proved as follows. Note that, $\mathcal{D}_{1i}$ and $\mathcal{D}_{2i}$ can be alternatively written as
-$\mathcal{D}_{1i}:=\{r\in(r_{\ast},+\infty)|\quad 2^{-i-3}<r-r_{\ast}<2^{-i}\},$
-and
-$\mathcal{D}_{2i}:=\{r\in(r_{\ast},+\infty)|\quad 2^{i-1-3}<r-r_{\ast}<2^{i-1}\},$
-for $i\geq 1$. Take a smooth function $\phi$ such that
-$\text{supp}\,\phi\subset\left(2^{-3},\,1\right).$
-Then we define
-$\phi_{1i}:=\phi\left(2^{i}(r-r_{\ast})\right),\quad\phi_{2i}:=\phi\left(2^{-i+1}(r-r_{\ast})\right)$
-and let
-$\psi_{1i}=\frac{\phi_{1i}}{\sum_{i\geq 1}\phi_{1i}+\phi_{2i}},\quad\psi_{2i}=\frac{\phi_{2i}}{\sum_{i\geq 1}\phi_{1i}+\phi_{2i}}.$
-Now $\{\psi_{1i},\,\psi_{2i}\}_{i\geq 1}$ is the required partition of unity.
-We can approximate $f\in H^{s}_{w}$ by the sequence $\sum_{j=1}^{K}\left(\psi_{1j}+\psi_{2j}\right)f$, $K\geq 1$, so that
-$\|f-\sum_{j=1}^{K}\left(\psi_{1j}+\psi_{2j}\right)f\|_{H^{s}_{w}}\rightarrow 0,\quad\text{if}\,\,K\rightarrow+\infty.$
-Furthermore, each of $\psi_{1j}f,\,\psi_{2j}f$, $1\leq i\leq K$, which belongs to the standard Sobolev space $H^{s}(\mathcal{D}_{1j})$, $H^{s}(\mathcal{D}_{2j})$ respectively can be approximated by $C_{0}^{\infty}$ functions.
-We thus conclude $C_{0}^{\infty}\left((r_{\ast},\,+\infty)\right)$ is dense in $H^{s}_{w}\left((r_{\ast},\,+\infty)\right)$.
-∎
+This quantity can be bounded straightforwardly through integrating (A.1a); for further details, we refer to [15].
+Theorem A.2 shows that the “first singularity” in the maximal future development of spherically symmetric asymptotically flat data for the SSEYM system must emanate from the axis.